This page: http://monitor.mikes.fi/ionclock/

TOC/References - Notes - Comb - Ion - Cavity - DAQ

Archived notes: index_2019 index_2020

2020 Archive

  • 2020-12-03: Initial measurement of repumper-shift: -2600 (C1) to -2900 Hz (C4)
  • 2020-11-18: New Ion page for clock-run data
  • 2020-11-04: clock run with new HelRes (now 14.424 MHz) and new B-beam direction, but only 800us probe
  • 2020-09-24: Move 2019 results to index_2019
  • 2020-09-24: Secular frequencies with tickler + photon-correlation
  • 2020-09-23: new 14.404 MHz Magic-RF Helical-Resonator
  • 2020-07-02: Fourier limited 4 Hz linewidth with 180ms probe-pulse. Night, no metro.
  • 2020-06-11: Second clock run, OD2 with 4ms probe, only C1, C2 peaks.
  • 2020-05-17: D5/2 lifetime, random-telegraph method with 405nm and fit
  • 2020-04-24: First clock run! 2.5ms probe, stability 2e-13/sqrt(tau_s)
  • 2020-04-01: trapping single ion again, after 405 laser broke
  • 2020-01-31: Move comb-plots to comb
  • 2020-01-15: For temperatures, see DAQ
  • 2019-12-19: Micromotion map, 3D pictures
  • 2019-11-29: Cavity mode number, Comb mode number. Cavity absolute frequency.
  • 2019-11-08: -C2 Zeeman component linewidth estimate of 800 Hz.
  • 2019-11-05: Zeeman peaks with B-shield (1.6 uT estimate).
  • 2019-11-04: Now with first B-shield. Laminated plywood + metglas.
  • 2019-07-12: Zeeman peaks in ambient (66uT estimate) B-field. First clock-spectrum!
  • 2019-07-10: Initial clock-transition search...
  • 2019-07-10: Page created.
  • 2019-06-18: ion photo and secular frequencies

  • 2021-01-24

    Temperature rise with thermal compound between helical-resonator and alu-mount.
    Adding thermal compound changed helres heating from 120 mK to 90 mK.

    2021-01-22

    Installed longer antenna: monitor voltage now 2.78Vpp! f = 14.422 MHz

    re-tune SWR close to 1.0. Monitor 60 mVpp with 10Vpp drive from siggen
    Install heatsink-compound between helical-resonator and alu-mount

    Install inner and mid-shields. Seems to fit OK. Beams go through shield-holes OK.
    Note back feet of mid-shield bottom left/right half did not fit the designed M3 holes - they loose between bottom shields and pump-chamber

    Tickler voltages measured after coupling capacitor (at eletrode) with 10x probe. Voltages are peak-to-peak

    Measured voltages seem low. Capacitors of the DC filter (fig. from OC18 guidelines) equal to approx. 0.5 or 0.25 ohms at 1 MHz and 2 MHz respectively.
    This seems to load the DDS output significantly

    Without attenuator

    f (MHz), DDS att., V

    1.0 5 distorted..
    1.0 10 7.4 mV
    1.0 15 4.4 mV
    1.0 20 2.9 mV
    2.0 5 distorted..
    2.0 10 3.0
    2.0 15 2.0 mV
    2.0 20 1.5 mV

    20 dB attenuator

    f (MHz), DDS att., V

    1.0 3 1.97 mV
    2.0 3 ~500 uV (noisy)

    2021-01-21

    Mu-Shields delivered. Inner and mid-shield unpacked - looks ok. Quite heavy duty material thickness!

    Tomorrow: Check ASE-beam alignment (analog camera), RF trap drive OK?
    15:03 no signal, dispenser off
    Intensifier gain turned up, PCO exposure ca 100-200ms
    14:51Z try to load with 2.0A. Now using image-intensifier and PCO camera.

    Field of view now larger, both top and bottom almost in same view
    NOTE: Image-intensifier input side is closer to the electrical wires. 3 wires: Orange (+5 to +15 V supply, 30 mA), White (Control voltage 0-5 V), Grey (Ground).
    Use the old PCO-camera, to image the intensifier output plane
    Install the (new) image-intensifier, this may help detect weak signals?

    12:56 no signal, dispenser off
    12:46 try load with 2.0A dispenser (2.3W initially)

    check Rb-lines, they look ok
    check 461nm with dispenser: 198.2 mA, 4.645 kOhm
    12:32Z no signal, dispenser off
    12:17Z try to load, 1.7A dispenser
    remove PMT-splitter, to get more signal to camera. Use old imaging-objective. re-align objective to electrodes
    (O-ring inside PMT-splitter-cube was loose?) blocking light thru lid is important

    11:07Z no signal, dispenser off
    10:53Z try to load, 1.7A dispenser

    09:46 no signal, dispenser off.
    09:32Z try to load, 1.7A dispenser
    with old objective A-beam window reflection is quite strong, and exits the objective-tube towards B-beam post
    new objective may have issues with beam window-reflections? Brass tubes used as spacers may need to be painted black?
    test new imaging objective - but no huge improvement in image-quality of electrode image with flashlight illumination
    realign camera to electrode image

    08:52 no signal, dispenser OFF
    08:37Z load with 1.7A dispenser
    trap RF as before. monitor waveform seems jumpy!?

    Optical power (slim sensor)
    422nm A-beam 0.8uW, B 1.1 uW, C ~1uW
    461nm A 200uW
    405nm adjust to 3.7 mW
    clearout 64 uW
    repumper 10.6 mW

    align A-beam with help of electrode reflections
    remove inner shield (outer shield was removed yesterday).

    win machine tilted? reboot.

    2021-01-20


    13:28 no signal, dispenser OFF
    adjust A-beam, based on output beam shape...(??)
    13:15 try 1.7A dispenser, 1.7W
    13:13 no signal, dispenser OFF
    13:01Z try again with 1.6A Dispenser. 422nm Rb-lines checked, 405nm power checked. A-beam output doesn't look very clean (hits some aperture somewhere?)

    12:50 no signal, dispenser OFF
    12:47 no signal in 10min..
    12:37:30Z dispenser on, 1.6A. 1.45W initially, rising slowly to 1.5 W
    Try to load with 1.6A dispenser current

    Trap RF from signal generator 10 Vpp, 14.424 MHz
    monitor 64 mVpp

    Preparing for trapping.
    461 dispenser check, OK at I = 198.3 mA, T = 4.645 kOhm.
    (New Thorlabs slim power sensor, model S130C)
    422 power: A exit: 0.85 uW, B: 1.1 uW, C: 1.0 uW. Detuning at 185 MHz
    461nm A-beam 195 uW
    405nm A-beam adjust to 3 mW
    repumper 13.5 mW (repeat 12mW) at ASE beam output
    clearout 65 uW (seems low?)
    (AW paper notebook clearout values: 75 uW on 2019-12-04, 77uW on 2019-11-25, 78uW on 2019-11-04)

    New Kubuntu linux install

    2021-01-19

    Installed new OS on new disk. Migrated all software.
    Artiq version now: v5.7144.9ed8f663. Conda enviroment: artiq-ion-v5-6.
    Ionclock ARTIQ now ident 5.7144.9ed8f663;mikes (autogenerated by M-Labs).
    New drift-DDS ARTIQ now ident 5.7144.9ed8f663;mikes-drift (built manually).
    Drift-DDS ARTIQ seems to work. Configured for external ref for both Kasli and Urukul.

    2021-01-16

    (slight improvement with lid-adjustment and heatsink)

    2021-01-14 Thursday


    13:50Z add a heatsink to the alu end-plate
    13:34Z improve slightly contact between antenna-lid and end-alumount. No visible reduction in monitor-voltage when loosening lid M3 screws (impedance matching still ok?)

    With added clamping, resonator heating is reduced 140 -> 120 mK. RF-feedthru heating ~few mK lower.
    (note the new year 2021 AD...)

    2021-01-13 Wednesday

    Install 2pcs hose-clamps that push the Cu-resonator into contact with the Alu-mounts
    12:13Z install clamps to breadboard on both sides of Alu helres-mounts

    Heating of the RF-feedthrough flange (10Vpp from signal-gen, 6+ hours settling time):
    Old mount, no sink 120 mK
    Old mount, Cu bars as heat-sink 85 mK
    New mount, 2 mounts clamped 66 mK
    New mount, mounts clamped, 2 hose-clamps around cu-helres 66 - ~few mK (results 2021-01-14!)

    Results with two alu-mounts clamped to breadboard, and no clamping-force between cu-resonator and alu-mounts.

    2021-01-11 Monday

    Install new alu-mounts that imrpvoe heat-ocnduction from Cu-resonator to breadboard

    11:26Z remove temporary copper heatsinks from helres

    2021-01-08

    (TLi) New vented rf shield (0.1 mm copper) with 6-mm slots top and bottom:
    ca 12Z move Pt100 ch. 6 (106) from rf shield to helical resonator cylinder, above adapter piece.

    With old shield:
    Signal generator (Vpp) Monitor (mVpp, average 32)
    10 66.4 (without averaging ca 70)
    9 59.5
    8 53.5
    7 47.0
    6 40.2

    rf off.
    New shield installed. RigExpert: min SWR 1.02 @ 14.4245 MHz, R = 52.1 Ohm, Q ~ 870 (SWR 2) ... 810 (SWR 3).

    rf on.
    Signal generator (Vpp) Monitor (mVpp, average 32)
    8 53.1
    10 66.8

    Both Q and monitor ratio as before within uncertianty.

    ca 13Z ready

    2021-01-07

    14:15Z RF OFF. will cool down by midnight, when computer turns ON.

    RF siggen to 10Vpp, turn ON for testing
    Menlo comb
    lab lights (partially)
    ca 09Z Turn on stuff after x-mas break

    2020-12-30

    7Vpp heating vs 10Vpp heating

    2020-12-28 Monday

    Set RF-amplitude to 7 Vpp. This should give 49% of heating compared to 10 Vpp?

    With cavity doors open it takes 5-10 days for the inner shield temperature to reach a new steady-state ca -8 mK colder than with doors closed.
    SIB04 D2.1.6 documentation gives cavity-frequency as -alpha(T-T0)**2
    with alpha=7(1)e-10 and T0=32(1) C
    For T-T0=8 mK this gives a frequency change 4.5e-14
    Compare to cavity drift of +20 mHz/s which gives 3.9e-12 / day

    Heat-sinking helres antenna-end with two Cu-rods reduces HelRes GND-end and RF-shield temperature-Deltas by almost 50%

    2020-12-23 Wednesday

    Note large ambient temperature drop affecting all channels, starting around Tuesday 19:00Z when Menlo-comb shut down (?)

    2020-12-22 Tuesday

    12:29Z install 2pcs 10mm x 40mm Cu-bars that conduct heat from the RF-input lid to the breadboard. No thermal paste between breadboard/Cu-bars/HelRes.

    2020-12-21 Monday

    13:02Z install copper foil cooling-strap. 50mm wide, only 0.2mm thick
    12:44Z Removed cooling fan.

    09:00Z add a cooling-fan at the driven RF-input end of the Helical-Resonator. Blows ambient air towards the monitor-connector.
    90mm square fan, 12 VDC

    RF on/off measurements, with 10Vpp from gen.

    2020-12-14 Monday

    laser stays locked with the door-test
    5 min fbeat 23 163 473 (27) Hz - OK
    600mA BOA gives 0.8V on cavity-transmission PD
    ca 08:20Z lock clock-laser

    Turn off lasers
    Ion loaded 2020-10-28, disappeared sometime between Friday 2020-12-11 evening and 2020-12-13 afternoon.
    try red-detuning 422nm, try 461nm - but no ion..
    re-lock Rb, try 405nm - no ion?
    07:55Z still no ion...

    2020-12-13 Sunday

    still no ion :( - leave as is
    14:13 try clock-laser, but comb probably needs warm-up time now?

    try 405nm?
    re-lock to Rb-line - no effect?
    14:00Z ion not visible!?

    2020-12-11 Friday

    leave locked
    the main-gain trimpot was initially pointing up (12-oclock) - slightly lowering gain reduces servo-bumps on beat-note and increases cavity throughput (0.75 V on detector)
    lab-door-test: let the door slam closed from ca 45-degrees open - laser stays locked OK
    11:56Z lock laser to cavity. 1348.052nm on wavemeter. 23 163 280 Hz beat note - OK(?)
    User error: FALC black on/off switch needs to be ON also!
    Check EOM-resonator with antenna-analyzer - looks OKish

    Can't lock Toptica to cavity? Reboot PDH-detector, but no luck..

    09:47Z ion still stored

    2020-12-09 Wednesday

    Tuning Toptical ECDL PDH-lock and FALC (KH).


    Turn OFF PMT
    Back to 10Vpp drive level. No overnight experiments



    Secular results:
    9.0 Vpp drive: z = 1 760 069 Hz (ID10819), r_hi = 890 069 Hz (ID10833), r_lo = 840 310 Hz (ID10828)
    9.5 Vpp drive: z = 1 863 345 Hz (ID10817), r_hi = 937 466 Hz (ID10807), r_lo = 890 328 Hz (ID10802)
    10 Vpp drive: z = 1 968 287 Hz (ID10794), r_hi = 985 241 Hz (ID10797), r_lo = 940 310 Hz (ID10798)
    Settings here: 13 dB attenuation for r-modes. 21-23 dB attenuation for z-mode
    16 kphotons per point (ca 4 s per point, with B-beam), 10 nbins


    ID10833 r_hi search 0.888-0.891 16% peak at 890 069 Hz - 90.3% of the 10Vpp result
    coarse r_hi search - weak peak at .89 MHz - radial splitting seems to increase at lower drive level?
    (the peak found below is r_lo, next search for r_hi)
    ID10828 r_hi 0.839 - 0.841 MHz, 15 dB att, 27% peak at 840310 Hz - 89.4 % of the 10Vpp result
    coarse r_hi search 20 dB (fixed) + 0 dB (artiq) att, resonance at 840 280 Hz (NOTE: this might be r_lo!!)
    ID10819 z 1.758 - 1.762MHz, 24 dB att, 24% peak at 1.760069 MHz - 89.4% of 10Vpp result
    coarse z-search: resonance at 1.76024 MHz
    set drive amplitude 9.0Vpp

    ID10817 z 1.862 - 1.865 MHz, 20% peak at 1 863 345 Hz - 94.7% of 10Vpp result
    ID10816 z 1.863 - 1.864 MHz (guided by the coarse method) - 23 dB att, peak found ok
    try searching with old method instead, faster. resonance at 1.8634
    ID10811 z 1.885 - 1.890 MHz,
    ID10810 z 1.88 - 1.885 MHz, no peak
    ID10808 z search 1.870 - 1.875 MHz, no peak
    ID10807 r_hi 0.9365 - 0.9385 MHz, 18% peak at 937 466 Hz - 95.15 % of 10Vpp result
    ID10806 r_hi search 0.935 - 0.938 MHz, peak found, higher frequnecy than expected?
    ID10805 r_hi search 0.922 - 0.927 MHz, no peak
    ID10804 r_hi search 0.927 - 0.931 MHz, no peak
    ID10803 r_hi search 0.931 - 0.935 MHz, no peak
    ID10802 r_lo 0.8895 - 0.8915 MHz - r_lo at 890 328 Hz - 94.7% of the 10Vpp result
    ID10801 r_lo search 0.888 - 0.892 MHz - 25% peak
    ID10800 r_lo search 0.894 - 0.898 MHz - no peak
    ID10799 r_lo with 9.5Vpp, search 0.892-0.893 MHz - no peak?
    set drive amplitude 9.5Vpp. Predictions by linear scaling: 1.8699, 0.9360, 0.8933 MHz

    ID10798 r_lo? search 0.939-0.941, 22% broad peak at 940 310 Hz - also ca 99.6% of the 2020-10-28 result

    ID10797 0.984-0.987, 13dB att. 14% peak at 985 241 Hz - this is 99.6% of the 2020-10-28 result
    ID10796 search lower 0.983-0.986 12% peak at 0.9853 MHz
    ID10795 r_hi? search 0.986-0.989 MHz, 15 dB attenuation - no peak?

    ID10794 21 dB attenuation 16 kphot. peak 20% (ok?). This is 99.75% of the result from 2020-10-28
    ID10793 24 dB attenuation, to get higher peak. peak now 14%
    ID10791 repeat 1.967-1.970 MHz, 30 pts, 8000 photons. 27 dB attenuation gives only 10% correlation peak
    ID10790 correlation method 1.965-1.975 MHz, 200pts
    use the coarse method to find resonance. This require removing the 20dB physical attenuator. Resonance around 1.969 MHz
    ID10786 search lower 1.94-1.96 MHz, 200pts (100Hz per point), same settings
    ID10785 try again 1.96-1.98 MH - no peak?
    abort, wrong range..
    ID10784 try to find z-secular: 2.02-2.06 MHz, 200pts (200Hz step), 27 dB att, only 4000 photons and 10 bins to speedup search

    Keysight siggen glitch-free from 10Vpp ...down..to... 3.6 Vpp should be safe

    DDS ch0 (topmost) is used as tickler. Physical 20 dB attenuator on output. Additionally 15/13/27 dB attenuation used for r_lo/r_hi/z seculars (2020-09-30)
    Previous secular result is 2020-10-28 lower radial: 0.94373, upper radial: 0.98916, axial: 1.97352, ax/rad: 2.042042

    uMotion 300kphot: A 1.4% in 51s, B 7.7% in 74s, C 14.0% in 63s
    uMotion 100kphot: A 2.5% in 18s, B 7.8% in 24s, C 13.9% in 21s
    uMotion measurement program: "Photon timestamp histogram"

    12:10Z stop collision monitor

    12:08Z ion still stored. loaded 28 Oct, so now 6 weeks?

    2020-12-07 Monday

    Camera images from 2020-12-02, with ASI 290MM camera, 2-binning mode (968 x 548 pixels) giving 2*2.9um = 5.8 um pixels.
    IMX290 has native 1936x1096 pixels, 2.9um pixel size, and 5.6 mm x 3.2 mm sensor size
    We are currently using a 400um pinhole in front of the PMT.
    Switching to a smaller pinhole, for better SNR and potentially shorter detection-window, probably requires an imaging objective with less aberrations
    narrow bar is 44 pixels or 255 um
    wide bar is 89 pixels or 516 um


    09:44Z ion still stored. surprisingly, camera-software still running

    2020-12-04 Friday

    ion still stored. collision-monitor running

    2020-12-03 Thursday

    15:43Z leave collision-monitor, with clearout ON
    counts 28/1.9 in 10ms
    states
    dark_durations
    bright_durations



    Frequency differences f - f_repumper_ON (Hz):
    +C1: 1302
    -C1: 1381
    +C2: 1246
    -C2: 1448
    +C4: 1349
    -C4: 1584


    -C4: 75 525 726, 1460, h. 38.8
    +C4: 76 392 990, 1323, h. 40.5
    -C2: 75 786 143, 1319, h. 57.1
    +C2: 76 132 898, 1167, h. 38.7
    Linecenter estimate 75 959 470
    -C1: 75 872 481, 1306, h. 41.7
    +C1: 76 046 459, 1260, h. 55.6
    ID10771: Re-initialize initial values and repeat peak search with repumper override ON.

    -C4: 75 525 800, h. 46 (fit fail, visual estimate).
    +C4: 76 393 036, 1247, h. 49.6
    -C2: 75 786 128, 1291, h. 46.2
    +C2: 76 132 939, 1126, h. 49.7
    Linecenter estimate 75 959 500
    -C1: 75 872 563, 1195, h 55.3
    +C1: 76 046 438, 1127, h. 60.3
    ID10769: Peak search with repumper override ON.

    ID10768: Peak search normally. Dark counts saved.

    Improvements to transition search program.

    -C1 75 872 597, FWHM 3304 (fit fails!?)
    +C1 76 046 390, FWHM 1797, h15.9 (fit fails!? at least wrt height) - this is shifted by -1202 Hz AOM
    ID10764 peak-search, same settings as before, 300us probe, but repumper ON
    Now, to estimate the repumper-shift, override the repumper-control to be allways ON

    B4 30.899 uT
    C4center 75 960 825
    -C4 75 527 310, 1299, h38.5
    +C4 76 394 339, 1438, h50.6
    B2 30.970 uT
    C2center 75 960 868 -C2 75 787 591, 1260, h 40.1
    +C2 76 134 144, 1337, h 56.2
    B1 30.987 uT
    C1center 75 960 811
    -C1 75 873 862, FWHM 1212, h55.1
    +C1 76 047 761, FWHM 1320, h39.5
    ID10763 peak-search again with new initial values. 300us probe

    B estimate 30.950 uT
    center-estimate 75 960 746
    -C1 75 873 900 (visually) fit fails?
    +C1 76 047 592, FWHM 1208, h 59.6
    ID10761 peak-search (old settings), 300us probe pulse
    13:17Z stop collision monitor

    Toptica PDH-lock tuning!?
    Previous settings 25 MHz output 2.8 dBm, 64.3 degrees from toptica
    With directional-coupler after photodiode, 102.9 degrees seems to work

    SYPD-52W phase-detector compares 357 MHz reference from ARTIQ DDS ch2 to the photodetector 357 MHz
    SYPD-52W spec is min 700 mV, typ 850 mV output with +7 dBm on RF1/RF2 ports
    Fiber-noise error signal now ca 7-800 mVpp slow sine-wave, when VCO tuned close to 178.5 MHz
    With -11 dBm input to the second ZFL-amp and 28 dB gain, theoretically +17dBm output expected - so we have ~10 dB extra gain here, which could be removed
    With SHP-300 HP-filter and two ZFL-amps signal is +8.2 dBm (Note Pout at 1dB compression is P1dB +7.9 dBm for ZFL amp). 2nd-harmonic increases significantly
    BOA 500mA, with one ZFL-amp gives -11 dBm 357 MHz signal. With +15VDC ZFL-gain is 28dB - corresponding to -39dBm input (~ok)
    .
    BOA-current adjust overall 674nm power, and causes variation in 357 MHz level:
    400 mA: -45 dBm
    500 mA: -41.3 dBm
    600 mA: -39 dBm
    when retro-reflector on ion-breadboard blocked, this goes down to -80 dBm - ok
    New OPA818 fiber-noise detector 357MHz signal at -42 dBm, measured after detector

    turn lights on in the back of the lab
    ion still stored - collision-monitor running
    08:18Z windows camera software crashed. restart

    2020-12-02 Wednesday

    B-beam counts 40/2.0 in 10ms
    15:28:06Z start collision monitor. B-beam, clearout ON!
    states
    dark_durations
    bright_durations


    Try new photodetector, with FDS025 ball-lens photodiode.
    Optimize fiber-noise detector.
    -42dBm
    Fiber-noise photodetector (old ET-2030A) signal levels at 357 MHz:
    0...+7.5 dBm fluctuates, with 2pcs ZFL-500LN + SHP-300 filter
    -27...-18 dBm with one ZFL amp
    -58...-48 dBm directly from detector BNC-connector

    12:25Z ion returned! switch to B-beam only
    12:20 ion seems dark/lost? try all beams ABC, and try with 405nm - but no immediate effect
    12:00Z Stop collision monitor.

    PDH-detector DC-output 118 mV
    Alignment improves transmitted intensity-monitor to 1.1 VDC
    Improved EOM tank-circuit. Now +2.8 dBm 25 MHz drive level from toptica. Phase 64.3 degrees. Remove directional coupler.

    07:40Z Open cavity door and remove EOM resonator for improvements.

    2020-12-01 Tuesday


    14:49Z close cavity door
    fbeat 23 165 431 Hz - (was 23 163 666 Hz on 2020-11-21)

    No clear camera-image of the side-bands, so they are quite weak?
    TEM00 mode at 23.2 MHz (SA marker). Side-bands in IR should be 12.5 MHz away, at 10.7 MHz and 35.7
    Can reduce amplitude in Toptica-GUI to 0 dB also? to avoid compression in the 5W PA?
    PDH-lock seems to work now, with 0 degree Phase in Toptica GUI
    comb f0-locking required a lot of adjustment of the 20 MHz f0 beat

    FALC servo settings:
    2 Input Offser range selector: both 1+2 to the right (adjustment range -0.2 to +0.2 V
    5 Slow Limited Integrator SLI (top right): switch 3 on
    6 Unlimited Integrator ULI (top left): R on, nr5 C on -> 20 (V/s)/V integrator speed
    7 reset/run switch (switched manually to lock laser to TEM00 mode)
    8 ULI offset
    9 ULI polarity: + ON (ULI non-inverting), others off -> output-range -5V..+5V
    10 Fast Limited Differentiator FLD: nr5 on.
    11 Fast Limited Integrator FLI: all off
    12 Extra slow limited integrator XSLI: nr1 on
    13 main-gain: turned up ca 1/3 from zero
    14 ULI-output: not connected (should be connected to DA2)
    15 monitor output: not connected
    16 main output, connected to laser-head current mod - we have 30dB attenuators here!?


    FALC connections:
    Invert-IN: PDH Err1 (from back of DLCPro)
    Noninvert-IN: 50R terminator
    Main-OUT: 20 dB att + 10 dB att -> ECDL laser-head current mod
    Mon-OUT: not connected
    Unlim-OUT: not connected
    DA2-output: to Fine-1 input of DLCPro

    PDH-photodetector -> ZFL-500LN G=24 dB -> Toptica In1

    12:56Z open cavity box

    ZHL-5W-1: Gain 40 dB, 1dB compression at +37 dBm output
    EOM Signal: Toptica Out1 (+12.8 dBm) -> VAT-3 attenuator (+9.8 dBm)-> ZHL-5W-1-SMA PA (+49.8 dBm?? well into compression/saturation) -> DC-32MHz LPF -> ZDC-10-1 directional coupler -> LC-tank -> EOM
    12:50Z turn on lights in the back of the lab, to work on PDH-lock (shold have little impact on PMT counts)
    Toptica software settings, PDH, channel 1: 25 MHz 12.8 dBm. Max Input level -10dBm, lock level 0V, Phase 12.9 degrees.
    12:48Z ion still stored and collision-monitor running
    camera-software crashed. new version available, so update. re-start camera software

    2020-11-27 Friday

    B-beam counts 48.0/2.0 in 10ms
    14:33:07Z start collision monitor. B-beam, clearout ON!
    states
    dark_durations
    bright_durations


    Lot of debugging. Turns out that EOM resonator is fried.
    PDH does not work. Open 'dog house' and reset detector. Detetor OK, still no lock. Detector DC level ok. Signal on detector channel while scanning. 25 MHz output to EOM ok. No signal on error channel??
    11:00Z: Stop collision monitor.

    2020-11-26 Thursday

    B-beam counts 51.2/2.0 in 10ms
    14:11:42Z start collision monitor. B-beam, clearout ON!
    states
    dark_durations
    bright_durations


    ID10740, ID10741: Rabi flopping on axial, no OD.
    ID10739: Scan for axial, no OD, att. 15dB, probe = 1.2 ms.


    ID10736: Rabi flopping on upper radial
    ID10734: Scan for upper radial, no OD, att. 15dB, probe = 0.6 ms.

    Fit: Center = 76.543174, h: 21.7, FWHM: 795 Hz.


    ID10733: Rabi flopping on +C1 lower radial, no OD, att. 15dB

    ID10730: Scan for lower radial, no OD, att. 15dB, probe = 0.6 ms.

    Fit: Center: 76.52069, h. 19.7, FWHM: 771.

    ID10729: +C1 scan with OD2, att. 15 dB, probe = 0.5 ms

    ID10728: Rabi on +C1, OD2 (incoherent), tau_pi ~1.8 ms. tau_pi/4 = 0.45 ms, tau_pi/3 = 0.6 ms

    ID10726: +C1 scan with OD2, att. 15 db, probe = 2 ms

    Fit: Center: 76.049091, h. 33.2, FWHM: 418 Hz.

    674 powers for temperature measurement: [0.0225, 0.0324, 0.0049] * max_power
    With OD2: 0.063 uW = 1.7%
    674 max power: 3.67 uW.

    ID10721: Scan around with 1.2 ms (axial).

    Fit: 77.035913, h: 7.2, FWHM: 646 Hz
    ID 10718: Scan around 76.543832 with 600 us (upper radial).

    Fit: Center: 76.54318, h: 10.2, FWHM: 831 Hz (noisy).

    ID 10716: Scan around 76.5211176 with 600 us (lower radial).

    Fit: Center: 76.52078, h: 13.28, FWHM: 633 Hz.

    tau_pi for radials: 0.3 / [0.15, 0.18, 0.07] = [2.0, 1.7, 4.3]

    Predicted sideband locations: 76.0492526 + [ 0.94373/2, 0.98916/2, 1.9735/2 ] = [ 76.5211176, 76.5438326, 77.0360126]

    ID10715, +C1 scan, probe 300 us, att. 15 dB.


    ID10712: Rabi flopping on +C1, no OD filter, 15 dB att, tau_pi = 0.3 ms.


    ID10709: +C1, Probe 260 us.
    Fit: Center: 76.0492526 MHz. FWHM: 1454 Hz.

    Previous secular result is 2020-10-28 lower radial: 0.94373, upper radial: 0.98916, axial: 1.97352, ax/rad: 2.042042
    11:57Z Stop collision monitor.

    2020-11-24 Tuesday

    B-beam counts 45/1.9 in 10ms
    15:06Z start collision monitor. B-beam, no clearout
    states
    dark_durations
    bright_durations


    Log toptica DLCpro fine1 and PC-voltage to database - to diagnoze PDH unlock-events?

    BC-tests, and measuring the clock-cycle pulses with a photodiode

    09:15Z Stop collision monitor.

    2020-11-23 Monday

    15:47Z start collision monitor, B-beam, no clearout
    states
    dark_durations
    bright_durations

    B-beam fluorescence levels 48.2/2.1 in 10ms
    15:46Z turn off clock-laser, comb-branch

    (These scans with C1-pair, 2.5 kHz span and 100 Hz step)
    ID10650 repeat 700us -> max jumps 82/100
    ID10649 try 675 us -> max jumps 76/100
    ID10648 Try 725 us -> max jumps 75/100

    Try 750 us -> 74/100 (this is too long)

    ID10643 BC line-center sweep, 21dB AOM att, 600us probe
    Try +10% to 600 us -> highest point 68/100
    Try 650 us -> 71/100
    Try 700 us -> 79/100

    Maybe BC-probing has some additional losses, and slightly longer probe would give better p_exc?

    ID10642 BC line-center sweep (fixed Z-shift), 21 dB AOM att, 540us probe
    Z-shift from peak search 0.086544 MHz
    max p 60% (why no higher, sum of peak search is 44+45 = 89 %)


    ID10641 peak search, 21 dB AOM att, 540us probe
    +C1 76 048 874, 765, h43.9
    -C1 75 875 787, 782, h45
    This gives B 30.842 uT

    ID10639 BC line-center sweep over +C1, with 21 dB AOM att and 540us
    Gaussian fit height 60.5, FWHM 870 Hz AOM
    linecenter 75 962 335 Hz
    Z-shift 0.086541 MHz


    ID10637 21dB AOM attenuation, rabiflop on +C1, tau_pi ca 540us


    ID10630 try 18 dB AOM attenuation. visually tau_pi 330 us


    ID10628 try again rabiflop with 15dB AOM att. tau_pi ~230 us (similar to Friday - OK)
    ion-breadboard 674-power-servo setpoint ca 2.18V


    ID10626 rabiflop on +C1, 23 dB att.


    ID10623 peak search, 23 dB att, 600us probe. Fourier limit 667 Hz AOM
    +C1 76 048 804, FWHM 623, h 33.5
    -C1 75 875 721, FWHM -716, h 39.3
    This is a difference of 173 kHz, so the two DDS-frequencies in BC produce AM with a period of 6 us. There are 100 of these oscillations in one 600us probe pulse - ok!?

    14:06Z toptica fine-in at -0.3V, piezo 54.2V, gain -0.5V/V

    Initial BC servo lock run, ID10623, AOM att 23 dB, 600us probe, C1,2,4 peaks
    FWHMs ca 761, 711, 695, 738, 717, 698 Hz
    BC code development...

    08:13Z Stop collision monitor

    2020-11-21 Saturday

    16:00Z leave lab.
    15:55Z start collision-monitor, B-beam, no clearout
    Fluorescence 47.7/1.49 in 10ms
    states
    dark_durations
    bright_durations


    stop writing beat-notes to DB
    turn off comb-branch, wavemeter
    stop clock-run
    turn off clocklaser
    15:51Z laser unlocked..
    15:21Z Toptica fine-in voltage at 0.2V (5% of the 4V range?)
    clock-run continues at around servo-index 1290.
    fbeat now 23 163 666 Hz, similar to yesterday, so drift has settled to old value +20mHz/s ?(which is compensated for)

    Try more ARC Factor gain (was at -0.5V/V), now to -1.5V/V. Also more FALC-gain from the pot?
    Higher FALC-gain from the pot makes the beat-note servo-bumps higher than before - which may or may not be a good thing..

    14:58Z relock clock-laser
    Toptica IP 192.168.10.52 - could collect the fine-in value over SNMP for monitoring?
    Toptica win-program shows fine-in at limit +4V (this is why laser unlocks?)
    windows camera software crashed, re-start

    2020-11-20 Friday

    leave lab.
    15:32Z restart run, ID10578, forcing FWHMs to be positive
    15:17Z ID10577 start clock-run. 15dB AOM att, 230us probe. No OD-filter. B-beam
    Cycle-time ca 63s
    working with the cavity probably gives large cavity-drifts at the start of this run?
    Initial values from the peak-search: [76.04870728 75.87555485 76.13472217 75.78966187 76.39368566 75.53063792] [ 0.00177975 -0.00154911 0.00173629 0.00160046 0.0017985 0.00183405] [0.5 0.5 0.5 0.5 0.5 0.5]
    (centers, FWHMs, gains)

    All peaks Fourier limited (within measurement repeatability?) - and good height!
    -C4 75 530 638, 1834, h44.1
    +C4 76 393 686, 1799, h44.5
    -C2 75 789 662, 1600, h50
    +C2 76 134 722, 1736, h48.9
    C1 center 75 962 131 - NOTE: this is now shifted by 3 kHz AOM after disturbing the cavity, old value from a few days ago 75 965 455
    B1 30.854 uT
    -C1 75 875 555. 1549, h 53.7 (negative value!?)
    +C1 76 048 707, FWHM 1780, h41.6
    search peaks with 230us, Fourier limit 1739 Hz AOM
    Rabiflop gives rough estimate of tau_pi ~230 us?
    Try to search for clock-spectrum, with the new higher 674nm power
    Note beat-note shifted by maybe 4 kHz, now 23 163 700 Hz

    Opening cavity doors causes temperature swings.



    Close cavity doors ca 15:00Z

    Work to increase Clock laser powers at Ion

    Zeeman AOM 15 dB attenuation, BC splitter, no OD filter
    415.8 mA on TPA driver. 167 mA on ECDL. Breadboard intensity servo set-point: 212 mV, Tomberg intensity box: about 0.86 V

    Trap breadboard:

    B beam exit: 315 nW
    B beam fiber input: 1.18 uW
    After lambda/2 waveplate before 422-combiner: 5.32 uW
    After intensity servo splitter: 12.3 uW

    Dog house (doors opened ca 11:30Z):

    Before fiber input (to ion): 53 uW

    Splitter after isolator from laser (ca 550uW from laser total):
    To ion: 100 uW, 100-105
    to cavity: 450 uW, 440-450 uW
    Splitting ratio 0.22, 0.23 - 0.24
    old Splitter is R 20/ T80 or R25/T75 ???

    230-240 uW after drift AOM before angle mirror at end-breadboard
    215 uW after EOM (end breadboard)
    21.4-21.7 uW after OD1 ND-filter (end breadboard)

    New splitter is BSX04 R90/T10 - this should give 4.5-fold more to ion, and 8-fold less to cavity
    With new splitter:
    17.25 uW transmitted towards Cavity (26-fold less than before)
    525 uW reflected towards ion

    align light from cavity side-breadboard to ion
    now get 90 uW at intensity servo splitter. This is 90/12.3 = 7.3-fold more than before
    This gives 1.45 V on the intensity-servo detector (old set-point was 212mV)

    remove OD1 filter (end breadboard)
    Try to align light into the cavity
    5.9 uW at the angle-mirror/end-breadboard. (40-fold less, input to Drift-AOM was 26-fold less, so AOM loss is ~2 fold more, needs aligning)
    6.2 uW before angle-mirror/end breadboard.

    Increase the BOA current to 600 mA - close to the max current limit. This gives 10uW towards cavity, instead of 6uW with 400mA
    got some light through the cavity. PDH-lock works. transmitted power (tomberg detector) ca 0.3 V at first
    optimize to 0.42V with 9 uW towards the cavity (this was 0.86V but with ca 20uW towards the cavity)
    ion-breadboard intensity detector now gives 2.6V


    B-beam counts now 49.4/1.7 in 10ms
    11:14Z stop clock-run to work on more 674-power to the ion
    clock-run continues..
    08:50Z clock-laser relocked
    increase FALC-gain further?
    Wavelength is too high 1348.056nm, increase piezo-voltage to lower wavelength
    08:44Z ion still stored, clock-run ongoing, but clock-laser unlocked

    2020-11-19 Thursday

    leave 2ms clock-run

    Bichromatic scan over C1 pair, varying center freq, ID10565. Looks OK.

    Bichromatic scan over C1 pair with old program: ID10559. Looks OK.

    ID10552: 23dB att scan over -C1, clock_probe = 2 ms, FWHM = 278 Hz, Fourier limit 200 Hz AOM. Center = 75.87896540932638
    ID10552: 23dB att scan over +C1, clock_probe = 2 ms, FWHM = 300 Hz, Fourier limit 200 Hz AOM. Center = 76.05204433012109
    ID10550: Rabi flopping on +C1, no OD, DDS att. 23 dB, incoherent.. tau_pi maybe 2 ms

    clock-laser in TEM00, but 1348.049nm. This is 2 cavity-modes too low frequency?. Lower piezo-voltage to get higher wavelength
    08:35Z ion still stored (now >3 weeks?). clock-laser unlocked ca 00 UTC

    2020-11-18 Wednesday

    15:55Z clock-laser unlocked again. relock. leave lab

    test clock-run continues..
    09:48 re-lock, Try to turn-up the FALC-gain a bit (potentiometer)

    clocklaser unlocked after just a few minutes, to TEM01 mode
    re-start the clockrun. same initial values as yesterday, hopefully laser has not drifted too much
    re-lock clockaser
    ion is still stored, and clock-servo running, but the clock-laser unlocked after only 1-2h yesterday
    08:20Z camera-software crashed. re-start

    2020-11-17 Tuesday

    Try a clock-run with these settings.
    The initial values from peak-search are: [76.05202085 75.87892607 76.13797895 75.79292381 76.39701993 75.53388513] [0.00030165 0.0002824 0.0003064 0.00025496 0.00026495 0.00031834] [0.5 0.5 0.5 0.5 0.5 0.5]
    [Centers] AOM MHz, [FWHMs] AOM MHz, [servo-gains] (normalized)
    2ms probe, 15 dB AOM att, OD1-filter.
    Coils (127.7, 6.1, 18.4) mA
    Cycle duration ca 66 s
    started ca 15:15Z

    Quantum jumps




    ID10493 rabiflop on +C1, 15 dB att, OD1. fit gives tau_pi 2.1ms


    -C4 75 533 885, 318, h39.3
    +C4 76 397 020, 264, h29.0
    -C2 75 792 924, 255, h40.3
    +C2 76 137 979, 306, h28.7
    B1 = 30.8437 uT
    -C1 75 878 926, FWHM 282, h30.0
    +C1 76 052 021, FWHM 302, h29.9
    Peak search with OD1 filter, 2ms probe. Fourier limit 200Hz AOM
    Initial estimate of abs(B) = 30.84 uT
    Peak search without OD filter
    -C4 75 533 785, FWHM 611 Hz, h. 46.5
    +C4 76 397 253, FWHM 492 Hz, h. 40.6
    -C2 75 792 896, FWHM 513 Hz, h. 52.7
    +C2 76 138 073, FWHM 526 Hz, h. 44.0
    -C1 75 965 453, FWHM 470 Hz, h. 38.5
    +C1 76 052 029, FWHM 460 Hz, h. 60.5
    Change coil currents to: (127.7, 6.1, 18.4) (should give 30 uT)

    Repeat peak search without OD:
    -C4 75 675 705, FWHM 519 Hz, h. 49.5
    +C4 76 255 218, FWHM 526 Hz, h. 37.8
    -C2 75 849 629, FWHM 567 Hz, h. 54.2
    +C2 76 081 281, FWHM 525 Hz, h. 40.3
    -C1 75 907 337, FWHM 480 Hz, h. 40.4
    +C1 76 023 546, FWHM 550 Hz, h. 55.2

    Peak search without OD filter, probe = tau_pi = 0.8 ms, DDS attenuation 15 dB
    -C4 75 675 753, FWHM: 600 Hz, h. 46.9
    +C4 76 255 215, FWHM: 559 Hz, h. 31.6
    -C2 75 849 621, FWHM: 492 Hz, h. 54.5
    +C2 76 081 303, FWHM: 554 Hz, h. 37.3
    -C1 75 907 386, FWHM: 521 Hz, h. 42.6
    +C1 76 023 592, FWHM: 560 Hz, h. 46.8

    No OD Rabi flopping, tau_pi = 0.8 ms, Fourier limit 1000 Hz optical, 500Hz AOM AOM sampling 100 Hz


    Repeat OD1 peak search:
    -C4 75 675 718, FWHM 346 Hz, h. 33.3
    +C4 76 255 213, FWHM 314 Hz, h. 25.0
    -C2 75 849 624, FWHM 279 Hz, h. 40.7
    +C2 76 081 273, FWHM 300 Hz, h. 29.7
    -C1 75 907 346, FWHM 261 Hz, h. 33.3
    +C1 76 023 545, FWHM 271 Hz, h. 35.4

    OD1 Rabi flopping

    ID10475: +C4, incoherent...
    ID10473: +C2, tau_pi = 1.8 ms

    ID10469: +C1, tau_pi = 2.1 ms


    -C4 75 675 609, FWHM 345 Hz, h 37.8
    +C4 76 255 293, FWHM 336 Hz, h 21.3
    -C2 75 849 490, FWHM 248 Hz, h 35.2
    +C2 76 081 273, FWHM 246 Hz, h 32.7
    B1 = 20.704 uT
    -C1 75 907 362, 238, h 29.9 (not very good fit, highest point around 40)
    +C1 76 023 551, FWHM 210, h 40.2
    peak-search with 15dB att, OD1-filter, 2ms probe. Fourier limit 200 Hz AOM
    rabiflop on +C1, 15dB att, OD1-filter. tau_pi ~2ms visually
    rabiflop on +C1, 18dB att, OD1-filter. looks incoherent, max at maybe 2ms?
    Try OD1-filter

    -C4 75 675 783, 713, h 43.3 (peak far right in scan, B-estimate is a bit too high?)
    +C4 76 255 314, 495, h 28.5
    -C2 75 849 646, 689, h 30.8
    +C2 76 081 345, 543, h 35.3
    C1center 75 965 455
    B1 20.703 uT
    -C1 75 907 363, 751, h 22.5
    +C1 76 023 547, FWHM 573, h 37.2
    peak-search with 700us probe (not rabiflop with 15 dB att and peak-search with 18 dB - should use same values)
    rabiflop on +C1 now gives much shorter tau_pi!? tau_pi ~0.7 ms with p~50%
    look for C1-peaks with 4 kHz scan (using old B1 guess for abs(B))
    try coil current suggestion (82, 4.1, 12.3) mA. these currents seem to not change abs(B) much?
    rabiflop on +C1. visually tau_pi ~1.4ms, but only p=30%
    +C2 (fit fails)
    C1center 75 965 468 Hz
    B1 20.758 uT
    -C1 75 907 221, 298, h 28.7
    +C1 76 023 714, FWHM 333, h 28.8
    no OD-filter, B-beam peak search, 18 dB att, 1.8ms probe

    clock-laser jumped to TEM01..
    not much 674nm power available? - one blu fiber blocking the 674-beam
    Rabiflop on +C1, slow rise, no oscillation at <10ms
    use OD1 filter
    B1 = 20.761 uT - initialize search with this value
    -C1 75 907200, FWHM 280, h26,4
    +C1 76 023 713, FWHM 277, h35
    ID10437 B-beam peak search. 1.8ms probe, no filter! 18 dB AOM att.
    Coils at (80, 15, 0) mA
    B-beam counts 51.2/1.8 in 10ms
    Align 674nm into B-beam

    2020-11-16 Monday

    Turn off comb-branch and PMT. leave everything else
    To be continued. No overnight experiments..

    -C4, h 26.4, FWHM 274
    +C4, h 10.3, FWHM 250 (S-state mJ asymetry same as for C2, expect higher -C4)
    -C2, h 39.0, FWHM 249 (better height now, but large +Cn/-Cn imbalance?)
    +C2, h 14, FWHM 249
    B1 = 20.698 - OK at (80,15,0) mA. (This is +0.34uT from I2 += 5mA)
    -C1 h 12.3, FWHM 458? (maybe sampling at 100Hz is too coarse?)
    +C1 h 46, FWHM 265, (peak ok in center of scan)
    B1 = 20.688 - reinitialize with this
    -C1 h 19.15, FWHM 285, why so low? (peak to the left of scan)
    +C1 h 57 (?), FWHM 240, (peak to the right of scan)
    Try (80, 15, 0) mA, with initial B1 = 20.48 uT)
    -C2 h 30.8, FWHM 263
    +C2 h 12.5, FWHM 308 (up from only 7 last run?)
    B1 = 20.3443 uT at (80, 10, 0) mA (ca +0.14 uT from I2 += 5mA)
    -C1 h 21.1, 252 Hz - why so low?
    +C1 h 52.3, FWHM 248 Hz AOM, (peak centered in scan - OK)
    initialize with 20.335 uT
    +C2 (peak is far right of the scan, no fit)
    B1 = 20.335 uT
    -C1 h 23.9, FWHM 280
    +C1 h 42.0, FWHM 310, (still to the right of the scan)
    Try again with initial B = 20.2 uT estimate
    +C1 h 52.4, FWHM 299, (peak to far right of 1.5 kHz wide scan)
    not much better C2s, try I=(80,10,0) mA
    -C2 h 23.6, 284 Hz
    +C2 h 7.2, 254 Hz
    B1 = 20.1094 uT - OK at (80, 5, 0) mA
    -C1 h 24.8, FWHM 311 (why so low?)
    +C1 h 48.3, FWHM 275 (peak in center of scan)
    B1 = 20.109 uT - better?
    B1 = 19.989 uT - wrong
    -C1 h 41.3 (peak far left in scan?)
    +C1 h 50.6
    Try coil currents (80, 5, 0) mA - to find better B-field where C1,2,4 heigts ok

    ID10407 rabiflop +C2 -> max p 20%, no clear oscillation? increasing probe to 10ms doesn't help
    ID10405 rabiflop -C1 -> 1.8ms (but max p 36% - why?)
    ID10403 rabiflop +C1 -> tau_pi 1.8ms (OK!)

    -C4 75 685 500, 302, h33.3
    +C4 76 245 426, 259, h22.5
    -C2 75 853 537, 361, h13.3
    +C2 76 077 342, 395, h7.4
    C1 75 965 450
    B1 = 20.007 uT
    -C1 75 909 309, 285, h 29
    +C1 76 021 590, FWHM 267, h 46.5
    ID10402 OD1, 1.8ms, 18 dB AOM att, peak search
    Using A-beam (B-beam probably has very low 674nm power att the moment?), and OD1 filter

    422 power now higher? A-beam counts 51.8 / 1.6 in 10ms

    422 polarization adjustment.

    SDR tracker 'tuning':
    New values: CH1: 23166800, CH2: 1666800. Measured frequencies around 200 Hz.
    Previous freq. values: CH1: 23165500, CH2: 1665500, measured frequencies around 1500 Hz.

    Drift comp back on at 10:59Z
    Drift comp off temporarily at 08:55Z

    ID 10394 (C beam): No resonance found.
    09:58Z: ID10394 24 dB att, 2.02 - 2.06 MHz, 400pts (100Hz/pt), C-beam
    Try C-beam instead (more projection on z?)
    ID10393 24 dB att, 2.03 - 2.055, 60 pts. no clear resonance
    ID10392 tickler 27 dB att, 2.040 - 2.045, 40 points. no clear resonance?
    ID10391 tickler 13 dB att, 0.942 - 0.945, 30 points. Fit 0.94439, Lorenzian FWHM (2g) 599 Hz
    ID10390 tickler 13 dB att, 0.988 - 0.991, 40 points. Visually resonance ca 0.9892. Fit 0.98919 MHz, Lorentzian FWHM (2g) 1071 Hz
    Note: now using A-beam
    Previous secular result is 2020-10-28 lower radial: 0.94373, upper radial: 0.98916, axial: 1.97352, ax/rad: 2.042042

    08:30Z uMotion ('photon timestamp histogram') 100kphot: A 7% in 23s, B 3.8% in 54s, C 2.6% in 38s

    Ion still stored
    Win-machine camera software crashed. re-start
    08:25Z stop clock-run
    Clock-laser unlocked ca 00UTC Monday morning (it was locked since Friday afternoon/evening - possibly the longest laser/comb dataset we have now)

    2020-11-14 Saturday

    15:24Z leave lab
    15:19Z start clock-run ID10386, A-beam, OD1-filter, 1800us probe - initial values from peak-search below
    cycle-duration ca 65.9 seconds

    Quantum jumps




    -C4 75 685 395, 281, h30.5
    +C4 76 245 407, 337, h20.4
    -C2 75 853 488, 330, h13.9
    +C2 76 077 293, 210, h11.8
    C1c 75 965 401
    B1 20.0066 uT
    -C1 75 909 263, 287, h33.6
    +C1 76 021 539, FWHM 242, h47.1
    OD1 peak-search with 1800us probe. Fourier limit is 222 Hz AOM
    ID10383/4 rabiflop on +C1, tau_pi ca 1800us, as expected 3-fold from 600us
    Use OD1 filter

    ID10382 rabiflop on +C1 - looks incoherent
    Use OD2 filter

    -C4 75 685 416, 633, h41.4
    +C4 76 245 422, 652, h23.7
    -C2 75 853 470, 677, h14.2
    +C2 76 077 303, 797, h8.7
    C1center 75 965 436
    B1 20.02 uT
    -C1 75 909260, FWHM 659. h 40.5
    +C1 76 021 613, FWHM 632, h55.8
    ID10381 peak-search with old settings?

    windows camera-software had crashed - restart
    14:50Z end clock-run from y-day, it ran for 1257 servo-cycles.

    2020-11-13 Friday

    ID10376 clock-run. A-beam, 600us probe, no OD-filter, 18 dB AOM att. lock-to-transition v3 program
    cycle-time ca 64s

    Quantum jumps




    A-beam polarization seems to not give as high C1,C2,C4 peaks as yesterday with B-beam? (both at B = 20 uT)
    -C4 FWHM 760, h 32
    +C4 FWHM 702, h 22
    -C2 FWHM 731, h 16
    +C2 FWHM 565, h 11
    -C1 FWHM 580, h 42
    +C1 FWHM 663, h 49
    Try 600us probe

    B1 = 20.0076 uT
    -C1 FWHM 561, h40
    +C1 FWHM 558, h 49
    Peak-search, 700us
    Try no OD-filter, 18dB AOM att, fit gives tau_pi = 700 us. Fourier limit 571 Hz AOM

    Not so great operating-point for clock run?
    -C4 FWHM 308, h28
    +C4 FWHM 328, h18
    -C2 FWHM 255, h16
    +C2 FWHM 299, h10
    C1-line-center 75 965 298
    B1 = 20.009 uT
    -C1 FWHM 260, h 34
    +C1 FWHM 250, h 43
    Peak-search with OD1, 18 db AOM att, 1.9 ms. Fourier limit 210 Hz AOM
    Try OD1 filter and 18 dB AOM attenuation. This gives tau_pi = 1.9ms for +C1

    -C1 FWHM 229, h20
    +C1 FWHM 210, h36
    ID10360 search again for C1, C2, C4
    -C4 FWHM
    +C4 FWHM fails... (no FWHM initialized..)
    -C2 FWHM 166, h 12
    +C2 FWHM 211, h 8
    -C1 FWHM 208, h 23
    +C1 FWHM 198, h 32
    Scan again with C1, C2, C4 peaks

    linecenter 75 965 314
    B1 = 20.0107 uT
    -C1 75 909 164, FWHM 190, h 26.6
    +C1 76 021 464, FWHM 174 Hz AOM, h 36.8
    ID10355 peak-search with 3.5 ms probe. 800 Hz span, 50 Hz step, 17 pts per peak

    ID10353 rabiflop with OD2 filter
    fit gives tau_pi = 3.5 ms

    ID10351 rabiflop, start_delay 1ms, with OD1 filter
    Fit gives tau_pi = 1.3 ms


    ID10350 400 ms
    13:12Z all runs completed ok!?
    ID10349 300ms
    ID10348 200ms
    ID10347 100ms
    ID10346 75 ms
    ID10345 50 ms
    ID10344 40 ms
    ID10343 30 ms
    ID10342 20 ms
    ID10341 10 ms
    ID10340 start_delay 1ms
    Try heating rate? no OD-filter

    ID10339 rabiflop on +C1 with OD2, some oscillation with maybe tau_pi ~3ms
    ID10338 rabiflop on +C1 with OD3 - looks incoherent..
    Try OD2+OD1 = OD3
    -C1 FWHM 263 Hz, h 22
    +C1 FWHM 207 Hz, h 35.2
    try turning off 674 breadboard-power servo (this changes rabi-freq and tau_pi?)
    linecenter 75 965 291
    B1 20.025 uT
    -C1 FWHM 224 Hz, h 21.2
    +C1 FWHM 236 Hz, h 30.5
    turn off extra dc-psu on upper shelf, labjack (optical table), lab laser warning light, wavemeter (upper shelf)
    linecenter 75 965 280
    B1 20.0298 uT
    -C1 FWHM 257 Hz, h18.7
    +C1 FWHM 249 Hz, h25.8
    Turn off 461nm laser PSUs
    linecenter 75 965 295
    B1 = 20.0307 uT
    -C1 FWHM 192 Hz AOM, h 25.6
    +C1 FWHM 191 Hz AOM, h 34.8
    Turn off extra laser PSUs near comb (also one AOM-PA, quite warm)

    linecenter 75 965 303 Hz
    B1 = 20.029 uT
    -C1 FWHM 221 Hz AOM, h 24.5
    +C1 FWHM 245 Hz AOM, h 28.4
    OD2, 3.4 ms probe peak-search, A-beam
    I1=80mA

    AC back to 20 Hz.

    AC to 10 Hz.

    I1 = 80 mA

    ID10330:

    Linecenter = 75.96530528070917

    B estimate = 20.03
    [76.02150422 75.90910635] [0.00023497 0.00017473]

    ID10331 repeat:

    Linecenter = 75.96530773806427

    B estimate = 20.02
    [76.02148532 75.90913016] [0.00021624 0.00022599]

    10332 repeat:

    Linecenter = 75.96531141516333

    B estimate = 20.03
    [76.02150585 75.90911698] [0.00021235 0.00022191]

    I1 = 0 mA

    ID10325:

    Linecenter = 75.96530309293766
    B estimate = 2.56

    [75.97251141 75.95809478] [0.00021166 0.00015859]

    ID10326 repeat:

    Linecenter = 75.96531035169087
    B estimate = 2.56

    [75.97249743 75.95812327] [0.00023511 0.00020959]

    ID10327 repeat:

    Linecenter = 75.96531401558983

    B estimate = 2.56
    [75.97250911 75.95811892] [0.00028921 0.00024493]

    AC to 20 Hz.

    Set I1=0 and search peaks:

    ID10322:

    Linecenter = 75.96531228253994
    B estimate = 2.56

    [75.97248611 75.95813846] [0.00019777 0.00020291]

    ID10323 repeat:

    Linecenter = 75.96529298950406
    B estimate = 2.56

    [75.97247664 75.95810934] [0.00020126 0.00022937]

    ID10324 repeat:

    Linecenter = 75.96530258717237
    B estimate = 2.57

    [75.97250705 75.95809812] [0.00021801 0.00015827]

    Peak search around I1 = 80 mA setting from Thursday.

    ID10307

    Linecenter = 75.965306
    B estimate = 20.03 uT

    [76.02151995 75.90909274] [0.00019264 0.00027958]

    ID10308 repeat

    Linecenter = 75.96530643773868
    B estimate = 20.03

    [76.021506 75.90910688] [0.00022745 0.00023002]

    ID10317 repeat

    Linecenter = 75.96531968076974
    B estimate = 20.03

    [76.02152236 75.909117 ] [0.00022078 0.00023357]

    Probe length 3.4 ms above. OD2, A-beam.
    AC to 20 Hz.
    422 needed adjusting.

    2020-11-12 Thursday

    turn off clocklaser and comb-branch
    17:41Z leave lab. no overnight experiments. A-beam cools

    -C4 FWHM
    +C4 FWHM 260, h12
    -C2 also weak
    +C2 (weak?)
    B1 = 20.030 uT
    C1 center 75 965 168
    -C1 75 908 963, FWHM 260, h18.9
    +C1 76 021 373, FWHM 232, h25.9
    ID10304 peak-search, OD2, 3.4 ms probe. 6 kHz span, 200 Hz step (large?), 17pts per peak

    ID10302 +C1 spectrum with 3.4 ms probe. Fourier limit is 118 Hz AOM
    800 Hz span, 20 Hz step, 41 points
    Gaussian fit FWHM 261 Hz AOM

    ID10300 rabiflop on +C1 center, with OD2 filter. Fit gives tau_pi=3.4 ms

    ID10297 +C1 spectrum with 1.4 ms probe
    Center 76 021 358
    Gaussian fit gives FWHM 329 Hz AOM. Fourier limit is 285 Hz AOM

    ID10295 rabiflop on +C1, with OD1 filter. fit gives tau_pi=1.4ms

    maybe too difficult.. abort
    ID10292, OD2 filter, 7ms probe, peak-search. Fourier limit is 57 Hz AOM
    Now try OD-filters, if the linewidth would be better at B=20uT??

    This looks OK, all peaks C1,2,4 decent height and near Fourier limit(?)
    -C4 75 684 879, FWHM 568, h 50
    +C4 76 245 444, FWHM 562, h 37.9
    -C2 75 853 095, FWHM 649, h 34.2
    +C2 76 077 231, FWHM 602, h 30.7
    C1 linecenter = 75 965 152
    B1 = 20.021 uT
    -C1 75 965 152, FWHM 529 Hz, h 27.3
    +C1 76 021 332, FWHM 561 Hz, h 35.6
    Try searching for C1, C2, C4 peaks with I1=80 mA. 700us probe, Fourier limit 571 Hz AOM
    4 kHz span, 200Hz step, 21 pts per peak

    I1=80mA -> fit fails, +C1 76 021 346, -C1 75 909, B1 = 20.02 uT manually
    relock to Rb line
    I1=80mA -> 422 Rb-lock lost!?
    80 mA prediction 19.9080 uT (this would be ca 10-fold the B0 field)
    I1=30mA -> 9.0904 uT
    30mA prediction 9.0994 uT
    I1 = 20mA -> fit fails +C1 = 75 984 505, -C1 = 75 945766, B1 = 6.903 uT manually
    20mA prediction 6.9426 uT - 6.9938 uT
    I1 = 10mA -> B1 = 4.7294 (center 75 965 176)
    I1 = 8mA -> B1 = 4.2825 uT (center 75 965 156)
    I1 = 6mA -> B1 = 3.847
    I1 = 4mA -> B1 = 3.419
    I1 = 2mA -> B1 = 2.9767 uT
    (leave I2=0, I3=0, for the next test)
    peak-search 700us probe, 6 kHz span, 300 Hz step. 21 pts per peak
    I2=0mA, sanity check, B1 = 2.556 uT (OK +6nT from last I2=0 result). 700us probe
    I2=8mA fails, very low peaks
    I2=6mA -> fit fails, manually B1 = 3.437 uT (low peak height) I2=6mA -> 422 Rb-unlocked?? re-lock. B1=
    I2=5mA -> fit fails? +C1 75.9742 -C1 75.956 -> 3.243 uT
    I2=4mA -> B1 = 3.0503 uT
    I2=3mA -> B1 = 2.8757 uT
    go back to I2=2mA and search for peaks
    B1 = 2.72366 uT - OK
    Note: now using A-beam (B may have issues with polarization..)

    I2 = 4 mA
    ID10233 peak-search, 700us
    fit does not work..

    ID10225 with I2 = 2 mA Rabi flopping (tau_pi = 0.7 ms)


    I2 = 2 mA
    ID10215:
    Linecenter = 75.96523481944708
    B estimate = -2.723181743382869
    [75.97287607 75.95759357] [0.0010741 0.00081136]

    I2 = 0
    ID10213:
    print:[75.97235898 75.95804976] [0.00094527 0.00096566]
    print:Linecenter = 75.9652043694033
    print:B estimate = -2.549752605433608

    Coil current - B field sensitivity (uT / mA), from 2020-08-19-coil-characterization.
    uT / mA
    1: 0.2169
    2: 0.3256
    3: 0.2085

    ID10209: -C1 No OD filter 0.4ms probe, 200 Hz AOM sampling, A/C 10 Hz
    Fourier limit 2000 Hz optical, 1000 Hz AOM
    Gaussian fit Center = 75.95816487100387, FWHM = 1300 Hz


    ID10207: +C1 No OD filter 0.4ms probe, 200 Hz AOM sampling, A/C 10 Hz
    Fourier limit 2000 Hz optical, 1000 Hz AOM
    Gaussian fit Center = 75.97242222964371, FWHM = 1000 Hz

    ID10205: A-beam Rabi flopping without OD filter, AC 20 Hz (tau_pi = 0.4 ms)


    15:47 Cooling laser back in lock, ion is still trapped. Temperature has started to decrease.
    15:31 Air conditioning was at 10 Hz for a couple of hours, resulting in a rise in temperature. Returned back to 25 Hz.
    AC 10 Hz
    ID10196: OD2 filter 3.4ms probe, 25 Hz AOM sampling, A/C 10 Hz
    Fourier limit 235Hz optical, 117 Hz AOM
    Gaussian fit FWHM = 211 Hz

    ID10194: OD2 filter 3.4ms probe, 25 Hz AOM sampling, A/C 10 Hz
    Fourier limit 235Hz optical, 117 Hz AOM
    Gaussian fit FWHM = 225 Hz


    AC off
    ID10192: A-beam Rabi flopping with OD2 filter, AC 0 Hz

    ID10189: OD2 filter 3.4ms probe, 25 Hz AOM sampling, A/C 0 Hz
    Fourier limit 235Hz optical, 117 Hz AOM
    Gaussian fit FWHM = 182 Hz

    ID10187: OD2 filter 3.4ms probe, 25 Hz AOM sampling, A/C 0 Hz
    Fourier limit 235Hz optical, 117 Hz AOM
    Gaussian fit FWHM = 161 Hz (noisy measurement)


    ID10185: OD2 filter 3.4ms probe, 25 Hz AOM sampling, A/C 20 Hz
    Fourier limit 235Hz optical, 117 Hz AOM
    Gaussian fit FWHM = 208 Hz

    ID10183: A-beam Rabi flopping with OD2 filter (tau_pi = 3.4 ms)

    ID10181: OD1 filter 1.5ms probe, 50 Hz AOM sampling, A/C 20 Hz
    Fourier limit 533Hz optical, 266 Hz AOM
    Gaussian fit FWHM = 281 Hz

    ID10177: A-beam Rabi flopping with OD1 filter (tau_pi = 1.5 ms)

    ID10174: No OD filter 0.6ms probe, 100 Hz AOM sampling, A/C 20 Hz
    Fourier limit 1333Hz optical, 667 Hz AOM
    Gaussian fit Center = 75.9724886 FWHM = 626 Hz

    ID10171: A-beam Rabi flopping without OD-filters (tau_pi = 0.6 ms)


    Switch beam back to A
    After: power at A-exit: 340 nW
    Before: power at A-exit: 40 nW, power at B-exit: 350 nW

    08:16Z: Stop collision monitor.

    2020-11-11 Wednesday

    15:54Z leave collision-monitor, '2020-11-11-run1'. B-beam, no clearout
    states
    dark_durations
    bright_durations

    B-beam fluorescence levels 45.2 / 1.4 bright/dark in 10ms

    A/C 20 Hz
    ID10163: Rabi flopping with OD2, incoherent..
    ID10160: OD2 filter 8ms probe, 20 Hz AOM sampling, A/C 20 Hz
    Fourier limit 100Hz optical, 50 Hz AOM
    Gaussian fit FWHM = 265 Hz


    ID10158: Rabi flopping with OD1

    ID10156: OD1 filter 3ms probe, 30 Hz AOM sampling, A/C 20 Hz
    Fourier limit 266Hz optical, 133 Hz AOM
    Gaussian fit FWHM = 212 Hz


    A/C 15 Hz
    ID10155: Rabi flopping with OD2, incoherent...
    ID10151: Rabi flopping with OD1

    ID10149: OD1 filter 3ms probe, 30 Hz AOM sampling, A/C 15 Hz
    Fourier limit 266Hz optical, 133 Hz AOM
    Gaussian fit FWHM = 220 Hz

    ID10147: OD2 filter 8ms probe, 20 Hz AOM sampling, A/C 15 Hz
    Fourier limit 100Hz optical, 50 Hz AOM
    Gaussian fit FWHM = 170 Hz


    A/C 10 Hz
    ID10145: OD2 filter 8ms probe, 20 Hz AOM sampling, A/C 10 Hz
    Fourier limit 100Hz optical, 50 Hz AOM
    Gaussian fit FWHM = 210 Hz

    ID10142: Rabi flopping with OD2. Incoherent...

    ID10140: Rabi flopping with OD1

    ID10138: OD1 filter 3ms probe, 30 Hz AOM sampling, A/C 10 Hz
    Fourier limit 266Hz optical, 133 Hz AOM
    Gaussian fit FWHM = 245 Hz


    A/C 0 Hz -> clear backflow of air!
    ID10136: Longer Rabi flopping with OD1

    ID10134: OD1 filter 3ms probe, 30 Hz AOM sampling, A/C 0 Hz
    Fourier limit 266Hz optical, 133 Hz AOM
    Gaussian fit FWHM = 200 Hz

    ID10128: Rabi flopping with OD1, tau_pi = 3 ms


    uMotion contrasts: A: 8%, B: 7.8%, C: 4.9%

    ID10119: OD2 filter 8ms probe, 20 Hz AOM sampling, A/C 0 Hz
    Fourier limit 100Hz optical, 50 Hz AOM
    Gaussian fit FWHM = 185 Hz

    ID10117: OD2 filter 8ms probe, 10 Hz AOM sampling, A/C 25 Hz
    Fourier limit 100Hz optical, 50 Hz AOM
    Gaussian fit Center = 75.972474, FWHM = 200.6 Hz


    B-fluorescence weak 40/10ms, rises to 80/10ms when heating B-fiber by hand..
    ion still stored
    re-lock 422 to Rb-line. Rb error signal weak and with DC-offset?
    13:30Z no ion?

    2020-11-10 evening..

    22:57Z leave lab, no experiments over night.

    +C4 higher ca 30/100, FWHM 220 Hz AOM
    C2s very weak?
    C1 l-c estimate 75 965 119 Hz
    B1_estimate 2.5806 uT ID10100 peak search with OD1, 2.4 ms

    ID10096 OD1, 2.4ms
    Fit 75 957 890, FWHM 241 (Fourier is 333 optical, 167 Hz AOM)


    ID10094 no-OD, 800us
    Fit 75 957 960, FWHM 545, h41.8

    ID10093 try OD2 and 8ms probe

    ID10092 try OD3 and 26ms probe (Fourier 15 Hz AOM, sample at 2Hz)

    ID10089 Try OD4 and 80ms probe


    center around 75 957 870
    strong 50/100 Hz sidebands may be a problem here..

    Fourier limit is 100Hz optical, 50Hz AOM
    Try to see narrow line with OD2 filter and 8ms probe
    start 23:44 local, and queue a few scans of the -C1 peak


    clock-laser unlocked...
    this give center 75 965 100 and B = 2.602 uT
    +C1 at ca 75 972 400
    -C1 at ca 75 957 800 ID10082 scan for peaks

    fbeat 23 166 671 Hz
    20:55Z no collision monitor running(?)

    2020-11-10 Tuesday

    13:45Z leave collision-monitor, ID1 '2020-11-10-run1'. B-beam, no clearout
    states
    dark_durations
    bright_durations

    B-beam fluorescence levels 66.1 / 1.6 bright/dark in 10ms

    Rabi flopping, B-beam:
    ID10078: Start delay 500 ms.
    ID10077: Start delay 100 ms.
    ID10076: Start delay 1 ms.

    Code cleanup day.

    06:45Z: Stop collision monitor.

    2020-11-09 Monday

    15:15Z leave collision-monitor, ID10005 '2020-11-09-run1'. B-beam, no clearout
    states
    dark_durations
    bright_durations

    B-beam fluorescence levels 50.5 / 1.8 bright/dark in 10ms

    NOTE: this may not work, probing-cycle doesn't seem to get longer as more delay is added? ('start_delay' in the rabi v3 program)
    ID10002, 200ms start delay
    ID10000, 150ms
    ID9998 +C1 rabi-flop, 100ms start delay
    ID9996 +C1 rabi-flop, 50ms start delay
    ID9994 +C1 rabi-flop, 1ms start delay


    -C1 75 957 774, FWHM 166(?), h24.6
    +C1 75 972 228, FWHM 270, h28.7
    Peak-search with 3ms (Fourier FWHM 133 Hz AOM) probe and OD1
    Try OD1-filter
    Rabi-flop on +C1, no OD-filter gives 1.1 ms tau-pi

    B1 2.578 uT
    C1 75 965 039 (line center)
    -C1 75 957 805, FWHM 536. h36.6
    +C1 75 972 273, FWHM 501, h40.3
    14:26Z peak-search 800us probe
    Turn off 4 GHz scope and extra PSUs causing air-currents at the setup

    tests of 422/674 polarization.

    clock-laser unlock..
    -C1 h = 29.2
    +C1 h = 35.4
    ID9963 peak-search, 2 kHz range, 100Hz step, 800us probe

    Initialize to these values, try again
    B1_estimate = 2.646 uT
    linecenter: 75 965 020
    -C1 75 957 595 (957 685 initial guess), FWHM 519, h= 33.0
    +C1 75 972 446 (971 829 initial guess), FWHM 561, h= 32.8
    ID9959 try peak-search
    set coils to zero (0,0,0) mA

    12:07Z: Stop collision monitor.

    2020-11-08 Sunday


    ca 2500s per sweep
    delete the stuck 9942 -> now started running 9944

    clock laser unlocked - relock
    ID9940 done, ID9942 clock-cycle v2 "running"-state, but nothing is happening? where did 9941 disappear?
    15:57Z ion seems still stored. artiq stuck Friday evening?

    2020-11-06 Friday

    ID9955 collision monitor, '2020-11-07-run1', starts automatically on 2020-11-08, B-beam, clearout off
    Started automatically 2020-11-08 20:10:12.807841
    states
    dark_durations
    bright_durations


    14:50Z: Start coil characterization (see O:-drive readme_2020-11-06-B-field.txt).

    Rabi flopping heating rate measurements on +C1 with B-beam.

    No OD filter, 0.8 ms pi pulse, varying start delay.

    ID9921: 1 ms


    ID9923: 100 ms


    ID9925: 500 ms


    Strange noise in beat-note spectrum (at approx 21.5 MHz.)


    2020-11-05/6 Nightlife

    leave a few wide scans running
    also leave Toptica Pressure Compensation on (but it gives abrupt jumps in the piezo voltage?)


    no narrow peaks seen - end.

    23:12Z peak-search, now 6ms probe (rabi-flop indicated this might be tau_pi)
    expect 133 Hz optical, 67 Hz AOM FWHM!?
    +C1 FWHM 174 Hz

    go back to ARC factor -0.5 V/V
    clock-laser unlocked :(
    23:02Z try again peak search

    ion went dark for a while!?
    22:59Z peak-search
    +C1 FWHM

    22:39 +C1 spectrum again..



    22:40Z peak-search
    +C1 FWHM 205

    ca22:36Z check spectrum of +C1. OD2, 8ms probe, 800 Hz span, 20 Hz step, 41 points

    gaussian fit FWHM 217 Hz

    22:21Z peak search
    +C1 75 971 829, FWHM 224 Hz
    -C1 FWHM 207
    +C2 FWHM 240
    -C2 FWHM 252
    +C4 255
    -C4 273

    uMotion 100kphot: A 7.9% in 10s, B 7.5% in 21s, C 4.6% in 16s

    22:12Z ID9874 peak search
    +C1 FWHM 199 Hz - still too wide.
    -C1 FWHM 179 Hz
    +C2 FWHM 254 Hz

    22:07Z Rabi-flop on +C1 seems to give incoherent result - still metro?

    21:53Z run peak-search again
    +C1 FWHM 233 Hz - still much too wide - metro noise!?
    +C2 FWHM 245 Hz

    Based on found C1-peaks, predicted positions are: [75.97184693 75.95769907 75.97887799 75.950668 76.00005692 75.92948908]
    Repeat this until see Fourier-limited peaks?

    21:41Z last westbound metro seems to leave Otaniemi on liikenne.live map
    OD2 filter, 8 ms probe, DB '2020-11-05-OD2'
    +C1 75 971 846, FWHM 199, h 18.2 (FWHM should be 100 Hz optical 50 Hz AOM?)
    -C1 75 957 700, FWHM 191, h 26.5
    C1 75 964 773
    B1 2.521 uT
    +C2 FWHM 235
    -C2 FWHM 279, at 75 950 649
    +C4 (no peak found?)
    -C4 (no peak found?))
    No ND filter, 800us probe. 4 kHz span, 150 Hz step -> 27 points per peak
    ID9864 search peaks
    +C1 75 971 849, FWHM 527, h 34.6
    -C1 75 957 700, FWHM 524, h 54.2
    C1 75 964 775
    B1 2.52 uT
    +C2 75 978 912, 474, h 42.5
    -C2 75 950 705, 468, h 24.2
    +C4 75 998 875, 280(?), h 3.14 (3) (maybe peak-fitting failed?)
    -C4 75 930 811, -31(??), h0.16?? (fit fails?)

    21:16Z relock clock laser. Set ARC Factor -1.0 V/V

    2020-11-05 Thursday

    For fun, leave servo running
    ID9861, DB '2020-11-05-run1'

    Quantum jumps




    -C4 75 930 138, FWHM 591, h 33.2
    +C4 75 999 442, FWHM 584, h 42.3
    -C2 75 950 929, FWHM 521, h 25.0
    +C2 75 978 636, FWHM 450, h 40.3
    (TODO: improve program to output linecenter-estimate and B-estimate after finding +C1, -C1)
    -C1 75 957 866, FWHM 550, height 43.9
    +C1 75 971 714, FWHM 505, height 39.7
    ID9860 Try peak-search, B-beam, 800us probe, to check everything works
    4 kHz span and 150 Hz step, 27 points per peak

    5 min fbeat 23 166 296 (26) Hz. ch1+ch2 sum check 100MHz (0.7 Hz) OK
    Try lower ARC Factor -0.3 V/V instead?

    uMotion 100kphot: A 7.4% in 10s, B 7.1% in 22s, C 4.7% in 16s

    No!? jumped quickly to TEM01...
    Try even more ARC Factor, to keep in lock? set to -1.6 V/V
    decrease piezo-voltage to go to higher wavelengths
    14:07 clock laser now in TEM00 at 1348.049 nm (two 500 MHz cavity-modes from the right mode?)
    camera-program crash. restart

    Leave C-beam, it has higher fluorescence than B.
    ion still stored
    06:10Z end clock-cycle run. clock-laser was unlocked at some point, now in TEM00, but wrong mode?

    2020-11-04 Wednesday

    ID9852 new try, v3-clock-servo program, DB '2020-11-04-run2' with 800us probe. (previous clock-run was 2020-06-15)
    Cycle time 64s, so for 6 peaks 10.7s per peak
    100+100 shots per peak, so 53 ms per shot
    NOTE: the 'all2' figure shifts the +/-C peaks closer to center, for better visual effect. peak-pair-centers not shifted.

    Quantum jumps




    Abort run, it was probably with 40ms probe...
    ID9851 try clock-servo run with the found peaks, DB '2020-11-04-run1'
    First clock-run with B-beam and the new helical resonator
    B-beam, t_probe 800us, AOM att. 15 dB,
    cycle time is ca 111 s, or 18.5 s per peak.
    100 B + 100 R shots per peak, so 18.5/200 = 93 ms per shot - something wrong? maybe using t_probe = 40 ms now??


    -C4: 75 930 045
    +C4: 75 999 371
    -C2: 75 950 875
    +C2: 75 978 577
    -C1: 75 957 766
    +C1: 75 971 639
    ID9850 Try again with improved values: center 75 964 751, and B = 2.479 uT
    -C1 found at: 75.957 794
    +C1 found at: 75.971 707
    From first B-beam spectrum of the day, -C1 75.9577, +C1 75.97165
    this gives center 75.964675 and B = 2.46uT (? morning estimate was 1.8 uT ?)

    Second: 'Search for transitions v1'
    First run: 'Calculate peak positions from B field v1'. This takes line-center and B-field as input.
    try automatic peak search?

    re-lock. try even more ARC Factor = -1.3 V/V
    clocklaser unlocked gain, in TEM01 mode
    A-spectrum peaks very low, max 30%

    1 hour cavity-drift vs. maser ca +17 mHz/s. Driftcomp fixed at 2*(-10.45 mHz/s) = -20.9 mHz/s
    note: two DDS-channels and splitter still connected to the Zeeman-AOM PA. Probing with DDS 3rd channel
    11:44Z ID9834 run4, clock-spectrum with A-beam. Same settings as with B and C: 80 kHz span, 150 Hz step. 800 us probe

    Fluorescence 139.8 / 2.0 bright/dark in 10ms
    Try A-beam

    09:48 ID9831 run3, C-beam spectrum, same settings as ID9820/run2


    C-beam counts 81.4/1.9 bright/dark in 10ms
    Try a C-beam clock spectrum!?

    uMotion 100kphot: A 8.9% in 10s, B 8.4% in 36s, C 6.7% in 15s

    re-lock, try even more ARC Factor: -1 V/V
    09:27 clocklaser jumped to TEM01

    Note +Cn and -Cn peaks unequal height, so B-beam linear polarisation isn't so linear?
    09:22 scan completed ok, 422 and 674 lasers stayed locked
    08:25 ID9820 re-try the clock-transition scan.
    duration 2400s = 40 min
    Visually B-field ca 1.8 uT


    B-counts 46.2/1.4 bright/dark in 10ms
    08:23 re-lock Rb-line -> strong fluorescence again!
    weak fluorescence with all beams ABC?
    abort scan..
    08:16Z ion dark moment during scan!? This gives 100% jumps...

    Note AHM3 prediction updated to -3.1e-12 from latest atomic bulletin
    ID9815 clock-transition scan around predicted center 75.964 MHz. span 80 kHz AOM, step 150 Hz. probe 800 us. randomized order
    533 frequency points, duration ca 2000s

    B-beam fluorescence a bit weak 44.3/1.4 bright/dark in 10ms

    Try more slow gain on Toptica, to keep it in lock? ARC Factor set -0.7 V/V

    set Zero coil currents: (0, 0, 0) mA

    5 minute fbeat = 23 166 211 (31) Hz
    Toptica ARC Factor = -0.5 V/V. Slew Rate Enable OFF. Set current ILD=167 mA. Piezo actual voltage 60.5V
    re-lock.
    clock-laser now in TEM03 mode, maybe ~1348.054 nm on wavemeter.
    coil runs from the night failed, clocklaser probably unlocked early in the evening Tuesday
    07:50Z ion still stored.

    2020-11-03 Tuesday

    14:00Z: Start coil characterization (see O:-drive readme_2020-11-03-B-field.txt).
    All experiments until now with zero coil currents.

    ID9784: B-beam Rabi flopping in +C1, longer range.

    ID9778: Detuning scan, B-beam
    ID9776: B-beam fluorescence: bright: 506.0 dark: 13.5 in 100 ms
    ID9775: B-beam Rabi flopping on +C1.


    ID9759: Rabi flopping with B-beam.

    674 nm power at A output: 160 nW
    422 nm power at A output: 0.6 uW
    ID9747: Clock sweep over +C1, A-beam with optimal tau_pi=0.8 ms as clock probe (Fourier 1000 Hz). Gaussian fit: center:75.9717 MHz , FWHM: 438 Hz AOM.
    ID9744: Detuning scan, A-beam.
    ID9743: A-beam fluorescence: bright: 1509.91, dark: 14.49 in 100 ms.
    ID9740: Rabi flopping on +C1, A-beam.


    Beat note 23.166161 MHz.
    ID9738: run2, A-beam, wide scan. 1ms probe, 15 dB AOM att.

    9:22Z Stop collision monitor.

    2020-11-02 Monday

    15:29Z ID9729 collision monitor, '2020-11-02-run1'
    states
    dark_durations
    bright_durations

    leave collision monitor, B-beam, no clearout. Counts 55.38 / 1.3 bright/dark in 10ms
    shut down clock-laser and comb

    no peaks found. also clock-laser jumped out of lock...
    ID9726. run6, A-beam, wider scan. 2ms probe, 15 dB AOM att.


    fbeat = 23 167 582 Hz
    ID9722 scan with A-beam (bright 120 counts/10ms). 1 ms probe
    1/100 jumps across the spectrum, but no clear peak

    B-beam gives ca 60 counts/10ms
    ID9720 clock-transition search, now with some small 674nm power in B-beam
    B-beam alignment needed after aligning the 674nm mirrors (why?)

    ID9706 run2, clock-transition search, now with AOM att 15 dB - no quantum jumps, something basic is wrong.. no 674nm in the B-beam
    Note the DDS-connections are now for bichromatic probing, with two DDS-channels via splitter to the AOM
    2020-04-16 has notes om 674nm Zeeman AOM-efficiency vs. DDS attenuation
    coil currents I1, I2, I3, all zero
    13:31Z re-lock clocklaser

    slow PDH servo, aka "ARC Fine In 1" is set to -0.5 V/V. Disabling the slew-rate limit possibly improves the servo lock?
    Laser again jumped to TEM01 mode...
    ID9705 try to find clock transition
    1ms probe pulse, 23 dB AOM att. 75.965 MHz center, 10kHz span, 100 Hz step, random sampling

    10:00Z stop collision monitor

    old AHM3-model predicts linecenter at 75965976 Hz AOM

    start clock-laser. Ilaser 170.9mA. I_amp 400 mA
    Vpiezo 60.13 V gives 1348.052 nm on wavemeter
    fbeat 23.2 MHz - but jumped to TEM01 mode.
    SDR ch2 mixing-note is at 78.5 MHz
    relock, fbeat = 23 167 608Hz

    09:36Z ion still stored, collision-monitor running

    2020-10-31 Saturday

    09:06Z re-start collision monitor. B-beam, no clearout
    DB '2020-10-31-run1'
    states
    dark_durations
    bright_durations


    uMotion 100kphot: A 5.9% in 8s, B 8.3% in 13s, C 5.8% in 13s

    B-beam counts 95.95 / 1.47 bright/dark in 10ms
    B-beam alone does give fluorescence, but weaker than A
    09:00Z stop collision monitor, re-lock 422 to Rb-line. ion appears with all beams, but not with B?

    2020-10-30 Friday

    Leave collision-monitor on. B-beam, no clearout
    Fluorescence levels 54 / 1.6 bright / dark, in 10 ms
    14:41Z start monitor.
    DB '2020-10-30-run1'
    states
    dark_durations
    bright_durations


    Weird noise problems with comb line measurement...
    About 07:20Z 422 unlocked. Stopping collision monitor.

    2020-10-29 Thursday


    Leave collision-monitor on. B-beam, no clearout
    Fluorescence levels 74 / 1.4 bright / dark, in 10 ms
    15:13Z start monitor.
    DB '2020-10-29-run1'
    states
    dark_durations
    bright_durations


    ID9674 A-beam 3200 kphot (in 195 s) - fit fails again (now first bin overpopulated?)
    ID9673 A-beam 1600 kphot (in 147 s) - note modulo(1/RF) histogram-fit fails, because last bin is overpopulated? why?
    ID9672 A-beam 800 kphot (in 105 s)
    ID9671 A-beam 400 kphot (in 53 s)
    ID9670 A-beam 100 kphot (in 13 s)
    use only A-beam where fluorescence is highest for this test
    More photons should give better statistics for the g2-histogram?

    ID9666 C-beam fluorescence levels: 619.63 bright, 13.89 dark in 100ms
    ID9665 C-beam time-stamps, 100kphot
    ID9664 C-beam detuning, 100ms, -23 +10 MHz

    ID9662 B-beam fluorescence levels: 523.26 bright, 12.02 dark in 100ms
    ID9661 B-beam time-stamps, 100kphot
    ID9660 B-beam detuning, 100ms gate, -23 +10 MHz.

    ID9657 A-beam fluorescence levels: 1001.29 bright, 13.69 dark in 100ms ID9655 A-beam uMotion 100kphot, timestamps saved. t_pulses variable should have time-stamps
    ID9655 A-beam detuning scan, 100ms gate, -23MHz +10MHz from nominal 208 MHz - no drop in fluorescence at the end? (218MHz is mid-resonance)

    Note new RF-drive is 14.424 MHz and 10 Vpp to the tuneable resonator
    Contrasts 100kphot: A 2.6%, B 3.6%, C 3.3%
    Back to: -6.2, 0.2635, 27.0703, 46.7861

    result is Optimal voltages (Common mode, diff): [-0.27877341 23.45517975 42.63258276] - but this seems wrong!?
    ranges [0.2, 0.3], [26, 28], [45.5, 47.5]
    12:07Z ID9649 run uMotion optimizer again, 100kphot, 4x4x4 grid

    12:01Z uMotion with doors, 300kphot: A 2.1% in 38s, B 2.7% in 57s, C 3.1% in 59s

    B-beam again required alignment after working with the objective-door...
    install outer shield doors

    uMotion with 2 shields, 100kphot: A 2.6% in 12s, B 2.5% in 18s, C 3.7% in 18s

    Outer shield B-holes work OKish - maybe not worth changing the old shield.
    B-beam required alignment after installing the outer shield
    inner-shield tends to push the imaging objective, not all plastic M6 screws attached
    11:30Z Install old shields

    uMotion 100kphot: A 3.1% in 11s, B 3.5% in 13s, C 2.7% in 18s
    ID9627 Optimal voltages set to (Common mode, diff): [ 0.2635054 27.0703494 46.78614235]
    -6.2 V is the new V12cm value.

    422 power at A exit is 1 uW! Reduced to 0.60 uW.
    umotions all around 35% now!
    Ion re-appeared after re-locking 422.

    07:52:00Z: Stop collision monitor. 422 unlocked during the night.

    2020-10-28 Wednesday

    15:26:00Z: Leave collision-monitor with B-beam with clearout OFF
    63 bright, 1.3 dark counts in 10ms, without the holy-blanket
    DB '2020-10-28-run-1'
    states
    dark_durations
    bright_durations


    uMotions contrasts at optimum: A: 1.6%, B: 1.7%, C: 2.2%
    ID9560: NPL uMotion minimization. Optimal voltages (Common mode, diff): [ 0.26023989 26.98180953 48.72613037]
    Seculars (correlation): lower radial: 0.94373, upper radial: 0.98916, axial: 1.97352, ax/rad: 2.042042
    Seculars (coarse): lower radial: 0.94372, upper radial: 0.98932, axial: 1.97368 ax/rad: 2.042047
    umotions: A: 7%, B: 15%, C: 28%
    Voltages: -6.2, 0.2214, 26.3703, 49.7532
    Adjust V34com by + 1.3*0.1
    V12cm to -6.2 V

    Seculars (correlation): lower radial: 0.9407, upper radial: 0.98625, axial: 1.97595, ax/rad: 2.0509
    Seculars (coarse): lower radial: 0.9406, upper radial: 0.98632, axial: 1.97632, ax/rad: 2.05127
    umotion: A: 6%, B: 10%, C: 23%
    Compensation voltages: -6.3, 0.2214, 26.3703, 49.7532

    Adjust V34diff by -18e-3 * 0.8 V
    A: 6%, B: 22%, C: 22%
    umotions: A: 6%, B: 21%, C: 22.3%
    Adjust V34cm by -1.3*0.8 V
    Set V12cm to -6.3 V

    Seculars with correlation (B-beam): lower radial: 0.95028, upper radial:0.99539, axial: 1.96657965, ax/rad 2.02149
    Seculars after adjustment with old tickler: lower radial 0.9504, upper radial 0.9955, axial 1.96736, ax/rad 2.0220566318927
    umotions: A: 9%, B: 8%, C: 8%
    Adjusted V12_common to -6.0V, and compensated accordingly by decreasing V34_common by 1.3/2 = 0.65 V
    Seculars using old tickler program: 0.9654, 1.0098, 1.9519, ax/rad = 1.975, had to use a lot of RF power (3 dB) on upper radial???

    Adjust A beam 422 power to 0.3 uW.

    uMotion: A: 4.1%, B: 4.4%, C: 5.5%
    ID9435: NPL optimization over lunch: comm-diff: [ 0.25898505 27.87242426 49.79449256]
    Optimized beam directions after umotion compensation
    Coarse micromotion compensation done by moving iteratively in cardinal directions.
    New voltages: [0.26711, 28.05231, 49.42779] A = 8%, B = 13%, C = 17%
    uMotion contrasts: A: 50%, B: 34%, C: 47%
    Previous compensation voltages (comm-diff): [-5.5, 0.3043, 27.7037, 47.2835] Optimize beam positions. All beams ok after adjustment.
    Blink test: 2 levels.
    Detuning set to 208 MHz.

    08:18:00Z start load, dispenser 1.60 A. Ion! after 3 min 30s.
    08:03:00Z start load, dispenser 1.60 A. No ion after 13 min.

    422 detuning set to 185 MHz for loading.

    2020-10-27 Tuesday


    15:17:30Z ion! in only in one minute!? at least two ions, at roughly center of camera (so camera-alignment to electrodes worked ok)
    15:16:30Z try again, after re-aligning camera without inner B-shield
    Image of inner electrodes is with higher magnification than before. Top and bottom electrodes found after some
    no ion..
    14:28:30Z start load, dispenser 1.60 A. No ion after 19 min.

    461 dispenser checked and adjusted to 198.65 mA, 4.645 kohm

    14:09Z start load, dispenser 1.60 A. No ion after 15 min.
    12:36:00Z start load, dispenser 1.60 A (1.468W). ASE on at 12:40:40Z. No ion after 14 min (12:54:15Z).

    405 spontaneously switched off during warm-up for couple of seconds!

    12:05: 422 power approx 0.6 uW in A beam exit. 461 power approx 150 uW.

    12:00Z: Check 461 dispenser. Ok with I = 199.21 mA, R = 4.645 kohm.

    Resonator adjusted to 14.424 MHz

    08:55Z KSK 105 AC increased from 20Hz to 25 Hz - to make lab cooler

    2020-10-05 Monday


    ion disappeared again while installing the inner shield..

    new B-holes in the outer shield seem ok, beam enters and exits ok
    a after some dark moments the ion re-appeared suddenly... (Rb-lock checked, ionization lasers tried, etc.)
    detuning was at 185MHz, but now back to 208 MHz as ion re-appeared

    camera focus-micrometer screw position at 6mm + 300um ("30") gives sharp ion image. removing the objective requires defocusing by 4-6 turns (2-3mm)
    lost ion (?) while installing old inner-shield (now with new B-beam holes)

    Cardinal voltage directions do not seem very stable or repeatable!? Why?
    Maybe voltage-ranges for uMotion scan needs to be optmized further?
    Cfit9327_inv =
        0.2372    2.3100    1.9737
        0.4131   -1.5132   -0.0264
        0.6644    1.4648   -3.8875
    Cfit9320_inv =
        0.0655    1.7460    1.6995
        0.2458   -0.9212    0.3168
        0.3429    1.0064   -2.8921
    inv(Cfit_2020-10-01) = (new B-beam direction)
        0.0159    0.2947    0.2067
        2.0259   -1.1843    1.3154
        1.9228   -0.1466   -3.5292
    

    at this point uMotion(100kph): A 5.2% in 16s, B 13.4% in 22s, C 3.7%% in 23s
    Optimal voltages [ 0.30428835 27.70370948 47.28350979] 08:56Z ID9327 ranges: 0.25-0.35, 27.4-27.8, 47.1-47.5 V
    Try again around the new optimum?
    At this point uMotion 100kphot: A 2.4% in 23s, B 11.1% in 27s, C 0.8% in 27s
    08:53Z done. Optimal voltages (Common mode, diff): [ 0.29125006 27.67571599 47.28898777] (V12cm -5.5 V as usual)
    08:42Z point 56/63
    07:32Z ID9320 uMotion optimization, 4x4x4 grid, ranges: 0.2-0.3, 27.25-27.75, 47.9-48.5
    Duration (22+25+25) s/point, total duration ca 77 minutes (finish by 11:45 local?)

    uMotions (100kphot): A 17.8% in 21s, B 23.5% in 24s, C 23.5% in 27s
    currently voltages: 0.2495, 27.5755, 48.2091

    B-beam bright/dark counts now 44/1.2 in 10ms
    07:26Z stop collision-monitor (this one started 20:00local on Friday, there was a second collision-monitor in the queue - intentional?)
    07:24Z ion still stored (9+ days, longest so far?)
    07:23Z ASILive camera-program on win-machince crashed. restart.

    2020-10-02 Friday

    Scheduled to start 17:00Z:
    Leave collision-monitor with B-beam with clearout OFF
    52 bright, 1.3 dark counts in 10ms, using the holy-blanket
    DB '2020-10-02-run-1'
    states
    dark_durations
    bright_durations


    422 unlocked from Rb-line? lights on or working in the lab causes this?

    10:10Z Stop collision monitor.

    07:10Z ion still stored, collision-monitor running.
    This ion was loaded 2020-09-25 Friday at 11:54:50

    2020-10-01 Thursday

    14:42Z leave collision-monitor with B-beam and both ASE-sources
    56 bright, 1.3 dark counts in 10ms, using the holy-blanket
    ID9301 DB '2020-10-01-run-1'
    states
    dark_durations
    bright_durations


    422 lineshapes:

    (compare to A-beam result from 2020-05-04)

    ID9280, A beam 1000 ms gate ASE on/off at 208.08 MHz
    ID9284, A beam cooling lineshape scan, 1000 ms gate, starting point 208.08 MHz

    ID9286, B beam 1000 ms gate ASE on/off at 208.08 MHz
    ID9287, B beam cooling lineshape scan, 1000 ms gate, starting point 208.08 MHz

    ID9289, C beam 1000 ms gate ASE on/off at 208.08 MHz
    ID9290, C beam cooling lineshape scan, 1000 ms gate, starting point 208.08 MHz

    422 powers:
    A beam power: before chamber 0.47 µW, after chamber 0.38 µW
    B beam power: before chamber 0.50 µW, after chamber 0.45 µW
    C beam power: before chamber 0.48 µW, after chamber 0.42 µW

    Vopt [ 0.24950152 27.57550502 48.20908736]
    at this point 100kphot contrasts: A 1.6 % (20s), B 1.6% (26s), C 1.1% (30s)
    The columns of inv(Cfit) are the voltage directions VA, VB, VC, that change only one beam-contrst by 100%.
    (Compare to 2020-04-23 and 2019-12 results)
    When manually searching for uMotion-null, maybe 10% or 20% of these vectors is a useful step.
    For the least-squares optimization, maybe 50% of these vectors is a good range?
    inv(Cfit_2020-10-01) = (new B-beam direction)
        0.0159    0.2947    0.2067
        2.0259   -1.1843    1.3154
        1.9228   -0.1466   -3.5292
    inv(Cfit_2020-04-23) =
        0.1235    0.3759    0.4353
        2.1165   -1.2290    1.1596
        1.8881    0.4130   -2.7941
    inv(Cfit_2019-12) = 
        0.1239    0.4418    0.4723
        2.4101   -1.6631    1.0045
        2.4619    0.5661   -3.6602
    

    08:52 done.
    08:25Z progress at 42/63. duration per point is (20+27+30) s -> will finish in ca 27 minutes
    uMotion-optimization, to find the new VA, VB, VC directions. Currently at [0.2555, 27.2649, 49.0708]
    smaller scan range 0.2-0.3, 27.0-27.5, 48.75-49.25, 4x4x4 pts scan, 100 kphotons/pt
    ID9257

    optimize beam-directions - but only slightly improved fluorescence levels

    07:20Z stop collision-monitor

    2020-09-30 Wednesday

    16:09Z leave lab
    16:04Z leave collision-monitor with the new B-beam. 50/3 bright/dark counts in 10ms gate. Both ASE-sources on
    Note first start of program gave ARTIQ-error, second run OK
    states
    dark_durations
    bright_durations


    axial might be at 1.9485 MHz.
    1.947-1.950 MHz, 27 dB att, 50 pts, A, B, C is ID9244, 9245, 9246

    measure secular frequencies and FWHMs, with new B-beam direction and good micromotion
    Lower radial seems lower than before around 0.9606 MHz
    0.960-0.962 MHz, 15 dB att, 40 pts, 40 kphot/point. A, B, C is ID9208, 9209, and 9210
    1.003-1.007 MHz, 13 dB att, 50 pts, 40 kphot/point. A, B, C is ID9211, 9212, and 9213
    1.941-1.945 MHz, 27 dB att, 50 pts, 40 kphot/point. A, B, C is ID9214, 9215, and 9216


    The Cfit matrix, and inv(Cfit) (whose columns are the va,vb,vc voltage directions) is similar to previous results, but maybe not correct, since the optimum voltage is outside the scan range
    Optimum at: [ 0.2337378 27.2069111 49.21050508]
    Contrasts at new optimum: A: 2% (22s), B: 5% (25s), C: 13.2% (32s)
    Move along Vc to [0.255465, 27.26488,49.070795]
    Now A: 2.1, B: 4.1, C: 8% !

    08:53 still running, point 37/63... duration per point (20+24+29) s = will finish in about 30 mins...
    Try automatic search around manually found (roughly AB-optimal) point 0.3876 27.653 47.4867
    ID9129 search 4x4x4 grid 0.28-0.48 26.5-28.5 46.5-48.5, 100kphot
    08:04Z start.

    Back to A-optimal point [0.2601, 29.1, 46.0387]
    From 2020-04-16 we have voltage directions that change only one beam-contrast
    These vectors change contrast by 100%, so use only a fraction of them
    VA04 0.1235 2.1165 1.8881
    VB04 0.3759 -1.2290 0.4130, useful change is maybe 10% of this, 0.1*VB = vb = (0.04, -0.12, 0.04) V
    VC04 0.4353 1.1596 -2.7941, 0.1*VC = vc = (0.04, 0.12, -0.28)
    If we move along VB or VC, this should not change the A-contrast!?
    at [0.2601, 29.1, 46.0387] A 5.8%, B 66%, C 70% - good, we are at A-optimal point
    move +vb to 0.29859 28.9771 46.08 - get: A 5.0%, B 62%, C 69% - good, A did not change
    move +vb to 0.3362 28.8542 46.1213 - get: A 5.1%, B 56%, C 70% - B-seems to change, so more of the same
    move +2vb to 0.4114 28.6084 46.2039 - get A 4.0%, B 39.6%, C 70.6 % - more vb...
    move +2vb to 0.4865 28.3626 46.2865 - get A 1.6%, B 17%, C 70.4 % - more vb
    move +2vb to 0.5617 28.1168 46.3691 - get A 0.6%, B 8.6%, C 70.6%
    The larger than previously V12diff is probably wrong, so the right move to optimize C is now -vc!?
    move -2vc to 0.4747 27.8849 46.9279 - get A 5.1%, B 2.0%, C 62% - seems to be the right direction?
    move -2vc to 0.3876 27.653 47.4867 - get A 10.6%, B 3.5%, C 47%

    C is largest, so optimize it?
    no clear minimum found by scanning small ranges around current set-point..

    at [0.2601, 29.1, 46.0387] get(100kphot): A 5.8% in 23s, B 66% in 25s, C 69% in 27s

    try scanning V34diff, with A-beam, only 10-20kphot to make it faster
    A-optimum at ca V34diff 29.1V

    Try older values from 2020-09-24, optimal values [ 0.26014207 28.08051673 46.03867207]
    Here contrasts: A 39% in 24s, B 57% in 25s, C 59% in 28s

    At the new optimal point, contrasts:
    100kphot: A 55.5 % (23s), B 70%, C 66% in 35s(blurry weak image)

    scan-range was 0.15-0.35, 26.5-29.5, 47.5-60.5, 6x6x6 grid
    uMotion optimizer suggests Optimal voltages (Common mode, diff): [ 0.20202339 24.60426097 57.7988483 ]
    07:06Z ion still stored.

    2020-09-29 Tuesday

    ID9133: Overnight NPL uMotion optimization search 6x6x6 grid, 100k photons.

    Optimize B-beam intensity and measure seculars.
    Axial: 1.942-1.945 MHz, 27 dB att, B ID9121
    Higher radial 1.0055-1.0085 MHz, 13 dB att, B ID9128
    Lower radial 0.961-0.964 MHz, 15 dB attenuation, B ID9127
    Results:



    uMotion contrasts now 65-75%!!!
    Relocated B-beam 10deg closer to camera axis, increased height 160 -> 185 mm

    Comparison Tuesday (Monday):
    Axial FWHM: A, B, C: 228, 294, 270 Hz (135, 361, 412) Hz
    Upper radial FWHM A, B, C: 995, 371, 425 Hz (899, 434, 514) Hz
    Lower radial FWHM A, B, C: 447, 671, 379 Hz, (445, 1023, 353) Hz

    Axial 1.948-1.961 MHz, 27 dB att, ABC ID9039 9040 9041
    Higher radial 1.006-1.009 MHz, 13 dB att, ABC ID9036 9037 9038
    Lower radial 0.963-0.965 MHz, 15 dB attenuation, ABC ID9033, 9034, 9035


    Repeat tickler secular sweeps, to compare with results from Monday

    06:45Z stop collision-monitor. ion still stored. lots of collision events during the night?

    2020-09-28 Monday

    14:12Z start collision-monitor
    A-beam 65 bright / 1 dark counts per 10 ms. Both ASE-sources on
    states
    dark_durations
    bright_durations


    ID9028 C-beam, 15 dB att,
    ID9027 B-beam, 15 dB att,
    ID9026 A-beam, 15 dB att, maybe slight blur, diffraction-rings visible at resonance,
    repeat lower radial with more attenuation, 40 kphotons/point, 0.962-0.965 MHz, 50 points


    ID9025 C-beam 13 dB att, c-max 23%, FWHM 739 Hz, Lorentzian fit: [2.23711492e-04 1.00740753e+00 3.69715782e-04]
    ID9024 B-beam 13 dB att, c-max 17%, sharp image, FWHM 727 Hz, Lorentzian fit: [1.50025783e-04 1.00741820e+00 3.63464740e-04]
    ID9023 A-beam 13 dB att, 1.006-1.009 in 50 steps. 13 dB att. c-max 13%, FWHM 1327 Hz, Lorentzian fit: [2.35154904e-04 1.00739944e+00 6.63353354e-04]
    ID9022 A-beam 13 dB att, higher range 1.006-1.008 MHz. this is the broadest resonance. c-max only 13%. Image stays sharp through resonance. Direction might be along the camera axis??
    ID9021 A-beam, 13 dB att
    Now higher radial 1.005-1.006 MHz scan


    ID9020 C-beam 13 dB att, slight blur (more than A-beam?), c-max 39%, FWHM 471 Hz, Lorentzian fit: [2.63558885e-04 9.63508909e-01 2.35579404e-04]
    ID9019 B-beam 13 dB att, c-max 22%, FWHM 696 Hz, Lorentzian fit: [2.31554074e-04 9.63493373e-01 3.47986538e-04]
    ID9018 A-beam 13 dB att, 0.962-0.965, 50 pts. maybe slight blur at resonance, c-max 28%, FWHM 619 Hz, Lorentzian fit: [2.37744064e-04 9.63548187e-01 3.09470795e-04]
    ID9017 lower radial 0.963-0.969 40 points, 13 dB att, A-beam

    Run all three beams with 27 dB attenuation, and more photons for better statistics
    50 kphotons
    ID9014 A-beam. 27 dB, FWHM 134 Hz
    ID9015 B-beam. 27 dB, FWHM 361 Hz
    ID9016 C-beam. 27 dB, FWHM 412 Hz


    Summary of 9005, 9008, 9011 Axial mode FWHM comparison:
    A-beam 27dB att, faint blur, c-max 15%, FWHM 222 Hz
    A-beam 25dB att, still blurry, c-max 18%, FWHM 274 Hz
    A-beam 23dB att, 660 counts / 100 ms gate. Blurry! c-max 21%, FWHM 288 Hz
    B-beam: 23dB att, 660 counts / 100 ms gate. c-max 23%, FWHM 439 Hz
    C-beam: 23dB att, 350 counts / 100 ms gate. c-max 30%, FWHM 477 Hz

    ID9013 A-beam, 27 dB att, slight blur remains, diffraction-rings faintly visible through resonance
    Lorentzian fit: [5.21626436e-05 1.94964964e+00 1.10884548e-04]

    ID9012 A-beam, try 25 dB att, still a blurry image at resonance
    Lorentzian fit: [6.73627956e-05 1.94959299e+00 1.37258450e-04]

    ID9011 A-beam, 23dB att. max contrast 21%, image is blurry at resonance! asymetric shape.
    Lorentzian fit: [1.01563463e-04 1.94961471e+00 1.44198864e-04]
    ID9008 B-beam, 23dB att. max contrast 23%.
    Lorentzian fit: [1.43360487e-04 1.94959719e+00 2.19437360e-04]

    ID9005 C-beam, 23dB att (same as ID9003). Max contrast 30%, 1s exposure camera image stays ~sharp during resonance.
    Lorentzian fit: [1.68185078e-04 1.94965074e+00 2.38378880e-04] FWHM 477 Hz (why higher than 9003 below?)
    Compare beam fluorescence-levels, and FWHM at a low tickler level. Axial mode.


    ID9003 23 dB att, [1.74523068e-04 1.94961955e+00 1.95524956e-04] FWHM 391 Hz
    ID9002 22 dB att, [1.78054036e-04 1.94969403e+00 1.87024488e-04] FWHM 374 Hz
    ID9001 21 dB att,
    ID9000 20 dB att
    ID8999 19 dB att
    ID8998 18 dB att
    ID8997 17 dB att (still asymmetric resonance) Lorentzian fit: [3.62632810e-04 1.94959744e+00 2.47898271e-04]
    ID8996 16 dB att (blurry image at resonance, asymetric/hysteretic shape?) 1.9485-1.951 MHz, 40pts. Lorentzian fit: [3.86085904e-04 1.94972414e+00 2.58231443e-04] FWHM 516 Hz
    ID8995 C-beam, axial tickler 1.948-1.951, 50pts, 20kphot. 20+15dB att. Lorentzian fit: [4.28908909e-04 1.94968264e+00 2.58508669e-04] FWHM 517 Hz
    Does tickler drive-level affect the observed FWHM?

    08:35Z stop collision-monitor.

    2020-09-27 Sunday

    15:13Z leave lab.
    15:09Z try to start collision monitor. Now it seems to start ok!? (see Friday below for data)
    A-beam, 10ms gate, 70bright/1dark counts. 2020-09-26-run1 DB tag. Both ASE-sources are on

    Lower radial FWHM (A, B, C): (572, 809, 468) Hz (Friday result 737 Hz for A-beam)
    Upper radial FWHM (A, B, C): (1295, 633, 604) Hz (Friday 1414 Hz A-beam)
    Axial FWHM (A, B, C): (315, 508, 656) Hz (Friday 259 Hz A-beam)

    Axial mode, 15 dB att, scan 1.947-1.951 MHz
    ID8989 A-beam (at resonance camera image blurry - too high tickler level?)
    print:Lorentzian fit: [2.04808225e-04 1.94986872e+00 1.57526180e-04]

    ID8990 B-beam, axial mode
    print:Lorentzian fit: [3.34966326e-04 1.94984179e+00 2.54115110e-04]

    ID8992 C-beam axial mode. (at resonance camera image blurry)
    print:Lorentzian fit: [4.57426818e-04 1.94979897e+00 3.27820460e-04]

    Higher radial mode
    ID8984 1.005-1.007, A-beam. too low range? try ID8985 1.006-1.008 MHz with 11 dB attenuatio, prelim Lorentzian fit: [2.31810846e-04 1.00751707e+00 5.25566911e-04] but scan-range still too low
    ID8986 1.006-1.008, A-beam, 11dB, 40 pts
    A-beam upper radial Lorentzian fit: [2.59266870e-04 1.00753200e+00 6.47697943e-04]

    ID8987 B-beam, 11dB att, 1.006-1.008 MHz
    print:Lorentzian fit: [1.68944514e-04 1.00753194e+00 3.16734569e-04]

    ID8988 C-beam, 11dB att, 1.006-1.008 MHz
    print:Lorentzian fit: [2.57580818e-04 1.00752836e+00 3.01935747e-04]

    ID8983 C-beam
    parameters are [A, f0, g] in the lorenzian A*(1/(g*np.pi))*( g**2 / ( ( f-f0 )**2 + g**2 ))
    FWHM = 2g
    print:Lorentzian fit: [2.39217324e-04 9.63620925e-01 2.43054054e-04]

    ID8982 B-beam
    print:Lorentzian fit: [2.20101447e-04 9.63596411e-01 4.04463962e-04]

    ID8981 tickler-sweep, A-beam, lower radial mode 0.962 to 0.965 MHz, 50 pts. 13 dB att. center. 20 kphotons/point
    print:Lorentzian fit: [2.22614323e-04 9.63654977e-01 2.85932046e-04]

    From Friday expect the modes to be at 0.963 (13 dB att), 1.006 (11 dB att), and 1.948 MHz (15 dB att)

    14:06Z ASI camera program crashed. re-start. ion seems still stored.
    No collision monitor active in ARTIQ? error message

    2020-09-25 Friday

    12:50: Overnight secular sweeps around 0.962 MHz: 40kPh, 13 + 20 dB att. Ca 10minutes per sweep. ID8879-8979 (approx. 18hr). Save Lorentzian fits to DB (desc 2020-09-26-weekend).

    Afterwards dark-state monitor on over the weekend (Starts automatically around noon on Saturday):
    Fluorescence level for A-beam ca 70 counts/10ms
    states
    dark_durations
    bright_durations


    Correlation seculars (20kPh, 20 + 13 dB): Radial: ~0.9626 MHz: ID8873
    (20 kPh 20 + 11 dB) Radial: ~1.0064 MHz: ID8874
    Axial: ~1.948 MHz: ID8875, 15 + 20 dB; ID8877: 18 + 20 dB.

    Results:



    Drop 422 power to 0.25 µW

    Seculars with old sweep program (A-beam): Radials: 0.9625 MHz, 1.0064 MHz. Axial: 1.9482 MHz

    11:39:00Z Load ion. Idisp 1.6A (1.44W). Ion! at 11:54:50.

    Keysight 33522b CH2: 14.404 MHz, 10Vpp (23.8 dBm on Siglent SA), monitor voltage Vpp: 67-68 mV (previously 71-73 mV)
    11:30Z Switch RF drive to straight from generator

    08:52Z still 3 runs of the secular-sweep in the queue... last one is ID8837.

    Results of the overnight monitoring of the lower radial secular frequency
    Secular frequency is anti-correlated with lab-temperature, because of RF PA gain-variations?

    2020-09-24 Thursday

    14:00Z: Overnight secular sweeps around 1.019 MHz: 40kPh, 13 + 20 dB att. Ca 10minutes per sweep. ID8737-8837
    Seculars with correlation:




    Correlation seculars (20kPh, 20 + 13 dB): Radial: ~1.0187 MHz: ID8718
    (20 kPh 20 + 11 dB) Radial: ~1.0606 MHz: ID8727
    Axial: ~2.0 MHz: ID8735: 18 + 20 dB more sampling; ID8734: 23 + 20 dB; ID8733: 18 + 20 dB; ID8732: 15 + 20 dB

    Seculars vs drive amplitude:

    These four drive-levels correspond to trap parameters
    qzv =
        0.4080
        0.4152
        0.4226
        0.4300
    az = 0(1)e-4 
    trap asymmetry epsilon = 0.0190(8)               
    trap voltage V0 =
      379.6754
      386.4195
      393.2914
      400.1285
    V0/Vgen = Vratio = 12680(20)
    From HelRes-characterization (Q, kappa), and PA-gain -> Vratioexp = 12283
    Vratio/Vratioexp = 1.03
    

    Timestamps saved at secular 1.019 MHz: 10k: ID8711, 200k: ID8712, 400k: ID8713, 800k: ID8714, 1M: ID8715
    try lower 422-power, 0.2 uW

    422 power in A-beam ca 450nW
    ID8708: Attenuation: 3 + 13 db, 20 kPh per point
    ID8706: Attenuation: 3 + 13 dB
    10:45Z: Photon timestamp correlation seculars (10kPh per point): Radial around 1.02 MHz: ID8705. Attenuation: 3 dB + 20 dB.
    Broad peak, too much power?

    08:50Z ID8676 uMotion-optimization over lunch
    4x4x4=64 pts. 100kphot
    Optimal voltages (Common mode, diff): [ 0.26014207 28.08051673 46.03867207]
    10:00Z: uMotion with new optimal values (100kphot) A 3.5%, B 1.4%, C 1.7%
    Previous values: [-5.5, 0.2788, 28.0406, 44.7618]

    08:42Z: uMotion (100kphot) A 23.9%, B 14.6%, C 25.1%

    axial secular freqs with old tickler program:
    60 mVpp, 2.0757 MHz (25 dB attenuation, 2.074-2.078 sweep/200pts, 50ms gate). Visible motion on camera (0.5s exposure)
    61 mVpp, 2.11372 MHz (26 dB att)
    62 mVpp, 2.15234 MHz (26 dB att)
    63 mVpp, 2.19086 MHz (23 dB att)
    and now radials: (prediction is ca 49% and 51% of axial?)
    63 mVpp, 1.07364 MHz (13 dB, broad resonance) and 1.11328 MHz (10 dB att, A-beam cooling stronger, so need more tickler?)
    62 mVpp, 1.0558 MHz (13 dB), 1.0959 MHz (10dB)
    61 mVpp, 1.0376 MHz (16 dB), 1.0785 MHz (9dB)
    60 mVpp, 1.020 MHz (15 dB), 1.0614 MHz (11dB)

    uMotion 100kphot (note RF 14.404 MHz), A 21.5%, B 13.6%, C 21.8%
    Repeat: A 20.1%, B 12.1%, C 24.6%
    06:17:00Z stop collision monitor

    2020-09-23 Wednesday, new Magic Helical-Resonator

    14:30:17Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 76 counts/10ms
    states
    dark_durations
    bright_durations

    60mVpp: 2.072 MHz axial, with 21 dB attenuation
    60mVpp: 1.019 (scan 1.01-1.03 = 20 kHz and 401 points) and 1.0608 MHz (7dB DDS attenuation)
    Lower RF-drive to 60mVpp from siggen (this was predicted drive-level for maintained secular freqs)
    Seach for secular freqs

    Reapeat uMotion after optimizations. 100kphot
    A 7.6%, B 2.5%, C 8.7%

    B gives no signal, needs alignment, after alignment 120counts/10ms
    C first 40counts/10ms, optmize to 60counts
    uMotion 100kphot (note RF 14.404 MHz), A 8.2%, B 3.6%, C 15.2%

    optimizing ASE-direction increases to 80counts/10ms! further align, to 100counts/10ms
    PMT signal: move camera in X/Y, got ca 50counts/10ms

    Focus camera in Z-direction, to get similar ion-image as before (still far left and up in image)
    increase RF siggen to 90mV, now a stable and bright image? Detuning 208MHz now.
    11:03Z try to load ion. Idisp 1.6A (1.4W)
    All beams ABC on
    11:18Z weak signal (far left in camera image)? disappears?
    Previous bias-voltages set: (-5.5, 0.2788, 28.0406, 44.7618) V
    B-coils zero (0,0,0) mA

    405nm, adjusted to 2.6mW (after warmup)

    Cooling-laser Rb-signal OK 2Vpp
    2A amp-current
    ca 0.4 uW in A-beam. 185 MHz detuning

    461nm checked against dispenser. R=4.646 kOhm, I=199.21 mA
    ca 187 uW in A-beam

    New tuneable helical-resonator on magic-RF 14.404 MHz
    Monitor signal (SMA-connector stub) is weak, 60 mVpp from signal-gen, thru RF-amp, gives 73 mVpp monitor
    Monitor components: SMA-connector, SMA L-piece, SMA-BNC adapter, probe-adapter, Keysight 10x-probe

    2020-08-26

    shut down lasers
    13:14Z ion lost while placing holy-blanket?

    11:50Z stop collision-monitor, to test 3d printed shield

    08:10Z camera software on win machine crashed? restart.

    2020-08-25

    13:26:15Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 70 counts/10ms
    states
    dark_durations
    bright_durations


    Tried to optimize B-beam fiber input coupler to align polarization axis. All beams aligned OK for power. Need to check polarization.
    09:45:00Z Stop collision monitor.

    08:34Z (visit lab to check all is ok), collision monitor running without B-shields (3D printed inner shield fitted)

    2020-08-24

    14:36:40Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 70 counts/10ms
    states
    dark_durations
    bright_durations


    B-beam aligment lost! A and C ok.

    3D printed inner shield, camera side: all beam-holes ABC need to move closer to camera hole

    reduce 422nm in A-beam to 200nW, this gives bright counts 70/10ms
    A-beam uMotion 6.7% (100kphot)
    with the high loading intensity, A-beam bright counts 172/10ms (0.5/10ms dark)
    set detuning to 208 MHz
    12:25Z start load, dispenser 1.60 A
    12:29:30 ion!

    RF settings not changed, 70 mVpp from siggen, 1.67 Vmonitor
    422nm detuning 185 MHz, maximize power to ca 0.8 uW in A-beam

    12:21Z load ion, before removing B-shields (in order to test-fit new design 3D printed dummyshields)

    06:05Z: Graphics card driver crash during weekend, stop collision monitor.

    2020-08-21

    13:23:10Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 70 counts/10ms
    states
    dark_durations
    bright_durations


    ID8292, ID8293, ID8294: Bichromatic line center scan, tau_pi = 2300 us, AOM att 23 dB, center: 75.96446 , split: 0.0052616 MHz
    2020-08-21-run11
    Center: 75.9644626 MHz, FWHM: 211 Hz AOM, height 63 (%)


    Select tau_pi = 2300 us.
    ID8290: Linecenter scan on -C1, FWHM Hz AOM
    ID8288: Linecenter scan on +C1, FWHM 207 Hz AOM (optical Fourier limit 348 Hz).

    Rabi flopping tests with increasing attenuation.
    ID8284: att 18 dB

    ID8286: att 23 dB

    AOM amplifier goes to compression with two signals -> reduce RF power.

    ID8251: AOM att 18, tau_pi = 1100 us
    ID8250: AOM att 15, tau_pi = 1100 us
    ID8249: AOM att 15, tau_pi = 800 us
    ID8248: Bichromatic line center scan. Linecenter: 75.964429 MHz, separation: 0.005395883 MHz (2020-08-21-run7), AOM att 6. Tau_pi = 800 us
    ID8244: Linecenter scan on -C1 (2020-08-21-run6)
    ID8243: Linecenter scan on +C1 (2020-08-21-run5)
    ID8241: Rabi on +C1:

    ID8237: Bichromatic line center scan with 2.5 kHz separation
    ID8236: Bichromatic line center scan. Linecenter: 75.964397 MHz, separation: 0.005428 MHz
    ID8224: Rabi on +C1, corresponding to AOM att 18 dB. tau_pi = 1100us:

    ID8229: -C1 linescan with tau_pi = 800 us
    ID8227: +C1 linescan with tau_pi = 800 us
    ID8224: Rabi on +C1:

    Clock-spectrum scans.
    ID8211: I: (0,0,5) mA (green), ID8213: I: (0,0,-5) mA (red).

    08:09Z Stop collision monitor.

    2020-08-20

    13:21Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 70 counts/10ms
    states
    dark_durations
    bright_durations

    2020-08-19

    13:30Z Start coil characterization. Description on O:-drive (2020-08-19-coil-characterization).
    10:03Z Stop collision monitor.

    2020-08-18

    13:21Z start collision monitor with clearout on!
    Fluorescence level for A-beam ca 70-90 counts/10ms
    states
    dark_durations
    bright_durations


    ID8078: Rabi on +C2:

    ID8076: clock-spectrum scan. C1s are maybe very low, C2s visible. B-field direction has changed over summer?
    Fit to previous spectrum gives 1.7 uT


    C-contrast 1.3 %
    B-fluorescence changes from 30 to 110 counts / 10ms when heating the B-fiber with one finger
    A-beam uMotion 3.7% (much better than before). So this is a single ion!?
    detuning to 208 MHz -> much brighter image
    12:38Z load with 1.60A
    12:43:40Z signal!

    422-detuning to 185 MHz
    check 461nm against helper-dispenser, 199.18 mA, 4.646 kOhm.

    seem to have trapped one bright one dark ion? uMotion correlation-histogram not sine-shaped. 60% contrast. Bright ion can sometimes suddenly move sideways on camera image
    12:12:30Z try to load ion, dispenser 1.65 A, 1.6W, Ion! after 5 min.
    12:01:00Z try to load ion, dispenser 1.6 A, 1.5W, no ion after 10 min.
    11:49:30Z try to load ion, dispenser 1.6 A, 1.5W, no ion after 10 min.

    100kphot: A 68%, B 66% in s, C 68%. Maybe multiple ion? Blink test shows only 2 levels. Maybe dark isotope?
    Check 461 dispenser, OK, I = -198.93 mA.
    10:26:00Z try to load ion, dispenser 1.65 A, 1.6W, Ion! at 10:35Z (9 min)

    Maximize 422 power (now 0.4 uW at A exit) and try again:
    10:12:00Z try to load ion, dispenser 1.60 A, 1.50W, no ion after 10min.

    09:58:00Z try to load ion, dispenser 1.60 A, 1.50W, no ion after 10min.
    Ion lost around 8:00Z

    ID8025 No ND line-scan with 600us probe
    50 kHz span, 150 Hz step, duration ~1500s

    06:34Z stop collision monitor

    2020-08-17 Fall-season opening

    13:55Z start collision monitor
    422-amp to 1.8 A, fluorescence level for A-beam ca 70-90 counts/10ms
    states
    dark_durations
    bright_durations


    13:30Z: ID8015: Clock sweep at expected center, 25 kHz, no filter, 150 Hz step. C1 separation: 75.9747 - 75.954 = 20.7 kHz AOM, 41.4 kHz optical.

    B field = 7.4 uT ?!?!? (6-fold larger than old field of 1.3 uT)

    13:00Z: Micromotion contrasts:
    Compensation voltages: V1 = -5.7788, V2 = -5.2212, V3 = -16.7212, V4 = 72.8024, Com-Diff: (-5.5, 0.2788, 28.0406, 44.7618)
    100kphot: A 2.9% in 12s, B 2.3% in 14s, C 3.3% in 14s

    12:32:30Z try to load ion, dispenser 1.60 A, 1.50W
    12:37:00Z ion! only 4min 30s loading time


    ASI-camera usual settings 968x548 pixels (Bin2), High-gain, 1s exposure, histogram stretched/zoomed
    bias-voltages not touched
    coil-currents not touched, all zero (I1, I2, I3)

    RF-drive 70mVpp 16.770 MHz, monitor 1.67 Vpp

    12:29Z check A-beam power levels.
    405nm 2.7 mW (reduced from max 6-7 mW)
    461nm 200 uW
    422nm 0.4 uW

    07:40Z Current driver changed for 461 nm. Good parameters: I = -198.68 mA, T = 4.646 kohm.

    2020-07-03

    13:05Z stop experiment for summer break

    300kphot: A 7.1% in 85s
    300kphot: C 107s, last bin of histogram overpopulated!? why?? can't fit sine-wave to histogram
    100kphot: A 7.6% in 28s, B 4.7% in 71s, C 11.4% (!) in 42s

    Check micromotions over lunch - did not work? - low counts only 40/10ms for A-beam?

    09:30Z ion still stored. clock-laser in TEM01.

    2020-07-02






    ND-filterRabi-fTau_piDecoherence rateT_ion
    no filter709 Hz
    ND1256 Hz2.1 ms0.807 Hz0.598 mK
    ND2109 Hz5.8 ms12 Hz ?2.3 mK ?
    ND3 ID795928 Hz20.6 ms (note 10x ND1 result)11 Hz1.15 mK
    ND3 ID796131 Hz20.5 2.6 Hz2.1 mK
    ND3+ND1 ID796710 Hz61.8 ms0? Hz2.2 mK

    leave scans with ND3+ND2 and 180ms over night
    (ion has dark moment now - let's hope it recovers...)


    ID7970 ND3+ND2 flop-on-center
    low Pexc, not hitting the resonance??

    ID7969 ND3+ND2 line-scan with 180ms probe
    40 Hz span, 1Hz step, 41pts * 22s/100shots = duration 900s
    only 24/100 jumps at most?


    ID7967 ND3+ND1 flop 2nd try at lower freq 75.968039, 100shots

    ID7965 ND3+ND1 flop-on-center 75.968045, from 0.05 to 300ms in 20 steps
    low excitation? not hitting the carrier frequency?

    ND3+ND1 predict 21ms*sqrt(10) = 66 ms tau_pi
    ND3+ND1 line-scan with 66ms probe, Fourier limit 12 Hz
    100Hz span with 1 Hz step, 101pts * 11s/100shots = duration 1100 s = 18 minutes
    optimistically central feature ca 6-7 Hz AOM wide!? lots of side-bands.. at 50 and 100 Hz?


    ID7961 ND3 flop to 150ms, 20pts

    ID7959 ND3 flop-on-center, 0.05 to 40ms in 20 steps (maybe too dense sampling, maybe longer pulses needed?)


    ID7957 ND3 line-scan, 15 ms probe
    300 Hz span, 5 Hz step, 59 pts
    now get side-bands, central feature maybe 5pts = 25 Hz AOM wide. Fourier limit for 15ms is 53 Hz
    Gaussian fit does not work for this data.


    ID7954 ND2 flop on center 75.968055, to 40ms


    ND2 predicted tau_pi = 1800*sqrt(10) = 5.7 ms
    ID7949 line-scan with ND2 and 5 ms probe (Fourier limit 160 Hz)
    500Hz span 16 Hz step, 31 pts.
    Fit FWHM 95 Hz AOM, center 75.968055


    ID7947 ND1 flop on center 75.968052, 50us to 30 ms


    (23:35 local, no trains?) ID7945 +C1 with ND1 1800us probe
    1 kHz span, 30Hz step, 33pts
    Gaussian fit 222 Hz FWHM AOM (at Fourier limit)


    rabi-osc-fit now with fixed A=0.5 (?)

    possibly to be continued after 23:30 local...

    ID7942 wide scan with 1.8ms probe
    duration is ca 8000s = 2+ hours
    50 kHz span with 30 Hz step, 1667 points
    +C1 75.96807
    -C1 75.96059


    ID7917 flop-on-center, ND1-filter, to 3 ms probe pulse
    predicted rabi-freq is 709Hz/sqrt(10) = 224 Hz, fit gives 258 Hz


    ND1 filter, now tau_pi maybe 1800us?
    line-scan to check +C1 center frequency:
    fit 75.96803


    ID7910 flop to 12ms, 300 shots.


    adjust 422-power down to get ca 60counts/10ms
    Previous flops with ca 120 photons/10ms. Try a lower cooling-laser power, to get colder ion? (but fit-function now limits lowest Tion to 0.5mK...)

    ID7906 probe pulses to 12ms, with 300shots per pulse for better statistics


    ID7904 probe pulses to 6 ms

    ID7900 flop-on-center
    behold, the new rabi-osc-fit-figure:


    ID7897 +C1 line-scan
    Fit 75.967965, h46, FWHM 719 Hz


    ID7896 Try C1/2/4 peak search, with B=1.31uT and linecenter=75.964275
    +C1 75.968047, h44, FWHM 710 Hz AOM
    -C1 75.960564, h37, FWHM 739 Hz AOM
    +C2 75.971737, h41, FWHM 618 Hz
    -C2 75.956842, h50, FWHM 611 Hz
    +C4 75.982886, h16, FWHM 743 Hz (low peak height)
    -C4 75.945704, h16, FWHM 870 Hz (low peak height)

    From the 30 kHz scan over lunch:
    C1 pair split by 7.35 kHz AOM -> B1 = 1.310 uT
    C2 pair split by 14.76 kHz AOM -> B2 = 1.319 uT
    C3 pair splot by 29.7 kHz AOM -> B3 = 1.325 uT

    C3-pair just about fits into the 30 kHz span
    No ND-filter, 600us probe, search for clock-transition
    ID7894 predicted linecenter 75.9642, scan 30 kHz span with 150 Hz step
    201 pts to scan, ca 4.6 s per point


    reinstall the 2nd shield. install the doors.

    uMotions seem ok..
    re-run B contrast: 1.2% in 72s
    B-beam histogram fit fails for some reason? B gives only 20 counts/10ms gate - reposition the holy-blanket
    uMotion 300kphot: A 1.8% in 62s, B ?? % in 107s, C 3.7% in 72s

    re-lock clocklaser, fbeat now ca 23165894 Hz (stdev 24Hz), 100 MHz checksum stdev 0.0003 Hz
    ion is still stored

    ASILive camera program had crashed
    Comb f0 lock had unlocked during the night. Clock-laser in TEM01 mode
    08:03Z stop collision-monitor

    2020-07-01

    14:35Z start collision monitor
    states
    dark_durations
    bright_durations


    stopped collision-monitor at some point.. (sometime before 10:51Z)

    2020-06-30

    13:40Z start collision monitor. A-beam, no clearout, 70/2 bright/dark in 10ms. Note without 2nd B-shield lab lights need to be OFF to get low dark counts.
    states
    dark_durations
    bright_durations


    13:28Z uMotion 300kphot A 0.8% in 46s, B 2.7% in 64s, C 2.5% in 75s

    07:20Z stop collision-monitor
    07:19Z ion still stored

    2020-06-29


    ca 14:30Z ID7832 2020-06-29-run1 is collision monitor over night
    states
    dark_durations
    bright_durations


    Only A-beam. no clearout. With lab-lights off and holy-blanket over trap: 70 bright, 2 dark counts in 10 ms
    Leave collision-monitor over night

    investigate 2nd B-shield DC/AC shielding with B-meter. overnight measurement in Z-direction with 2nd-shield and mu-metal-plate bottom.

    12:51Z ion signal!? when adjusting 422-power slightly up!? (should use max 422-power allways when trap loading?)
    12:31:30Z try dispenser 1.60A
    12:46 no signal in 10+ minutes...

    check Rb-lock, check 461nm current 197.25mA
    12:12Z dispenser 1.55 A, 1.36 W
    12:25 no signal in 10+ mins...

    Load ion before removing 2nd B-shield for measurements
    A-beam power levels
    422nm 0.35 uW (208 MHz detuning)
    461nm 143 uW
    405nm 2.7 mW

    2020-06-18 Thursday

    11:18Z: Stopped B-field monitor

    2020-06-17 Wednesday

    13:24Z: Starting over-the-weekend magnetic field monitor:

    09:00Z: Ion lost. Switching off lasers.
    08:00Z: Clock laser was spontaneously unlocked during the night. Relocking.

    2020-06-16 Tuesday

    Data:
    run3
    run4
    run5
    run6
    run7
    run8
    run9
    run10
    run11
    run12
    run13
    run14
    run15
    run16
    run17
    run18
    run19
    run20
    run21
    run22
    run23
    run24
    run25
    run26
    run27 in progress 2020-06-17 ca 08 local
    run28
    run29
    run30
    fancy 3d plot of all the spectra?






    15:15Z leave lab.
    15:13Z start Bmeter on win-machine

    With 500Hz span and 2 Hz step (251pts) each scan takes ca 2100s
    run3 is the first one, queue about 30 scans overnight


    Leave this as a repeated scan over night to monitor noise?

    Rabi-flop maybe shows good Pexc with a long 35ms pulse?
    ID7787 scan with 2Hz step and 35ms probe

    Try ND3 and 10ms, only 16/100 jumps at most

    try to get a narrow linewidth?
    ND2 and 4ms gives 111 Hz AOM for +C1

    seems hard to find a coil-setting where C1,2,4 are strong..

    ID7775 rerun peak-search with B=1.436 uT and wider 5 kHz scan, hopefully to find C4
    5 kHz span, 150 Hz step, 33 pts/peak
    +C1 h35, 719 Hz
    -C1 h35, 753 Hz
    +C2 h48, 607 Hz
    -C2 h47, 670 Hz
    +C4 weak 20/100 central peak with two sidelobes..
    -C4 strange, multiple peaks

    ID7773 peak-search with (0,0,0) currents
    +C1 75.967400, h48, 624Hz (moved closer to linecenter?) -C1 75.959360, h35, 706 Hz. From C1-pair B=1.436 uT now
    +C2 75.971367, h47, 668 Hz (peak at far left of scan-range)
    -C2 75.955355, h46, 707 Hz (far right in scan-range), ca 1.43 uT from C2-pair
    +C4 no peak found, program aborts at Gaussian fit..

    Try zero coil-currents (0,0,0), so far we had (+1,0,0) mA from yesterday

    ID7771 re-run peak-search
    +C1 75.967936, h48, 627Hz (now at position where ID7765 found it!)
    -C1 75.958822, h46, 666Hz
    +C2 75.972437, h43, 609Hz (this one falsely identified as +C1 below in ID7768...)
    -C2 75.954234, h48, 672Hz
    +C4 low peak-height 16/100 jumps, with big sidelobes, no good gaussian fit
    -C4 center lower than sidelobes??

    B-estimate below is wrong (by 2x), field should be 1.625 uT ?

    ID7768 auto-search for peaks. 600us probe (Fourier 667 Hz AOM), 3kHz span, 150 Hz step, 21 pts per peak
    +C1 75.972372 (?), h55, 618 Hz (this may search for the peaks in the wrong place now??)
    -C1 75.954280, h49, 628 Hz
    +C2 75.981584, h40, 703 Hz
    -C2 75.945219, h36, 727 Hz
    -C4 no peak found
    +C4 no peak

    ID7765 line-scan for transition, no ND-filter, 600us probe
    the mu-plate may have changed the field a lot!?
    +C1 at 75.968
    -C1 at 75.9589
    From these peaks, B-field now 3.25 uT

    Improved B-meter data-collection program on win-machine. Probe is now on top of old trap, rouglhy vertically
    Note that meter has an arbitrary DC-offset applied now - absolute field not measured

    13:29 Try a mu-metal plate on top of the setup!?

    2020-06-15 Monday

    15:49Z leave lab.



    Quantum jumps



    15:44Z start clock-servo (v3 program), 4 ms probe
    Cycle time ca 67s

    ID7763 peak search, 4ms probe (Fourier 100 Hz AOM), 400 Hz span, 20 Hz step, 21 pts
    +C1 75.966944, h44, 117 Hz (better height now!)
    -C1 75.959637, h34, 110 Hz
    +C2 75.970567, h27, 137 Hz
    -C2 75.956002, h36. 132 Hz
    +C4 75.981502, h24, 201 Hz (broad and noisy)
    -C4 75.945069, h32, 195 Hz (also broad and noisy)

    ID7762 rabi-flop on +C1. visually tau_pi ~4ms

    ID7759 peak-search 6 ms probe (Fourier 67 Hz AOM), 300 Hz span, 15 Hz step, 19pts
    +C1 75.966938, h17, 302 Hz (low height, check probe-length with rabi flop)
    -C1 75.959640, h17, 208 Hz
    +C2 75.970573, h22, 124 Hz
    abort and do rabi-flops instead

    Switch to ND2 filter

    ID7758 peak search, 1800 us probe (Fourier limit 222 Hz AOM), 800 Hz span, 40 Hz step, 21 pts
    +C1 75.966949, h48, 202 Hz
    -C1 75.959645, h36, 198 Hz
    +C2 75.970598, h37, 222 Hz
    -C2 75.956008, h51, 233 Hz
    +C4 75.981529, h30, 308 Hz (looks noisy..)
    -C4 75.945104, h28, 279 Hz (better looking peak)

    Switch to ND1 filter

    ID7757 peak-search with 600us probe (Fourier limit 667 Hz AOM), 3000 Hz span, 150 Hz step, 21 pts
    +C1 75.966912, h50, 668 Hz
    -C1 75.959570, h45, 579 Hz
    +C2 75.970544, h40, 707 Hz
    -C2 75.955929, h52, 717 Hz
    +C4 75.981499, h27, 802 Hz
    -C4 75.945003, h50, 601 Hz (much better looking peak than +C4)
    set coil-currents (+1,0,0) mA

    ID7755, run7, carrier, same 16 kHz span

    ID7754 red secular sideband at 75.481 MHz

    ID7753 larger 16 kHz span, around expected blu radial (lower) sideband

    ID7752 spectrum around 76.445, predicted upper secular sideband?
    span is 8 kHz - so we may only see one radial mode?

    scan spectrum at carrier + 0.964 MHz
    Tickler with 10dB attenuation shows radial (lower?) mode at 0.964 MHz
    (on 2020-05-04 we have secular modes at 0.9644 and 1.0067 MHz)

    ID7748 run3, smaller span around carrier. 600us probe, 200 Hz AOM step
    Fit 75.963303, h 64, FWHM 1995 Hz

    ID7745 'run2' spectrum

    abort... correct signs are (-4.6061, +0.7881, +1.3572) mA
    ID7743 'run1', spectrum in zero field!?
    For fun, try a zero-B-field setting of I = (-4.6061, -0.7881, -1.3572) mA

    Notes for possible big box around experiment, 3rd Bshield:
    Trap Breadboard: 900mm x 600mm x 55mm
    C-beam post is highest component on breadboard, height 370mm
    oscilloscope-shelf is ca 660mm from optical table, instrument-shelf is 860mm from table

    06:43Z stop Bmeter logging

    2020-06-14 Sunday

    22:08Z (?) servo seems stalled at iteration 535?? why?
    htop shows 12.5/15.5 G RAM use (Monday morning 07 local)
    gnuradio-companion uses 25% of RAM, and Firefox ca 13% (WebExtensions?)
    diskspace seems ok, 350 G free

    09:45Z leave lab.
    09:43Z start the Bmeter logging on the win-machine. Same probe position as before, ca. vertical



    Quantum jumps



    09:29 run the clock-servo (lock-to-transition v3 program) for C1,C2,C4, with 600us probe
    Not sure how dark/bright threshold now set - find out!?
    This is ID7740 or clockrun3
    Clock-cycle duration is ca 64s

    ID7739 rerun peak search
    +C1 75.966709, 53, 624
    -C1 75.959544, 40, 650
    +C2 75.970336, 41, 688
    -C2 75.955943, 52, 642
    +C4 75.981083, 37, 634 (stronger sidelobes than C1/C2)
    -C4 75,945122, 41, 586

    recompute predicted peak-positions for all 6 peaks - this should prime the peak-search correctly
    FIXME: make this prediction use actual beat-note (cavity-frequency) and AHM3-model

    Try to find all thre pairs C1,C2,C4
    ID7736 peak search for three pairs, 600 us probe, 3kHz span, 150Hz step
    (center / MHz, height / jumps, FWHM / AOM Hz) - from Gaussian fits
    +C1 75.966766, 47, 634
    -C1 75.959584, 48, 614
    +C2 75.970399, 40, 673
    -C2 75.955970, 50, 656
    +C4 (program aborted... length 4 vs length 6 dataset vector problem?)

    ID7733 peak-search, 4 kHz span, 200Hz step
    +C1 center 75.966768, height 58, FWHM -622 (negative!? FIXME)
    -C1 center 75.959613, height 45, FWHM 599
    B1=1.278 uT
    +C2 center 75.970382, h 42, FWHM 760 Hz
    -C2 center 75,955975, h 44, FWHM 779 Hz
    B2=1.286 UT

    set coil currents (+1,0,0) to also see C4 peaks

    ID7731 (restarting an artiq-program works) run peak-search with 660us probe (Fourier limit ca 1300Hz optical)
    +C1 height 52, FWHM 653 Hz
    -C1 height 46, FWHM 635 Hz

    08:34Z Chinesium-camera software had stalled (win machine suggests updates need to be applied..). Restarting cam-software worked. Ion seems still stored. Clocklaser seems locked
    However cavity-camera on website shows a dark image - why?
    ARTIQ seems to have stalled? Last successful run is 7728. Now stalled on 7730 (0-0-5)? why?

    2020-06-12 Friday


    13:27Z: Long B-field run for the week-end. No filters, clock probe 600us, 34 measurements, 80 kHz span, 200 Hz step, approx. 17.5 h total
    Data saved to DB. Currents, RIDs and DB description fields: 2020-06-12-B-field.txt


    06:40Z check uMotions 300kphot: A 1.6% 42s, B 2.3% 51s, C 4.3% 63s

    06:37Z turn off B-field logging and the peak servo
    servo ran for ca 1207 iterations, ca 47 s per C1+C2 pair

    2020-06-11 Thursday

    Peak servo run starting ca 15:00Z

    +/-C1 and +/-C2 and centers in same figure, peaks offset for visual effect:


    Qauntum jumps for the peaks:


    ID7603 restart with PMT-counts set to 90 bright, 10 dark
    ID7602 servo-lock run to C1, C2 peaks (C4 was weak or weird, not used)
    OD2 filter with 4 ms probe

    B-field monitoring start: 14:55Z. Bartington-probe about vertically on IR-window-side of trap
    Meter voltage output seems to be (uT)/100 V, so 0.29 V corresponds to meter-reading of 29 uT
    Sample at 200 S/s and store in text-file on Win machine
    ID7599 Automatic peak search OD2, C1+: 119 Hz, C1-: 98 Hz, C2+: 148 Hz (?), C2-: 135 Hz (?)
    ID7591 line-scan with OD4 and 40 ms, 2 Hz step and 125 Hz range 63 steps, AC off

    ID7590 line-scan with OD4 and 40 ms, 2 Hz step and 125 Hz range 63 steps, AC off
    8.5s/100shots.

    Note: AC-off reverses air-flow, now cold air and dirt comes down from the ceiling...
    ID7588 line-scan with OD3 and 12 ms, 7 Hz step, AC off, FWHM = 81 Hz, two side bands, not a good fit.

    ID7586 line-scan with OD3 and 12 ms, 7 Hz step, FWHM = 72 Hz

    ID7585 flop, out to 30ms - noisy slope upwards..
    Switch to OD3 filter
    ID7584 flop-on center, OD2, 20 steps to 10 ms. tau_pi 4..5ms visually
    ID7582 +C1 line-scan, 5ms probe
    FWHM 100 Hz AOM (Fourier limit 160 Hz optical)

    ID7581 peak-search with 6ms probe, 20 Hz step - takes too long, abort
    Switch to OD2. Try 6 ms probe

    ID7580, A-beam, auto-peak-search, OD1, 2ms probe. 1kHz range 50 Hz step, 21pts/peak
    Gaussian fit FWHM AOM:
    +C1 202 Hz height 48.5, -C1 167 Hz h49.2
    +C2 193 Hz h44.2, -C2 175 Hz h49.5
    +C4 Gaussian fit failed - program abort.
    (+C4 spectrum noisy with no clear peak?)
    at new voltages uMotion 300kphot: A 1.9% 42s, B 2.2% 43s, C 4.5% 54s - A and B improved, C same or slightly worse
    Set the new optimal voltages (ID7566 result): V1 = -5.7788, V2 = -5.2212, V3 = -16.7212, V4 = 72.8024
    uMotion at old voltages, 300kphot: A 6.2% 42s, B 3.4% 48s, C 4.1% 54s
    (422-detuning kept at 208 MHz throughout loading and experiments)

    11:01Z flashing the 405 on/off a few times restored ion fluorescence!? (one of the first times this has worked!?)
    check 422 Rb-lock, seems fine
    10:55Z Ion seems to have vanished..

    NPL results: Optimal voltages (Common mode, diff): [ 0.2787832 28.04061488 44.76179374]
    (old voltages in use so far: [0.2763 27.9914 44.5533] )
    09:00Z over lunch: ID7566 NPL uMotion-optimizer with old scan ranges 4x4x4, 200kphot per beam
    B-beam was realigned to get some signal again

    ID7561 repeat auto-peak-search, now with fiber-noise and breadboard-intensity servos on
    Fit AOM FWHMs (2ms Fourier limit 400 Hz optical). 50 Hz AOM-sampling is only 4pts/FWHM
    +C1 183 Hz, -C1 189 Hz
    +C2 209 Hz, -C2 222 Hz
    +C4 720 Hz, -C4 1996 Hz (??)
    (-C2 far right in scan-range, because overnight cavity-drift not accounted for in the starting-values?)
    (+C4 double-peak or very strong rabi-sidelobe? low max Pexc 22/10 )
    (-C4 no clear peak found)
    peak-search run duration 690 s

    ID7560 try auto-peak-search (old starting values from yesterday). scan range is 1 kHz AOM around old peak value. 50Hz step, 2ms probe
    +C1 found to the left of scan.
    Gaussian fit FWHMs:
    +C1 198 Hz, -C1 (fiber-noise and trap-breadboard-intenstiy servos switched on mid-way...)
    abort and re-run

    fbeat 23164431 Hz
    Clocklaser. Toptica 170mA, piezo 58.638 V, amplifier 350 mA

    Fluorescence levels at 90 bright and 10 dark in 10ms seem still valid
    08:20Z stop collision monitor
    08:19Z ion still stored

    2020-06-10 Wednesday

    15:44Z Turn off clocklaser + comb amp(s)

    meh, no clock-run today, just leave collision monitor.
    15:36Z start collision monitor. 10ms gate, 90 bright, 10 dark photons. Only A-beam on. No clearout.
    states
    dark_durations
    bright_durations


    ID7556 peak-search again after re-locking clocklaser. Leave intra-cavity servo off, it may cause unlocking?
    AOM FWHMs from fits:
    +C1 211 Hz, -C1 179 Hz
    +C2 168 Hz, -C2 205 Hz - again clock laser in TEM01!!??? why?
    +C4 , -C4

    ID7555 try the automagic peak-search program with 2ms probe
    AOM FWHM outputs:
    +C1 192 Hz, -C1 202 Hz
    +C2 184 Hz, -C2. Clock-laser jumped to TEM01 mode!!?? argh. abort run
    +C4, -C4

    ID7552 run11, line-scan, OD1, 2ms, 40Hz step 1500 Hz range
    Fourier limit 400Hz optical, observe FWHM 186 Hz AOM (Gaussian fit)
    2020-06-10-run10.txt

    ID7551 try OD1, flop shows tau_pi ca 2ms

    ID7550 flop with comb-amp OFF
    ID7549 flop-on-center
    ID7546, run10, 4ms probe, OD2 filter. Wider 500Hz scan with 20 Hz sampling
    FWHM 187 Hz AOM - but Fourier limit is 0.8/4ms=200Hz??
    2020-06-10-run10.txt

    Try OD2 and 4ms

    ID7545 run9, OD3 with 13ms probe, same scan as before
    Noisy peak maximum with Pexc=26/100
    Expected FWHM is (0.8/13ms)/(2*2Hz) = 15 points wide
    61 points in 356 s
    2020-06-10-run9.txt

    try OD3, if 40ms<->OD4, then OD3 should have 40/3=13 ms probe with 61 Hz Fourier limit.
    ID7544, run8, repeat with 120 Hz AOM range, to also see +/-100Hz optical sideband? Re-center on previous scan presumed carrier 75.965777
    61 points in 521 s
    2020-06-10-run8.txt


    ID7543 line-scan with 80 Hz AOM range, 2 Hz step, run7
    Now see 50Hz side-band peaks!? They are spearated by 50Hz/(2*2Hz) = 12.5 steps(datapoints) from the carrier
    2020-06-10-run7.txt

    Try intra-cavity power servo. set-point 1.25V (1348nm Iamp=350mA)
    (fiber-stabilization, and probe-power-at-beamcombiner are already ON, from the start of the day)

    Turn comb-amp back on
    ID7542,run6, same line-scan 120 Hz AOM range, try switching OFF the Comb 1348nm beat-note branch amplifier
    61 frequency points, duration 521s
    2020-06-10-run6.txt


    repeats, 2020-06-10-run1,2,3,4,5
    OD4, 40ms probe, 2Hz AOM step, only 80 Hz AOM range (should see +/-50 Hz optical sidebands if they are present)
    Fourier limit is 0.8/40ms = 20 Hz optical, 10 Hz AOM, so with 2 Hz AOM sampling we should get 5 points/FWHM.
    Scanning 41 frequency points, 100shots/point, takes roughly 350 seconds, so 8.5 s/100shots
    data: 2020-06-10-run1.txt
    2020-06-10-run2.txt
    2020-06-10-run3.txt
    2020-06-10-run4.txt
    2020-06-10-run5.txt






    ID7536 line-scan with 40ms probe
    ID7535 rabi-flop on center, 20 steps to 100ms. oscillating rising curve to 30% @ 60ms, then smoother??
    ID7534 repeat line-scan with 2Hz step 55ms probe. Two peaks at Pexc=30/100 spaced only 12 Hz apart?
    ion had a dark moment... ID7530 aborted
    ID7530 line-scan with 2Hz step, 55ms probe
    ID7529 rabi-flop on center - low Pexc and strange shape
    ID7528 first OD4 line-scan, +C1
    Try OD3+OD1 = OD4 with a long probe

    ID7525 repeat peak-search with OD2, 4ms probe, now 20 Hz AOM step. AOM FWHMs:
    +C1 127 Hz, -C1 110 Hz
    +C2 149 Hz, -C2 129 Hz
    +C4 248 Hz, -C4 273 Hz, (noisy looking spectra..)
    Note C4 Zeeman shift is 5-fold more than for C1.

    ID7524: Automatic peak search. OD2, 4ms probe. FWHMs: +C1: 115 Hz, -C1: 129 Hz, +C2: 154 Hz, -C2: 125 Hz, +C4: 241 Hz, -C4: 325 Hz

    ID7520: +C1 line scan, OD2 ND filter, clock pulse 4000 us. AOM FWHM = 102 Hz. Fourier limit = 200 Hz.
    ID7517: +C1 Rabi OD2 ND filter. tau_pi ~ 4000 us

    ID7516: +C1 line scan, OD1 ND filter, clock pulse 1900 us. AOM FWHM = 228 Hz. Fourier limit = 421 Hz.
    ID7513: +C1 Rabi OD1 ND filter. tau_pi ~1900 us

    ID7511: +C1 line scan, no ND filter clock pulse 550 us. AOM FWHM = 700 Hz. Fourier limit = 1454 Hz.
    ID7510: +C1 Rabi no ND filter. tau_pi ~550 us
    Problems with B-beam alignment: almost no signal
    uMotion contrasts 300kphot: A 6.5%

    461 parameters check: 197.2 mA, 4.645 kohm

    11:03Z Ion at 8min!
    10:55Z Load with I=1.55A (1.35 W), 422-detuning at 208 MHz (because this worked OK yesterday)
    06:01Z collision monitor off.
    Ion seems lost, possibly due to Rb-lock problems (lab temperature induced?) with coolinglaser

    2020-06-09 Tuesday

    14:06Z collision monitor on. 10ms gate, 90 bright 3 dark counts.
    states
    dark_durations
    bright_durations

    ID7426 repeat beam shutter on trap bread board again. line-scan repeat with comb-amp on. Both ACs off. now FWHM 109 Hz AOM
    ID7424 beam shutter on trap bread board again. line-scan repeat with comb-amp on. Both ACs off. now FWHM 117 Hz AOM

    ID7422 repeat line-scan repeat with comb-amp on. Both ACs off. now FWHM 104 Hz AOM
    ID7420 line-scan repeat with comb-amp on. Both ACs off. now FWHM 108 Hz AOM
    ID7419 rabi-flop on center. Visually tau_pi ~5 ms.
    ID7417 line-scan repeat with comb-amp on. now FWHM 89 Hz AOM
    ID7415 line-scan repeat with comb-amp off. now FWHM 130 Hz AOM
    ID7414 line-scan 5 ms probe. FWHM 96 Hz AOM - Fourier limit is 160 Hz optical.
    ID7412 try OD2 filter

    ID7411 rabi-flop on center. Visually tau_pi ~2ms.
    ID7408 line-scan of +C1, 1900us probe. FWHM 205Hz AOM - Fourier limit is 421 Hz optical.
    Try 422-shutter further away from the ion. Spare shutter now mounted on optical table at 422-breadboard fiber-launch

    ID7399, ID7400-7401: +C5, clock probe 1900us. FWHM: 258 Hz AOM.
    ID7396, ID7397 (7403-7404 high res.): +C4, clock probe 1900us. Two peaks ?.
    ID7392: +C3, clock probe 1900us. FWHM: 250 Hz AOM.
    ID7390: +C2, clock probe 1900us. FWHM: 220 Hz AOM.
    ID7387: +C1, clock probe 1900us. FWHM: 210 Hz AOM.
    ID7386 OD1 wide scan 70 kHz, clock probe 1900 us. 2020-06-09-run14

    ID7383 OD2 line-scan 4500us, repeat, Gaussian fit 116 Hz FWHM AOM - Fourier limit is 178 Hz optical. Both ACs off. 461 drivers off. 2020-06-09-run10.
    ID7382 OD2 line-scan 4500us, 2 peaks ?!, Both ACs off. 461 drivers off. Comb amp off. 2020-06-09-run10.
    ID7380 OD2 line-scan 4500us, Gaussian fit 124 Hz FWHM AOM - Fourier limit is 178 Hz optical. Both ACs off. 461 drivers off. 2020-06-09-run10.
    ID7378 OD2 line-scan 4500us, Gaussian fit 140 Hz FWHM AOM - Fourier limit is 178 Hz optical. Both ACs off. 2020-06-09-run10.
    ID7377 OD2 rabi-flop on line-center - AC off -3k (KSK?), AC off behind the wall -2k (TK203).
    ID7376 OD2 rabi-flop on line-center - AC off

    ID7375 OD1 rabi-flop on line-center - AC off
    ID7373 OD1 line-scan 1900us, Gaussian fit 205 Hz FWHM AOM - Fourier limit is 420 Hz optical. AC off. 2020-06-09-run9.

    ID7372 OD1 rabi-flop on line-center -
    ID7369 OD1 line-scan 1900us, Gaussian fit 240 Hz FWHM AOM - Fourier limit is 420 Hz optical
    OD1 ND-flip-filter gives 30 nW
    674nm 270nW -> tau_pi = 600 us. FWHM 661 Hz - agrees with Fourier limit

    uMotion contrasts 300kphot: A 5.1%, 0.8% B %, 3.4C %

    run1: scan for +C1 peak, 1ms probe
    07:12Z stop collision monitor

    2020-06-08 Monday

    14:05Z collision monitor on. 10ms gate, 90 bright 3 dark counts.
    states
    dark_durations
    bright_durations


    Dark state monitoring: DB id: 2020-06-08-run1.

    ID:7334 Line scan, +C1, clock 10 ms. OD2 ND filter. Comb amplifier off. FWHM: 110 Hz.
    ID:7333 Rabi at +C1. OD2 ND filter. Comb amplifiers off.
    ID:7332 Rabi at +C1. OD2 ND filter.
    ID:7330 Line scan, +C1, clock 10 ms. OD2 ND filter. FWHM: 130 Hz.

    ID:7328 Line scan, +C1, clock 2.7 ms. Center: 75.650 MHz, FWHM: 160 Hz (AOM). OD1 ND filter.
    ID:7327 Rabi at +C1. OD1 ND filter.
    ID:7325 Line scan, +C1, clock 3 ms. Center: 75.650 MHz, FWHM: 154 Hz (AOM). OD1 ND filter.

    ID:7322 Rabi at +C1. No ND filter.
    ID:7320 Line scan, +C1, clock 1 ms. Center: 75.650 MHz, FWHM: 373 Hz (AOM). No ND filter.

    Line scan, clock pulse 1 ms, 2020-06-08-run2.

    uMotion contrasts 300kphot: A 6.5%, B 3.9%, C 4.2%

    12:05:15Z try to load again. 1.55A
    12:08:20Z signal! single ion

    11:51Z Idisp=1.55A (1.3W) - try to load again
    12:01Z signal! two ions.
    Glitch from signal-gen at 36mV - ions lost

    RF-monitor 1.69 Vpp - 70mVpp from signal-gen

    11:36Z Idisp=1.55A (1.3W)
    No signal in 10 mins

    461nm check dispenser - 198.07mA - a bit lower than before (199mA)?

    Check pulse-stretcher. Output is 25ns long

    2020-05-27 Wednesday


    -2 floor air-compressor seems to give strong 85 Hz noise component
    07:19Z reduce KSK105 VFD setting to 20 Hz

    Cavity heating power slightly higher than before 20.5 W + 3.2 W (heat-loss thru steel I-bar to the now colder A/C channel?)

    A/C VFD was on 25 Hz setting overnight. lab-temp shows slow oscillations around +23C

    07:09Z ISS to further reduce input-air temp. Biltema sub-floor shows 18.1C

    2020-05-26 Tuesday

    11:17Z Biltema temp sensor: 18.2C 08:25Z lower A/C airflow to 25 Hz setting
    08:20Z lab-temp falls below +23C, biltema-meter sub-floor +18.5C.
    Biltema-meter at back of lab shows 19.2C under the floor
    06:50Z ISS to lower temperature of incoming air, setting at +17.5C
    A/C 30Hz and original temp gave ca 23.5C lab-temperature (Vaisala DMP248 sensor)
    Now 5pcs D160mm holes in the floor completely open

    2020-05-25

    10:35Z Try a setting of 30 Hz
    Various tests with different A/C-VFD frequencies..
    Opened 4pcs D160mm holes in the floor. This lowers lab-temp by ca -0.5C with A/C-VFD at original RPM of 48 Hz

    2020-05-22

    With EOM PA outside the box the top coarse heater servo now works
    Ca 19W bottom heater and 2W top heater now (rightmost data in figure)

    2020-05-20 Wednesday


    17:19Z Turning off the experiment... ion storage time was 7+ days
    17:12Z A/C VFD back to 48 Hz


    ID7292 flop to 20ms, 100shots - compare with previous 7291, but this one is longer. seems noisy after 8ms or so
    ID7291 flop on center. 100shots, good oscillations out to 10ms
    ID7289 line scan 1.4ms probe
    fit FWHM 310 Hz AOM


    ID7288 flop with high probe power, first max ca 1400us
    for fun try a high probe power, ca 180 nW now

    not so great flop
    Adjust probe power with waveplate.
    previously 18dB att was ca 25 nW.
    now go to lower value, with 18dB att ca 5 nW

    ID7284 line scan 20dB att, 4ms probe
    FWHM 106 Hz AOM

    ID7283 flop 20 dB att, 100shots - max around 4ms
    ID7282 flop 24 dB att, 100shots - rises to 6ms but then incoherent/noisy?

    ID7281 flop-on-center 18dB att, 100shots - seems ok - tau_pi now maybe shorter than 4ms? compare to ID7276 with high 422-power
    ID7279 scan for line, 4ms probe, 18dB attenuation 20Hz step

    mis-align 422 input-fiber so we have 0.4 uW and get 100 counts in 10ms
    422-amp to 1.8A, ca 0.8 uW in A-beam, 160 counts in 10ms (previously 'low' cooling was 80 counts?)
    Clock-laser jumped to TEM01 mode? but maybe flop-run was done?
    We now have a 'high' cooling power 1.24uW, so next ty a lower value?
    ID7276 flop, 18dB att, 500 shots
    ID7274 flop, 26 dB AOM att - slow incoherent rise
    Note: comb beat-note stdev now at 23 Hz (lower than before?)
    ID7272 line-scan 8ms probe, 23dB AOM att, 10 Hz step.
    gaussian not a good fit

    ID7270 flop, 23 dB AOM att. tau_pi maybe 8ms
    ID7268 line-scan, 20dB AOM att, 6ms probe, intra-cavity servo on
    fit FWHM 87 Hz AOM

    ID7267 flop, 20dB AOM att, intra-cavity servo on (1.4V setpoint), tau_pi ca 6ms
    ID7266 flop, 20 dB AOM att
    ID7264 line-scan 6ms probe, 10 Hz AOM step
    Fit FWHM 96 Hz AOM

    ID7262 line-scan, 6ms probe, 20 Hz AOM step
    FWHM 103 Hz AOM

    ID7261 flop-on-center, tau_pi maybe 6ms now(!)
    ID7260 line-scan 4ms probe

    15:30Z turn down aircon VFD to 10 Hz (it was at 48..49 Hz)

    ID7252 line-scan with 2ms probe
    seems noisy, Pexc(max)=16/100. Fit FWHM 245 Hz AOM


    ID7250 repeat line-scan
    Fit FWHM 167 Hz AOM


    ID7248 line-scan, 18dB AOM att, 4ms probe. 180 counts bright-state (we have more 422-power than before)
    better blanket, but still a noisy spectrum!?
    Fit FWHM 186 Hz AOM

    ID7246 flop on center, 18dB AOM att
    ID7244 flop on center. 20dB AOM att
    ID7242 try positioning a blanket over the trap breadboard
    4ms probe, 18 dB att, 20 Hz step
    fit FWHM 159 Hz AOM (but Gaussian not a good fit)


    ID7240 line-scan without blanket. 18dB probe att, 4ms probe. 20 Hz AOM step. (this is not necessarily a pi-pulse - test rabi-osc also later)
    Gaussian fit gives FWHM 178 Hz AOM (Fourier limit 200 Hz optical) - so we have 1.78-fold FWHM

    Try a bigger holy blanket that minimizes air-currents around the trap breadboard?

    NPL-optimization 200kp, ID 7231.
    Optimal voltages (Common mode, diff): [ 0.36018227 28.15357435 44.74236194]
    Optimum outside scanning range!

    uMotion contrasts at the old bias voltages. 300 kphot
    Old: 0.2763, 27.9914, 44.5533
    A 7.2% (26s), B 7.6% (40.4s), C 6.5% (39s)

    11:10Z doors closed
    09:45Z doors open. Move the EOM PA outside the box.
    09:32Z collision monitor off.

    2020-05-19 Tuesday

    14:15Z collision monitor on. 10ms gate, 80 bright 2 dark counts.
    states
    dark_durations
    bright_durations


    ID7208 line-scan, 20 dB AOM att, 4 ms probe
    FWHM 171 Hz AOM. noisy?

    ID7207 flop, 20 dB AOM att, visually tau_pi 4ms
    build a finnfoam box for the ECDL

    ID7197-98 repeat, clock laser now repeatedly unlocks??
    Gaussian fit 140 Hz FWHM AOM (This was 79 Hz before lunch)

    ID7196 repeat 6ms scan, with 10 Hz step - clock laser unlocked
    ID7193 repeat 6ms line-scan, to compare aircon off vs on
    6ms probe, 23dB AOM att. 20 Hz AOM step
    Fit FWHM 56 Hz AOM


    Go back to the old voltages
    New uMotions 300kphot: A 14.6%, B 12%, C 8.4%. These are worse than before!?
    set the new voltages
    uMotion contrasts at the old bias voltages. 300 kphot
    Old: 0.2763, 27.9914, 44.5533
    A 10% (35s), B 9.6% (39s), C 7.4% (65s)

    Vopt found: (Common mode, diff): [ 0.31609226 28.105798 44.69042412]
    ID7174 NPL uMotion over lunch. 4x4x4 grid. C-beam gives quite weak fluorescence compared to A and B

    ca 12 local - aircon back ON

    ID7172 line-scan 12ms probe, 24 dB attenuation
    Gaussian fit FWHM 84 Hz AOM (not a good fit)

    ID7171 flop on center, 24 dB AOM att, tau_pi 12ms visually

    ID7170 line-scan 14ms probe 25 dB AOM att.

    ID7169 line-scan, 14ms probe, 27 dB AOM att.

    ID7168 try 27 dB AOM att. ID7167 flop on center, 29 dB AOM attenuation. taus out to 40 ms - no clear osc.

    ID7165 relock clocklaser, try again. same as ID7164, 14 ms probe 26 dB AOM att.
    FWHM 60 Hz AOM, but Gaussian is not a good fit


    ID7164 repeat same scan - clock laser unlock!
    ID7161 same scan, but smaller span
    Fit gives FWHM 59 Hz AOM


    ID7160 line scan, 14 ms probe, 26 dB AOM attenuation. 6 Hz AOM step
    Fourier limit now 57 Hz optical
    meh.. abort the long run

    ID7159 flop-on-center, now 26 dB AOM attenuation.


    ID7157 line-scan 6 ms probe, 23 dB AOM att, 20 Hz AOM step
    Fourier limit now 133 Hz optical
    Fit gives FWHM 79 Hz AOM


    ID7156 flop. increase to 23 dB AOM att. taus out to 20ms now. now tau_pi maybe 6ms
    ID7155 flop-on-center, seems ok tau_pi around 4ms
    ID7153 line-scan again, 4000us probe, 20 dB AOM att.
    FWHM 102 Hz AOM, close to the Fourier limit!


    08:12Z aircon OFF

    ID7151 line-scan 4000us probe, 20 dB AOM att.
    FWHM 178 Hz AOM


    ID7150 flop on center, try 20 dB AOM att. visually tau_pi 4ms
    ID7148 line-scan again. 3200us probe
    FWHM 175 Hz AOM


    ID7144: Line scan with 461 current, temp and dispenser powers off (acoustic noise/vibration?).

    FWHM AOM 154 Hz
    ID 7146: Rabi

    ID7143 Rabi at line center.
    ID7140 Line scan with doors closed, same settings as the day before.
    FWHM 178 Hz AOM (similar to last line-scan y-day!?)


    With the AA AOM PAs (ca 9W each) moved off the cavity-system and attached to the Thorlabs rails we still overshoot the cavity-top temperature
    Previously bottom-heater was at only 5-6 W, now at 14 W.
    Probably need to move the AOM PAs and the EOM PA outside the box?
    Top temperature overshoot is now +0.150 K. Previously it was around +0.8 K
    Positive Kikusui supply to the box is +24V 2.7A = 65 W


    07:23Z: Collision monitor stopped

    2020-05-18 Monday

    15:36Z restart monitor with 10ms gate time, 55/2 bright/dark counts
    states
    dark_durations
    bright_durations


    14:51Z start collision monitor
    Previous runs had all beams ABC. Now we have only A beam. Bright 18, Dark 1 count in 2ms
    states
    dark_durations
    bright_durations


    line-scan with doors closed, 3200us probe
    FWHM 178 Hz AOM

    14:34Z close cavity doors
    Move AOM PAs from top-breadboard to the inside of the Thorlabs-profile. Still inside the box, but not in contact with the cavity-system


    line-scan with 3200us probe, 25 Hz AOM step
    FWHM 220Hz AOM


    intra-cavity servo on seems to unlock the PDH-lock to the cavity (jumps to a TEM01 mode) - maybe lower gain, BW, or output-range on intra-cavity servo would help?
    line-scan again
    touching the cavity probably shifted frequency
    Try intra-cavity servo on. set-point 1.7V (low?)
    ID7122 flop on center, AOM att 20 dB (was 18dB previously)

    ID7120: line scan
    Fit 75.970640, FWHM 240 Hz AOM

    ID7119: Rabi
    turn on intensity servo

    line scan with 40nW, 2600 us (fourier limit 308 Hz optical). 30 Hz AOM sampling
    Fit: 75.970705, FWHM 221 Hz AOM


    ID7116 flop, 500 shots, 20pts out to 10 ms. visually tau_pi 2600 us
    ID7115 flop-on-center, 40nW. taus 0.05 to 5 ms
    lower probe to 40 nW, with waveplate

    ID7113 rabi-flop on center. ca 300 nW, visually tau_pi= 1300 us
    +C1 at ca 75.9708 MHz
    Fit 75.970787, FWHM 437 Hz AOM


    Fiber-noise servo on, but intensity-servo and intra-cavity servos off for now
    Beam powers: 674nm: 320 nW. 1ms probe, scan for peaks


    12:53Z: Dark state data collection stopped

    replaced detector with ET-2030A (from length lab. Now with two mini-circ amplifiers

    Photodiode-signal after amplification ca -15 dBm
    Try to optimize 357 MHz fiber-noise detection chain.
    Currently: unknown photodiode -> Filter SHP-300 HP 290MHz -> ZHL500LN +28dB -> SLP-450 -> ZHL500LN +28dB -> 6 dB Att -> ZHL550LN +28 dB
    +84dB - 6dB = 78 dB Gain!?

    07:08Z open cavity doors, to decrease heat-load (AOM PAs) inside box

    07:04Z ion still stored. collision-monitor running. Now back-of-lab lights on, hopefully doesn't affect PMT counts

    2020-05-17 Sunday

    14:20Z no dark states seen so far (12 minutes). Leave this running.
    14:08Z start new collision-monitor run, now with also Clearout ON. We expect to only see short events now!?
    This is 2020-05-17-run3
    states
    dark_durations
    bright_durations


    14:07Z end the random telegraph, ca 2k events collected. Turn OFF 405-laser.
    Try driving D5/2 state with 405nm laser. Visually turn-down 405-power to only 30uW to get nice quantum-jumps. A high 405-power results in completely dark ion (at 2ms resolution)
    'random telegraph' method
    2ms PMT gate. 28 bright, 1 dark counts. Threshold is geometric mean, rounded to an integer, so 5.
    Letchumanan thesis FIG5.6 has 97k events and lifetime of 389.3(2.1) ms
    We collect ca 60 events per minute, so 97k events in 27 hours
    dark_durations

    200ms get 2800/20, or SNR 140 (21.5dB). Scaling this to 2 ms gives 28/0.2 !
    20ms get 280/2
    2ms gate, 28/1 bright/dark counts still roughly OK.

    13:33Z stop collision-monitor run. It ran from around 15:23Z on Saturday until 13:33Z on Sunday, and recorded ca 110 events. Ca 70 of these (64%) are short-lived around 2-6 ms maybe.
    13:31Z ion still stored and collision-monitor running

    2020-05-16 Saturday

    Try to plot Likforman2016a FIG4 figure:
    states
    dark_durations
    bright_durations

    15:23Z vacuum collision-monitoring restarted with 2ms gate (bright 28 counts, dark 1 count), tag 2020-05-16-run4

    NOTE: influxdb needed re-starting (systemctl restart influxdb) after disk-space was OK
    Dispace was running low, about 360 GB SDR IQ data... deleted.
    ion still stored

    2020-05-14


    C1 not resolved, problem with fiber noise cancellation. Rerun:
    wide scan. center 75.963 MHz, span 2.2 MHz, step 200Hz. 11 kpoints, duration 58 ks = 16 hours



    Overnight run failed due to memory leak. Problem solved, rerun:
    wide scan. center 75.963 MHz, span 2.2 MHz, step 200Hz. 11 kpoints, duration 58 ks = 16 hours


    2020-05-13


    wide scan overnight. center 75.963 MHz, span 2.2 MHz, step 200Hz. 11 kpoints, duration 58 ks = 16 hours


    fbeat 23164789 Hz. predicted line-center 75,963 MHz
    Back to high probe-power of 370 nW. tau_pi=800us. probe-intensity servo now off
    FWHM 668 Hz AOM, Fourier limit 1kHz optical

    and now a 60s image-stack:
    Determining the Zernike polynomials that lead to these aberrations is left as an exercise for the reader


    ID6999 line-scan
    12:55 Center 75.966346 MHz, FWHM 361 Hz AOM (if line-center now has moved a lot, then we cannot compare ID6998 and ID6995 fairly!)
    ID6998 and ID6995 were run at 75.966328 MHz (line-center measured ca 12:10)


    12:39 ID6998 is comparison against ID6995, now with aircon recently started
    12:35Z aircon back ON
    12:32 ID6997 AOM-att 18dB, out to 20ms
    12:30 ID6996 AOM-att 24 dB (probe W?), 100shots, incoherent..
    12:20 ID6995 repeat with 500 shots. Pexc(tau_pi=2.5ms)=163/500 = 33%
    12:17 ID6994 flop on center - clear oscillation, tau_pi 2..3 ms (similar to 2020-05-08 result with 63 nW?)
    12:10Z ID6992 line-scan 3ms 30nW, 25Hz AOM step
    clock-laser unlocked (door-slamming...)
    FIT 75.966328, FWHM 249 Hz AOM

    ID6991 aircon OFF flop!?!
    12:06Z aircon OFF
    ID6990 flop on center. incoherent rise to 25%
    check probe power, still ca 30 nW
    try with airconditioning off!?

    ID6989 flop on center, reduced 422-power

    get ca 80 counts/10ms with PMT
    Reduce 422-power to ca 0.3 uW

    ID6989 line-scan again, 3ms 30nW probe
    low incoherent Pexc level?
    fit 75,966275, FWHM 379 Hz AOM


    ID6985 flop on center, 30pts to 10ms, 200shots - no clear oscillations
    ID6983 line-scan, 30nW 3ms probe
    FIT 75.966267 MHz, FWHM 271 Hz AOM


    ID6982 flop, to 10ms, 100 shots
    reduce to 30 nW, with the waveplate

    probe was ca 300 nW
    ID6980 rabi-flop on center, to 5ms, 100shots. tau_pi ca 1 ms visually
    scan for +C1 around 75.966
    Fit: 75,966291 FWHM 493Hz AOM


    scan for clock transition
    1ms probe. fiber-noise servo ON, probe-intensity servo ON, intra-cavity servo OFF
    with beat-note of 23.1647 MHz line-center should be around 75.963 MHz


    at Vopt uMotion 100kphot: A 7.4%, B 6.9%, C 6.0% - not much better than before?
    Optimization result is: Optimal voltages (Common mode, diff): [ 0.27629906 27.9913947 44.553304 ] 09:05Z 39/64 pts done. ion still alive and stable..
    ID6965: Try the NPL uMotion optimization. 4x4x4 grid.

    uMotion 100kphot: A 6.9%, B 7.3%, C 6.5% (weakest)
    detuning to 208 MHz

    optimize A-beam
    08:20Z now suddenly a stable ion image!!?
    stacked 10s image:

    trapping too unstable to optimize anything?? image comes and goes randomly, just few seconds bright, then dark for 10s or 1 min or whatever..
    strange blinking. try relocking to rb-line
    07:30Z signal! diffuse,single ion? 07:24Z dispenser 1.55A (1.3W)
    try to reload again
    detuning to 185 MHz

    Try aligning beams onto two ions, with RF 100mVpp
    turning off 422 for 30s or 60s does not remove second ion
    changing V34cm around moves ion sideways in image, but doesn't get rid of the second ion. ca 10V steps tried. But we have slow RC-filters on the bias voltages..
    turning off cooling for 10s or 30s does not remove the second ion
    RF-drive can be adjusted 40mVpp to 100mVpp without glitches. At 40mVpp ions move clearly further apart.
    stacked image, 40x 1s exposure:

    07:00Z signal! again two ions? now stable side-by-side for longer
    06:54Z dispenser 1.55A (1.3W) (lower is better?)
    cycle RF OFF/ON, and try to reload


    diffuse blinking image? briefly saw two sharp images side-by-side
    06:44:20 ion! diffuse image? went dark, then reappears
    06:37:30Z dispenser 1.6A (1.40W)

    422-detuning 185 MHz
    all beams on
    camera: 1s exposure, high gain, bin2
    405-power to 2.5mW
    tickler-cable terminated 50R
    Trap RF 16.77 MHz, 70 mVpp, monitor 1.69 Vpp

    beat-note around 23.1644 MHz now, maybe 1 kHz change from before closing the doors

    Bottom heater power dropped from 40-45 W to ca 6 W. Time-constant for coarse heaters around 4-6 hours maybe

    However cavity-box top is still too hot by +0.8 C.
    The AOM-amplifiers, 2pcs bolted to side-breadboards, consume about 9W each. Could be moved outside cavity-box?
    EOM-amplifier is on stone, so heats air inside box.
    Photodetectors, camera, ion-pump should dissipate only small amounts of heat?

    2020-05-12


    14:47Z shut down coolinglaser. Leave clocklaser, tracker, and comb running
    Overnight might see large cavity-frequency excursions as temperature settles?

    go back to the old Vbias values (-5.5, 0.2772, 27.9984, 44.4698)
    Try 405+461, but doesn't seem to help - ion gone/dark?
    Try to re-lock Rb-line. All beams on.
    Program only got to point 8/64, then ion went dark.
    Try to run uMotion-optimization in a 4x4x4 grid around previous optimum. 100 kphot.
    ID6876. 64-point scan. (A 11s + B 17s + C 19s) * 64 pts= 3000s duration

    run3: try again scanning +C1, 2 kHz region around peak in run1
    get Pexc of 5/100 as max. Maybe all Pexc is now in uMotion sidebands?

    uMotion 100kphot: A 66% !?, B 69%, C 68%
    These have changed a lot!

    do we have multiple ions? blink-test looks ok?
    went dark when starting clock-spectrum scan!? why?
    detuning to 208 MHz

    13:34:40Z ion! looks good!?
    13:29Z dispenser 1.6A (1.4W...1.418W)
    405-power to 3 mW...drifts to 2.5mW, try 1.6A. camera 2s exposure

    13:28Z no clear signal.. 13:16Z dispenser 1.55A. all beams on. 1.315W
    Try to reload again, with 2.6 mW 405-power, which worked previously

    12:56Z dispenser 1.55A, 1.306 W
    13:09Z signal, but flashing? (possibly multiple ions? long loading time and high 405-power?)
    reload ion. detuning back to 185 MHz

    scan for clock transition around 75.961, 1ms probe
    set probe to max 470 nW
    437 frequency points at 5.5s per point
    ion seems lost sometime after half of this scan!? (12:45Z maybe)
    +C1 at 75,96516 MHz


    detuning to 208 MHz for operation
    11:42:30Z dispenser 1.55A. (1.32W)
    ion! at 11:49Z

    detuning to 185 MHz for loading. Try max 405nm power of ~6 mW.

    RF siggen 70mVpp, monitor 1.7Vpp
    Try to load, with new camera. 1s exposure, high-gain, 2x2-binning, auto-histogram

    11:18Z close doors to cavity. Expect heater power to drop from 71-72W down. Previously drop was from 90W to maybe 30W.
    New wood-fiber floor-tile with hole for input fiber. Air-gaps to cavity-space sealed.

    2020-05-08

    shut down lasers, leave trap RF on, and comb oscillator on.



    ca 11:03Z airconditioning turned back on
    Clock-laser unlocked! beep from IR amplifier (?) won't turn on again (cable problem?)...
    ID6831 scan for peak again, 5ms 10nW, 'run1', get low Pexc, fit (Fourier limit 160 Hz)
    ID6830 scan for peak, 5ms 10nW,
    ID6829 flop on center P674=10nW, reduce sampling to 20pts, expect sqrt(60/10)=2.4-fold tau_pi at 5.9ms(?). Pexc very low - probably not at line-center?
    ion went dark for a while, after adjusting probe power - why?? re-lock 422 to Rb-line, just in case...
    reduce probe to 10 nW
    ID6823 flop on center. visually tau_pi ~2.4ms. P674 ca 63 nW
    This is possibly the slowest rabi-osc with good cotrast we have seen.

    ID6821 scan to find the line again. 3ms probe. Gaussian fit center 75.964460, FWHM 142 Hz AOM
    Rabi model FWHM 305 Hz

    10:16Z airconditioning turned off

    With airconditioning on, no contrast at 50nW probe visible


    ID6820 P422=200nW, P674=50nW. Rises to Pexc 25% at ca 2.5 ms
    Reduce P422 further? to 200 nW. PMT counts now 50/10ms
    ID6818 500shots, P674 50nW, P422 300nW. Only see 25% Pexc at expected tau_pi=3ms. Flat 25% after 3 ms
    Reduce probe further, to 50 nW, expect sqrt(3)~1.7-fold longer tau_pi (maybe 1.7*1.8 ms = 3 ms)
    ID6817 500shots, probe 150nW, P422 0.3uW. Probe reduced 2.4-fold, expect tau_pi increase 1.55-fold. visually tau_pi ~1.8 ms now
    reduce probe to 150 nW
    ID6816 repeat with 500 shots, P422 0.3 uW, probe 360 nW, visually tau_pi ~1 ms
    ID6815 same as ID6812 but with lower 422-intensity
    Try lowering 422-intensity to ca 0.3uW, PMT counts down to 80/10ms
    Currently 422nm ca 0.6 uW, PMT counts 150/10ms
    ID6812 rabi-flop on center freq. tau_pi=1ms, Pmax 50%, oscillation to about 5ms, then 25%
    flop with probe 0.05 - 10 ms in 40 steps
    ID6810 360nW, 1ms probe scan -> center 75.964383, FWHM 424 Hz AOM
    adjust probe to 360 nW - using waveplate for adjustments (Zeeman AOM DDS attenuation constant at 18 dB)
    ID6809 scan for peak, 1ms probe 3kHz range (this was at 40nW? probe, very low Pexc)
    422-detuning is 208 MHz ( 2*(208-218) = -20 MHz optical)

    2020-05-07 674-intensity-servo



    ID6807 should be repeat of ID6802, slow rise to incoherent 25% Pexc
    run9 +C1 peak with 3ms probe 45nW
    FWHM 238 Hz AOM (Fourier limit 267 Hz optical)

    ID6802 rabi-flop, 45nw, to 10ms probe length, 100shots. slow rise to 25% Pexc..
    reduce power to 45 nW
    high power was ca 450 nW
    Check +C1 peak at 75,963736, 490 Hz AOM FWHM.
    ID6799 rabi-flop high power, 100shots
    increase intensity-servo setpoint to 0.275 V
    turn on intra-cavity servo. signal 2.6 V
    peak now at 75.9637
    connect PDH-detector + following mini-circ amplifier to batteries. touching the cavity shifts frequency...

    ID6792 / run5, +C1 peak, 1ms probe 1 kHz scan 20 Hz step, 52pts, ca 300s
    Fit 445 Hz FWHM AOM (Fourier limit 800 Hz optical)


    ID6791 0.05 - 10 ms, 40steps, 200shots. tau_pi 1.1ms. damped oscillations.
    ID6790 flop, 0.05-2 ms, 25 steps, 100shots. extra faster modulation on the rabi-osc? tau_pi maybe 1ms
    go back to highest probe power possible, ca 390 nW

    ID6787 +C1 peak, 10ms probe 1kHz span, 10Hz step. ca 600s = 10min scan. run4
    Center 75.964994 MHz, FWHM 216 Hz AOM.
    Not a thing of beauty...

    ID6786 1-20ms, 100shots
    ID6785 rabi-flop 1ms-100ms, 100shots. rises to 25% at 10ms, then flat
    ID6784 rabi-flop 0.05-5ms, 100shots. slow rise, but low Pexc
    Reduce power with waveplate, to 12 nW
    ID6781 +C1 peak, 2ms probe 1kHz span 40Hz step, run2. Fit center 75.965011, FWHM 265 Hz AOM. Fourier limit 400 Hz optical

    ID6780 turn intensity-servo gain up.
    ID6779 flop 0.05 to 5 ms 100 shots with 100nW, visually slow rise to 25% Pexc, then flat.
    Reduce power with waveplate, to 100 nW
    Rabi-osc on +C1 at 75.96492
    ID6776 +C1 peak, 1.2ms probe 1 kHz span, 60 Hz step. Fit center 75.964989, FWHM 368 Hz AOM
    ID6774 18dB att, 0.05 - 10 ms, 25 steps, 500 shots. tau_pi=1.2 ms (visually), P 330nW
    ID6774 18dB attenuation, 0.05 - 5 ms probe, 25 steps, 100 shots. P674 ca 330 nW

    run1: wider scan 40 kHz AOM, 1ms probe, 18 dB probe att.

    2020-05-07-run1.txt

    ID6771 peak-scan, 2 ms, 20 Hz AOM step. Fourier limit ca 400 Hz FWHM optical
    fit gives 296 Hz FWHM AOM
    ID6770 AOM att 24 dB, rabi-osc.
    tau_pi 2 ms maybe

    ID6768 same rabi-osc, but now intensity-servo ON, at DC-setpoint 200 mV

    ID6767 rabi-osc. intensity servo off, detector DC-level 260 mV
    tau_pi ca 1.1 ms

    peak at 75.9582 MHz, 1 ms probe, fit gives center 75,958106, 451 Hz FWHM AOM
    measured 674nm thru chamber ~450 nW
    18 dB AOM-attenuation as usual

    Search for clock-transition. Double-pass zeeman AOM config changed, make sure it shifts 2*76 MHz
    (+1, +1) is the right diffraction from the AOM. Away from the SMA-connector, after the 45-deg mirror this means the +1 diffraction is towards the trap!

    uMotion contrasts A: 1.1% (24s), B: 1.4% (21s), C: 1.4% (22s)

    11:50:30Z try reload again. 1.6A (1.38W). all beams on. 422 Rb-lock checked
    11:55Z ion! sometimes looks cold, sometimes diffuse/blinking?
    strange blinking at first. then stable single ion!?

    11:31Z reload. 1.55A. All beams on.
    11:43Z no clear signal

    11:21Z load with 1.55A (1.27W)
    11:27Z see diffuse and weak cloud? Multiple ions!? (camera exposure 1.9s). Signal goes away at 208 MHz detunign - reappears at 185 MHz detuning
    only A-beam was on initially (most of loading time), all beams might be helpful?

    Built simple servo that measures 674-intensity at trap breadboard. Servo controls 178.5 MHz AOM drive level using variable attenuator.
    DC-signal ca 0.25 V (could be higher?), using uncoated UVFS plate as splitter.
    Detector is 'one-inch-detector' with S5791 photodiode, OPA657 opamp, and 1 MOhm transimpedance, bandwidth 4 MHz, using 1.9 MHz LPF on output.

    Trap RF 16.770 MHz, 70mVpp siggen, monitor 1.70 Vpp.

    Power in A-beam (through chamber):
    674: 260 nW (intensity servo off)
    422: 0.85 uW (185 MHz detuning)
    461: 200 uW, I=199.3 mA checked against helper-dispenser
    405: adjusted to 2.5 mW

    2020-05-05


    ID6635 SLI position 1+2 (SLI position 1 looks unstable)
    ID6634 (servo bumps at -30dB from carrier)
    change to 20dB + 10dB attenuation on toptica servo output (it was 20dB + 6dB)
    ID6632 rabi-flop (SLI position 2, it was at position 3) servo-bumps 25 dB down from carrier
    ID6631 rabi-flop (further increase gain, now see 2nd servo-bump peak also)
    ID6630 rabi-flop (slightly again increase PDH, gain)
    ID6629 rabi-flop (slightly increased PDH gain, gives clearer servo-bump peaks) - maybe rabi-osc now smoother?
    ID6626 rabi-flop, lower PDH gain, no clear servo-bump peaks

    Disconnect B-coils
    flop ID6622m 27dB att - rise to 25% then noise

    Try terminating (50R) the tickler-cable
    flop ID6621 (27 dB att)- nice rise to 2.5ms, then flat (incoherent?) noise at 25% Pexc

    +C2 (?) is strongest peak:
    18dB attenuation, tau_pi = 1.3 ms probe gave FWHM 333 Hz AOM (Fourier limit ~600 Hz optical)
    21dB attenuation, (ID6615 flop) tau_pi ~2 ms probe, gave 321 Hz FWHM AOM (Fourier 400 Hz optical)
    27dB attenuation, (ID6618 flop) not-so-great rise to max at maybe 5 ms, no clear oscillations. (Fourier limit 160 Hz optical).
    Noisy spectrum (like the top of the rabi-osc), FWHM 261 Hz AOM

    2020-05-04 May the Fourth be with you

    run9: leave long and wide run overnight
    Start below carrier C5, scan linearly up to and beyond upper axial secular sideband
    scan range [75.9379, 77.0629]. 5600 points (ca 200Hz AOM step, 400 Hz optical)
    with 600us probe Fourier limit is 1300 Hz so we get 5 points per FWHM
    16:41Z expected remaining duration is 25 ks = 7 hours


    ARTIQ NOTE: requesting clock-cycle scan of ca 2 MHz width and 200 Hz resolution gave error...

    Scan spectrum around expected secular side-bands
    ARTIQ NOTE: restarting the 'v2' clock-cycle program cancels the currently running program! ARGH.
    ARTIQ NOTE: running the tickler program messed up settings for the fiber-noise-cancelling DDS! ARGH.
    Go back to maximum probe power, ca 600us tau_pi now
    run6 carrier peaks around 75,9629 MHz, 45 kHz AOM range should capture C5 also?
    run7 upper r_lo sideband should be centered around 76,4451
    upper r_hi sideband should be centered around 76,4663
    upper axial sideband should be centered around 76,9314



    Look for secular frequencies with tickler. Helres monitor voltage is 1.71 Vpp
    Scans with 4 dB attenuation around 1 MHz
    one mode at 1.0067 MHz, lower at 0.9644
    Axial, 20dB att, many dips around 1.937 MHz
    reduce to 22dB att, still many peaks
    try 24 dB att, now maybe a single dip at 1.93692 MHz

    Try to reduce probe-power 10x, with the waveplate. Previous reading was 600 nW, now maybe <60 nW
    ID6485 100shot rabi-osc, to 5 ms. Now tau_pi = 2 ms maybe (not a good clear maximum)

    ID6486 Scan the line with 2ms probe, 40Hz AOM step (3 kHz span is too wide...)
    gaussian fit 75,965782 MHz (FWHM 305 Hz AOM), Fourier limit ca 400 Hz optical

    Try Rabi-osc on +C1
    Locate C1 at 75,965775 (FWHM 1131 Hz AOM)
    ID6480 rabi-osc to 1ms probe, 100shots - see only a rise to max Pexc?
    ID6481 to 10ms - too few points?
    ID6482 to 10ms, 50pts, 500 shots. Tau_pi ca 660 us. see attenuated oscillations out to 6-7 ms or so - good!?

    Probe length 400 us - power 'max'. Fourier limit ca 2 kHz optical
    scan for clock transition, 70kHz optical span, 400 Hz step
    line-center ca 75.9629 MHz (?). The 70 kHz (optical) scan is not wide enough to see C5
    NOTE1: run1 and run2 were unsuccessful because the fiber-stabilization servo was unlocked - need a monitor/alert for this!
    NOTE2: the (magnetic?) chair is now close to the door


    Detung-scan, 1000ms gate
    ID6472, A-beam
    ID6474, bright/dark/bright counts in 1000ms gate


    Move closer to target (97mm)
    A-beam counts in 10ms bright/dark 170/20 - with some darkening-improvements dark-counts go to 1-5

    Move camera closer, obj-to-cube now 102mm. re-focus imaging-optics (small z-movement)
    now PMT counts 180-190/10 in 10ms

    align ASE: A-beam counts in 10ms 80-90/4-5

    set detuning to 209 MHz
    A-beam counts in 10 ms ca 70/4 bright/dark
    uMotion 100 kphot: A 9.8% (18s), B 6.7% (47s), C 8% (28s)

    13:13:30Z Load ion, Dispenser 1.55A (1.26W)
    13:20:20Z ion! loading time 6min50s
    Image seems weak-ish on camera(?)

    Lock toptica to cavity. SDR tracker
    fbeat 23164630 Hz

    Goal: reposition camera focus-plane and PMT pinhole plane for better agreement, and thus better SNR
    Camera objective (front edge), to cube first edge 115mm, should be about 97 mm to have image-plane and PMT-pinhole agree

    A-beam power levels after chamber:
    405: 2.5 mW after adjustment
    422: 0.4 uW, 185 MHz detuning (power is decreased from max possible)
    461: 200 uW (4.645 kOhm, 199.32 mA)
    674: 660 nW

    Bias voltages back to previous values of 2020-04-28

    Helical resonator DC-modulator-box removed, N-Short grounding
    This raises the resonance frequency and Q. Resonance at 16.770 MHz. 70 mVpp drive gives 1.73 Vpp monitor

    Morning: RF investigations with Spectrum Analyzer..
    76 MHz drive to the Zeeman AOM
    357 MHz drive and 178.5 MHz VCO of the fiber-stabilizer
    16.7 MHz trap drive

    2020-04-28 Search for EMI

    Measured from tickler cable:
    Stable 'comb': 64 kHz, 124 kHz, 248 kHz, 372 kHz, 495 kHz, 620 kHz, 743 kHz, 867 kHz
    Intermittent noise: ~420 kHz, 320 kHz

    Record Vbias voltages before shutting them down
    Vdiff = [-5.5 0.2772 27.9984 44.4698]
    V(1:4) = [-5.77719 -5.2228 -16.4715 +72.4682]

    2020-04-27 Toptica

    Carrier peaks fit reasonably well to a sinc-squared model with B=1.52 uT (larger B-field than previously!)
    Linewidth is Fourier-limited, ca. 0.8/0.45ms = 1800 Hz

    Wide scan shows side-bands at both 525 kHz (why?) and 905 kHz (radial secular mode)

    Leave long and wide run overnight, ca14 hours, to maybe see secular sidebands.
    18 dB attenuation, tau_pi=tau_probe=450 us
    Ion lost mid-way, but some useful data was collected


    Shorter scans with sinc-squared visual fit:



    24 dB att, 1 ms probe, 516 Hz FWHM AOM (0.8/1ms = 800 Hz)
    ID6333 longer rabi-osc run 0.05 - 5 ms, 50pts, 500 shots, 18dB attenuation, high probe power
    Try 209 MHz 422-detuning?
    With high power now tau_pi = 350 us
    350us gives 1071 Hz FWHM AOM (0.8/350 = 2285 Hz)
    Test rabi-flop with different attenuations
    18 dB -> tau_pi = 400 us, ID6322
    24 dB -> tau_pi = 900 us, ID6323
    27 dB -> tau_pi = 1400 us, ID6324
    30 dB -> tau_pi = ? us, ID6325

    Increase intra-cavity feedback set-point to 3.0V - IR-amplifier ca 360 mA

    ID6313 scan at 2.5 ms probe, gives 277 Hz FWHM AOM

    intra-cavity power stabilizer ON, 2.27Vset
    ID6309 18 dB attenuation, 6 ms probe

    ID6307 18 dB attenuation, 6ms probe -> 328 Hz FWHM (AOM)
    24 dB attenuation, peak with 6 ms probe, ID6305 run5, ca 230 Hz FWHM (AOM) ? wide?

    small-power Rabi osc ID6304 0.05 - 20 ms, 50 pts, 100 shots, 24 dB attenuation
    Further reduce with DDS attenuation 24 dB
    Reduce probe-power with waveplate (works only for A-beam) to approx 10 nW
    search again for the peak (run4) ID6302

    Try intra-cavity power stabilization, set-point 2.34V - does not seem to stabilize power at trap-breadboard?
    back to 18dB attenuation, ca 510 nW probe, seems to fluctuate a lot?

    ID6295 0.05 - 20 ms probe, 27 dB DDS attenuation, 100 shots
    ID6294 Rabi-flop on 75,966785, 27 dB DDS attenuation, 100 shots
    ID6292 run3, search for peak again, 27 dB probe attenuation, 1.6 ms probe
    C=75,966785 (311 Hz FWHM AOM)


    ID6291 Rabi-flop with 27dB probe-attenuation (35nW?). 0.05 - 5 ms, 100 shots - abort, doesn't look very good..

    More attenuation:
    21dB 350 nW
    24dB 120 nW
    27dB 35 nW - try this?
    Re-measure probe power w. 18dB DDS attenuation: 610 uW (2-fold compared to start of experiments??)

    Rabi-flop with more statistics, 500 shots
    ID6286 Center 75.966945, probe 0.050 to 5 ms in steps of 0.05 ms

    run2/ID6284: 400us probe, 4 kHz span 50 Hz step. 81 frequencies.
    This might be Fourier limited? Max quantum jumps as high as 66/100?
    Center 75.966945, FWHM 975 Hz (AOM)

    re-check the peak-center with tau_pi=400us ID6283 or 2020-04-27-run1, 2kHz span/100Hz step -> only 21pts to scan

    ID6282 0.05 - 5 ms with 0.05 ms step (100pts), 100shots - oscillates with better contrast than before, at least out to 3-4 ms
    ID6281 rabi-flop 0.05-10ms probe - now we have too few tau_p-points (250us increase per point) -> aliasing?
    Now Rabi-oscillations, we should have 330nW/20nW = 16-fold more probe power
    ID6280 first rabi-flop. Now visually first max at tau_pi = 400 us

    Search for the +C1 peak
    Found a peak higher than expected, 75.966911 (FWHM 1283 Hz) with 1 ms probe

    uMotion contrasts (300 kphotons): A: 0.2% (45 s), B: 0.6% (90 s), C: 2% (69 s)
    Seems OK, no need to re-run uMotion optimization

    Set cooling detuning 210MHz

    Ion! At 10:54Z, loading time 7 min.
    Load ion. I = 1.55 A start at 10:47Z. 1.37 W dispenser power.
    RF: 70 mVpp drive, 1.56 Vpp monitor.
    405 power 3 mW at A exit.
    461 power 175 uW at A exit.
    422 power 0.450 uW at A exit. Reduced from maximum - for better temperature.
    Optimized 674 nm power on breadboard. Now 330 nW at A beam exit.
    Switched to Toptica as clock laser.
    Fiber-noise stabilizer on. Intra-cavity power stabilizer not on

    2020-04-24 First clock run!

    livestream: 128x speedup version
    livestream: 12 hour stream

    On 2020-04-23 a manual peak-fit gave -1.5(??) Hz, now we get -9.9(5) Hz
    Rough estimate of servo-noise, assuming 20 red/blue counts, poisson statistics, G=0.5 nodimensional gain, and 400 Hz FWHM
    0.5 * 400/2 * sqrt(40)/40 / SRS = 4e-14
    Observed noise is 5-fold larger (C2/C4 have worse poisson statistics)


    Y-axis below should be "f - SRS / Hz". The time-series should show aging of AHM3, but this is ca -2e-15/day or 0.4 Hz, not visible


    Visualization of the peak servo. In order to reach ion line-center with a downwards-drifting input-light, the servo compensates by drifting upwards. If we flip the sign of 2*Zeeman_AOM this matches the cavity drift well (assuming small AHM-drift)
    Traces offset for visual effect
    Note: this shows drift-compensated cavity frequency filtered from 1s counter data with sgolayfilt(f_optical_red,order=2, framelen=61) to ca. 60s, corresponding to the servo cycle duration.


    Magnetic field. Why is there a bias of -3 nT in the C2 and C4 estimates, compared to C1?


    Quantum jump counts. Stdev seems to follow Poisson statistics (?) and mean seems 50/50 red/blu distributed




    Clock laser plots during the run.
    Note1: the drift-compensation overcompensates, so the input to the peak-servo is a (slightly) decreasing frequency
    Note2: the predicted zeeman_aom line-center below does not take into account AOM-shift (ca 1200Hz)

    2020-04-23 Thursday

    Peak servo run starting ca 58962.5


    Qauntum jumps for the peaks:




    Manual search for C1, C2, C4 to prime the servo
    2.5 ms probe, this is tau_pi from the flipflops
    with lower 422-power we now have ca 70 counts / 10ms
    +C1 75,963436 (FWHM 234 Hz)
    -C1 75.956509 (220)
    +C2 75.966886 (240)
    -C2 75.9530925 (224)
    +C4 75.977215 (270)
    -C4 75.9427147 (273)
    B-field estimated around 1.23 uT
    Cycle length approx. 64 s.
    Measured 20nW probe in A-beam after chamber
    ca 1 uW 422-power
    ID6186 rabi-flop on +C1 at 75.96342 (shifted lower over lunch?)
    0.05ms to 20ms probe in 40 steps, 500 shots
    ID6187 detuning scan, 1000ms gate. background count estimate 350 cps

    lower 422-power to 0.35 uW, this halves fluorescence from 12 kcps to 6 kcps @ 210MHz detuning
    ID6190 detuning scan, 1000ms gate. background
    ID6191 rabi-flop (same settings as ID6187)



    Vopt = [0.27719 27.99845 44.46984]
    Here 300kphot contrasts: A 0.1% (32s), B 0.5% (54s), C 1.7% (40s) - an improvement!
    For previous C-matrices see 2020-04-16
    Note this run of the NPL-optimizer produced a C-matrix with opposite sign compared to earlier runs
    Cfit =-1*[   [-0.7486428 , -0.25501217, -0.16788016],
           [-1.92252664,  0.26163298, -0.16973487],
           [-0.86786103, -0.16630517,  0.19853788]]
    
    V0 = [  0.27719086,  27.99845308,  44.46984   ]
    

    ID6172 NPLv3 scan 4x4x4 points over lunch. 100kphot and voltage ranges around current voltages
    uMotion 300kphot: A 3.8% (31s), B 7.4% (76s), C 2.3% (40s)
    detuning 210 MHz as before

    +C1 peak with 8ms probe, A-beam (damped rabi-osc now shows max at around 8ms)
    tried spectra with B and C beams, but Pexc is low. B and C beams might be less focused?
    +C1 A-beam 8ms gives FWHM of 192 Hz AOM - not much narrower than previously...


    Memory usage during the peak-servo was rising up to 75% - investigate...
    Give servo-gain G in units of FWHM/2
    Try reversing sign of gain and increasing to G=0.05 (G=1 should be optimal according to theory)
    same starting centers/widths as y-day seem to work(?)
    add also B-field computation based on peak-separation - field is around 1.3 uT
    try G=0.5 (half of optimum, much higher than before) - seems to work

    Various gains/settings tried.
    Here +C1, -C1, and +C2, -C2 are shifted for visual effect. C1/2-centers are plotted without shifts

    Archive 8-hour plots of clock-laser here
    When the peak-servo works it should capture the residual cavity drift, here around +1 mHz/s AOM
    Long-term cavity drift here +18.8 mHz/s and we (over)compensate at 2*(-10.45) mHz/s = -20.9 mHz/s
    Drift compensation once per minute, a 60*(-10.45 mHz) = -627 mHz AOM step.

    2020-04-22 Wednesday

    sign of gain might be wrong here. Low red counts high blu counts should push AOM up!? (not down as we see here?)






    422-power was adjusted back to 'high' for these tests
    3.5ms probe length, 100shots per frequency. 10ms detection
    Scan of 4 peaks, cycle duration ca 79 seconds. 4 peaks * red/blu side = 8 frequencies to probe
    leave this overnight. at 80 seconds per cycle we can run about 540 cycles in 12 hours - so program 600

    Initial centers and FWHMs: (searching for these should be automated...)
    C1+ 75.963553 0.000187
    C1- 75.9562818 0.000174
    C2+ 75.9671580 0.000149
    C2- 75.952675 0.000179

    Try to program a peak-lock-servo
    While programming, saw mysterious dark disappearance act of the ion. Re-appeared maybe 5 min later... (a bit of 405nm was tried, but not sure if this had an effect?)
    The C1 pair takes about 40s to probe (red and blu side of +C1 and -C1)



    Try to lower ion-temperature, by reducing 422-power in the A-beam. Detuning kept at 210 MHz (-16 MHz optical)
    1 uW after chamber, 120/4 bright/dark in 10ms. Rabiflop ID6044, Detuning-scan ID6046
    0.3 uW after chamber, 60/2 bright/dark in 10ms. Rabiflop ID6054, Detuning-scan ID6051
    Rabi-flops on +C1 frequency from below 75,96353 MHz

    Locate the Zeeman peaks... (3.5ms probe, from yesterday)
    +C1 75,963533 (sigma 201 Hz) - run1 (6023)
    -C1 75.956234 (sigma 190) - run2 (6026)
    C1 center 75,959874
    Z-shift 3659 Hz AOM (1.307 uT)

    10minute mean driftcorr-freq is -330,4185162 MHz

    +C2 75,967154 (sigma 178 Hz) (run3)
    -C2 75,952596 (sigma 203 Hz) (run4)
    C2 center 75,959875
    Z-shift 7279 Hz AOM (1.3002 uT)

    +C4 75,978085 (sigma 236 Hz) (run5)
    -C4 75,941685 (sigma 178 Hz) (run6)
    C4 center 75,959885
    Z-shift 18200 Hz (1.3003 uT)

    at the end of this 6-peak quest we have drift-compensated 10min mean of -330.4185164 Mhz (a 0.2 Hz change from between C1 and C2 measurements)

    AHM3-shift estimated at 1241.1 Hz, ion-frequency at mj_squared=35/12 is 1239.6 Hz. (need to write down more significant digits from measurements in the future)
    Difference is -1.5 Hz or -3e-15





    10:05Z Check micromotion, 300kphot: A 3.6% (32s), B 7.8% (56s), C 2.0% (41s)

    10:00Z ion still stored. Overnight tracking-osc unlocked, and comb f0-signal seems weak

    2020-04-21 Tuesday

    16:40Z end. Leave ion trapped, turn off PMT

    And now for something copletely different, a note on the clock-laser performance over the past 4 hours:


    ID6013 rabi-flop on -C1 at 75.95584 MHz. 0.1 to 20ms probe. 500 shots

    C1 center is at 75.9594955 MHz Zeeman AOM. Latest value for drift-compensated light is -330.417743 MHz SRS
    Thus C1-line center is at: -330.417743 + 178.5 (fiber_AOM) + 2* 75.9594955 (Z_AOM) = +1248 Hz SRS
    AHM3 model from Bulletin-Atomique is -2.74565e-12 - 1.98725e-15 * (MJD - 58939)
    Now is roughly 58960.6, so we expect y_AHM3 = -2.78857e-12
    Naively AHM3_shift = y_AHM3 * SRS = 1240.3 Hz
    Remains to be explained (left as an exercise for the reader): observed +1248 Hz - expected 1240.3 = +7.7Hz (1.7e-14)

    Gaussian fit to +C1: 75.963145, FWHM 224 Hz (2.5 ms probe)
    estimate Z-shift of 3.5 kHz, so look for -C1 ca 7kHz lower?
    -C1 seems to be at 75.955846
    zeeman-split 7299 Hz AOM = 14598 Hz optical.
    2/5*muB*B/h = 5598.5 Hz/uT, so we have a field of 1.304 uT
    rabi-osc on -C1 suggest optimal exication at longer probe of 3.5ms

    15:12Z micromotion contrasts, 300kphot, A 2.7% (34s), B 5.5% (57s), C 2.9% (54s)

    ID5954 Rabi-flop, 500 shots, probe 0.1-20 ms, center 75,96318 MHz (+C1?)

    Modified ARTIQ programs: rabi-flop v2 where probe-pulses in order, clock-cycle with 2nd detection window, clock-cycle-plotting + ARITQ-logic-analyzer (coreanalyzer)view

    Ion still stored.

    2020-04-20 Monday


    No good ideas for overnight experiments..

    ID5866 10Hz scan of the line
    run29 ID5868 with 8ms probe with 18dB attenuation (too much!)
    run30 ID5869 with 8ms probe with 26dB attenuation
    run31 ID5870 2.5ms probe, 18 dB att


    ID5867 rabi-flop with probe-att 26 dB (maybe 10x less power)

    ID5884 rabi flopping at 75,963525MHz (cavity-shifted while closing doors?)

    2 ms probe
    run26 close doors (partly) to the cavity-box, fiber-servo ON, intensity-servo ON
    run24 fiber-stab ON (357 MHz from ARTIQ to mixer) and intensity-servo OFF
    run23 fiber-stab OFF (178.5MHz from ARTIQ to AOM) and intensity-servo OFF
    run22 fiber-stab OFF (178.5MHz from ARTIQ) and intensity-servo ON
    run21 fiber-stab OFF and intensity-servo ON (free-running VCO - not the way to do this measurement!)
    run20 fiber-stab ON and intensity-servo OFF
    run19 fiber-stab and intensity-servo on
    Shifts of +/-200 Hz due to touching cables going to the cavity -> cavity-system tilts??


    ID5814 rabi-flop 75.9633 MHz, 500 shots, 100 pts, 5.3s/shot = 44min

    run18 / ID5813, 1kHz scan, 25Hz resolution, 2ms probe


    ID5812 rabi-flop 75.96325 MHz, A-beam. 0.1-20 ms, 50 pts. 100shots. 2.6s/shots =

    run16 / ID5809, C-beam +C1 at 76.9633 MHz. ?/100 counts max
    run15 / ID5807, B-beam +C1 at 76.9633 MHz. 0/100 counts max
    run14 / ID5805, A-beam +C1 at 76.9633 MHz. 37/100 counts max

    run13 / ID5802 C-beam
    only 1 jump?? maybe repumper was overridden, always on??

    run12 / ID5799: Try the B-beam, it has different power-level for both 422nm and 674nm
    Only 2/100 jumps


    Look for motional side-bands
    run11/ ID5796 +C1 red side-band, 75.9633 - 0.457 = 75.5066
    30 kHz range, 100Hz step, 301 pts, 4.6s/pt


    +C1 Carrier 75.9633, 1 kHz scan 40Hz step, 25 pts. ID5790 / run6. MaxP=35/100
    blue: 75.9633 + 0.9424 = 76.9057 ID5791 / run7 - no dark counts - (but this is the wrong range!? AOM is double pass!)
    blue: 75.9633 + 0.9424/2 = 76.4345 ID5792 / run8 - single dark counts
    blue: 75.9633 + 0.9424/2 = 76.4345 ID5793 / run9 - 3kHz range, 100Hz step, 5 dark-counts at 76.4333
    run10 scan 76.42, 30kHz range, 100Hz step. This should capture both radial-lo and radial-hi modes? ID5794. 301 pts * 4.6s/pt = 23 minutes


    Lunch-break clock-spectrum, run5, ID5789
    36 kHz span, 40 Hz resolution, 900 pts, 4.6s/pt = 69 minutes
    rabi-flop shows tau_pi = 2.5 ms - so use that!





    ID5788 rabiflop again, 75.9632 MHz, 18 dB probe att. 0.1 - 10 ms probe, 50 pts.
    500 shots, 2.4s/shot
    clearout 10ms, cool 30ms, repumper 35ms, detect 10ms. bright/dark 120/10

    Check peak ID5786, 5787 (20Hz AOMres)



    ID5784 rabi flop at 75.96325 MHz. probe-att 18 dB - 0.1-5 ms probe, more points and 300 shots.
    ca 2.3 seconds/run (50 points).
    But here we might have lost the fiber-stabilization servo - because DDS-attenuation goes wrong..

    ID5783 rabi flop at 75.96325 MHz. probe-att 18 dB

    +C1 peak at 75.9632 MHz, 2 kHz span, 50 Hz resolution
    41 points, 4.6 s/point
    Good excitation-prob 37/100!
    Fourier limit should be 0.8/3ms = 267 Hz
    ID5752 (?)


    Clock-spectrum with A-beam
    Bright/dark levels 120/10 in 10ms
    Optimized clock-cycle with 30ms cooling, 35ms repumper. 3 ms probe, 75.9594 cecnter, 0.018 span, 100 Hz resolution. 18 dB probe attenuation
    179 points, 4.6s/point = 14 minutes


    Look for seculars with tickler
    0.92 to 9.96 MHz scan, 4dB att, shows dip at 0.9426 (upper radial?), 2nd radial not visible with A-beam
    B-beam scan 0.9 to 1.0 MHz: lower at 0.8964
    5dB att / 20ms detection: 0.89565 and 0.9424 MHz

    Vbias from last time (-5.5, 0.2563, 28.007, 44.332)
    uMotion 300kphot: A 2.7% (29s), B 4.2% (54s), C 3.0% (54s)
    Set detuning to 210 MHz. Fluorescence A strongest, B weakest, C middle.

    07:27:30Z load with 1.55A (1.34 W, slow rise to 1.36W) - 07:38:00Z ion signal! (10min30s)
    Camera exposure time 2 s, zoom out to see full image. Image has moving horizontal-stripes?
    Detuning 185 MHz

    A-beam power levels

    Clock-laser
    Beat 23.1612 MHz
    Driftcomp 79.762800 MHz
    Predicted line-center at 79.9594 MHz
    Intensity servo set-point 1.15V

    RF 70mVpp 16.709MHz, monitor 1.57Vpp

    Check 461nm. R=4.645kOhm. I=199.32mA (check with dispenser 5A, 2min)

    2020-04-16 Thursday

    Linear model for micromotion
    S = C *[ V - V0]
    where V0 are the optimal voltages
    2020-04-16 C-matrix (V0 = [ 0.25631175, 28.00738916, 44.33204023])
    Cfit 
          [[ 0.6155015 ,  0.25619819,  0.20220264],
           [ 1.68776957, -0.2430226 ,  0.16205872],
           [ 0.66540325,  0.13719351, -0.19730336]]
    
    Compare to 2019-12-19 C-matrix with V0 = [ 0.24380421, 26.67929547, 43.62223335]
    Cfit 
       [[ 0.6688455   0.22836817  0.14897131]
        [ 1.36885973 -0.19586379  0.12287398]
        [ 0.66160432  0.12331246 -0.15400281]]
    
    The inverse of Cfit gives voltage-directions (as columns) that change only one contrast-value
    These directions have changed by 2..7 degrees between the 2019-12-19 and 2020-04-16 datasets.
    Moving in the voltage-direction VA changes only the A contrast (not B or C). In other words VA lies both in the B-optimal and the C-optimal planes.
    Directions are (2019-12-19 and (2020-04-16) (V12dif, V34cm, V34dif)
    NOTE: changing voltages by this amount gives a contrast-change of +1 (100%) - not recommended experimentally, try only 10% or 1% of this vector


    Note1: VB and VC lie in the A-optimal plane (moving in VB or VC directions does not change A-contrast)
    Note2: moving along VA moves us 'maximally' away from the A-optimal plane


    All three planes and vectors:


    Zeeman AOM attenuation vs. 674nm power, measurements from 2019-01-31


    ca 14:30Z lost ion (why?) - shut down lasers

    Not-so-optimized clock-cycle in use:
    Estimated on/off for SRS shutter is 5ms, and for ASE 4ms-on, 8ms-off


    Distribution of PMT counts from the run4 dataset
    Poisson distributions fit by eye. 101 frequencies * 100 shots/freq = 10100 datapoints in histogram
    Fits roughly 285 counts/20ms bright state, and maybe 12 counts/20ms dark state
    Distribution of dark-counts inconclusive - maybe ambient light-conditions affect this


    run4 (ID5752)
    2 kHz range, 20Hz step, 101 points
    40ms probe, 24 dB DDS attenuation





    rabi-flop tests 5747, 5748
    18 dB DDS attenuation gave max P at ca 2.5 ms
    21 dB attenuation -> ID5749
    24 dB attenuation -> ID5750
    Inconclusive results - probability rises from 0 but then levels off (+noise). No clear oscillations

    focus on the +C1 peak at AOM 75.9532 MHz
    range 2 kHz, resolution 50 Hz, 41 points


    search again for clock transition 2020-04-16-run2
    same parameters as run1, but now we have the 2nd shield
    duration 180 points * 9 s/pt = 27 min


    Install 2nd magnetic shield. (no collar on the 1st shield)
    re-align PMT
    Contrasts 300kphot: A 2% (27s), B 0.3% (43s), C 3.3% (52s)

    Clock-transition search. 2020-04-16-run1 (ID 5729)
    Bright 300, dark 45, in 20ms detection window
    75.9595 MHz center, 18 kHz span, 100Hz resolution, 180 points.
    20 ms probe length


    Clock laser
    Clock-laser cavity intensity servo. Set-point at 1.4V. Amplifier at 370mA gives measurement also around 1.4V
    Amplifier current-limit is 450mA - servo may drive it beyond limit -> unlock
    Fiber-stabilization servo, reboot ARTIQ-DDS producing 357 MHz reference signal (careful when setting DDS attenuation?)

    Apply the optimal voltages (0.2563, 28.007, 44.332)
    Now get: 300khpot A 5% (26seconds), B 5.5% (48seconds), C 3% (43s)

    To get more fluorescence from B/C, set detuning to 210 MHz
    (ID5717) Try NPL algorithm around the optimum. 3x3x3 = 27 points. ranges 0.2...0.3 27.5...28.5
    A 9s, B 16s, C 14s = 40s/point * 27 points = 18 minutes
    Vopt [0.2563, 28.0074, 44.332]



    From visual fit, relative intensities A/B/C = 1.0 / 0.2 / 0.4
    Fluorescence collection efficiency now 1/1150 vs. 1/500 in 2019-12-05 data.

    NOTE: A-beam has enough power to produce a wider line! Lineshape is narrower with the weaker B and C beams.
    Detuning scans, from 205MHz -20...+14 MHz, 51 points.
    ID-5704, A-beam, 1000ms gate
    ID-5705 is PMT counts/20ms with ASE on/off (at 205MHz)
    ID-5707, B-beam, 1000ms gate, detuning scan
    ID-5708, PMT counts/20ms, B-beam, ASE on/off, 205MHz
    ID-5710, C-beam, 1000ms gate, detuning scan
    ID-5711, C-beam PMT counts/20ms, ASE on/off

    Now at (-5.5, 0.24, 28, 44.25)
    Try 300kphot, to confirm optimal point:
    A: 3.6 %, 3.3%, 3.9% (ca 32seconds to collect 300kphot)
    B: 4.5 % (73 seconds)
    C: 2.6 % (63 seconds)

    Try moving in A-optimal direction (+0.3, -0.45) -> (28, 44.25) gives A 3.6%, B 2%, C 3%
    Back to (0.24, 27.7, 44.7) -> A 3% (11s), B 12% (23s), C 11% (19s)

    09:53Z result is Vopt ( 0.458, 29.067, 44.961)
    This Vopt gives 100kphot contrasts A 48% (10seconds), B 18% (21seconds), C 20% (18seconds)

    08:50Z Try NPLv3 program. ID5660, 4x4x4 matrix = 64 points
    scan range 0.2..0.3 26.7..28.7 43.2..46.2
    measurement time A 10s, B 26s, C 19s = 55s, so duration ca 60min (?)

    repeat contrasts after B-focus adjustments
    A 3%, B 13%, C 13%

    Try to focus B-beam tighter, to get more fluorescence
    lens 1/8-turn closer gave 80-100 counts/20ms. 1/16-turn closer gave ~90 counts
    try 1/8-turn outwards: 60-80 counts
    (A beam gives >200counts, all beams ca 250 counts)

    scan V34dif to find A-optimum at (-5.5, 0.24, 27.7, 44.7)
    100kphot A 2.9%, B (very weak) 13%, C 12%

    try change of (+1, -1.5) = (27.7, 44.85)
    A 4%, B 14%, C 14%

    Try (25.7, 47.85) change is (-1, +1.5)
    A 3%, B 67%, C 70%

    A-optimum line slope is -1.5 V34diff / V34cm (result from 2019-12-04)
    Look for A-optimum V34dif-scan, optimum at (-5.5, 0.24, 26.7, 46.35)
    A 5%, B 57%, C 55%

    Try voltages from 2019-12-19 (0.24, 26.7, 43.6)
    A 48%, B (very weak, align?) 21%, C 7%

    Vbias from yesterday (-5.5, 0.2668, 22.705, 45.994)
    Contrasts (100kphot) A 74%, B 82%, C 75%

    Blink test with 200uw to 500uw of 405nm. Looks OKish with 2 levels 400/40 bright/dark in 20ms

    all beams ON give 150 counts/20ms at 185MHz detuning.
    205 MHz detuning gives 380 coutns/20ms

    (2nd try) more 405-power 2.56 mW
    1.55A 7:13:30Z start. (1.33W) 07:22:07Z signal! (~9min 30s)

    (1st try) 1.55A (1.32W .. 1.34W)
    07:01Z start. stop 07:11Z (10min) - no signal (405-power drifted down to 0.9mW)

    RF 70mVpp, monitor 1.61Vpp. No change in frequency 16.709MHz

    A-beam through chamber:
    422nm 1uW, detuning 185MHz for loading
    461nm 200uW
    405nm 1.2 mW (manually reduced fiber-coupling efficiency
    Try to load with lower 405nm power. With the old laser we had around 1 mW of 405nm

    2020-04-15 Wednesday

    Conclusion(?): new 405-laser significantly changed trap loading efficiency, in 4 tries we always got 2 or more ions (sometimes dark ions)

    matlab figure from 2019-12-04 with A-opt and B-opt lines
    modify this (offset lines to new values, same directions as before) and look for crossing of lines -> (0.2668, 22.705, 45.994)
    Contrasts 100kphot, A 4%, B 7%, C 10%

    strange blinking - two ions?

    Vbias (0.2668, 24.7, 43)
    A 5%, B 72%, C 64%%

    Vbias (0.2668, 24, 43)
    A 22%, B 64%, C 64%

    (4th load) load with 1.55A (1.3W) - 5min13s gave signal
    at 185MHz detuning blink test gives 100/30 bright/dark levels (in 20ms)
    at 205 MHz more fluorescence 250/30 bright/dark in 20ms

    (3rd load) load with 1.6A - weak signal - 5min 40s gave first a cloud, then suddenly 2 (?) cold ions?
    blink test (500uW of 405nm) shows 3 levels ca 0, 175, and 350 counts in 20ms
    emptied trap with 3Vpp of DC-modulation (26.8Vpp probe)

    Test at what DC-modulation we loose the ion. RF is at 70mVpp, monitor 1.59Vpp. back to V34 = (24.8, 43)
    1Vpp Dc-modulation (probe 9.2Vpp), ion moves sideways. sometimes large-jump, so we have a dark ion also?
    405 restored fluorescence again
    1.5Vpp (probe 13.5Vpp) - still see ion, moving sideways
    2Vpp (probe 18.5Vpp) - still see ion
    3Vpp - lost ion

    strangely ion goes dark, blinks, and then is recovered by 405/461 light. This repeats a few times.

    After optimizing beams, S-contrasts A=37%, B=77%, C=71%
    Detuning now 205 MHz

    Try DC-modulation at 1Vpp , probe 9.2Vpp. Not much movement of ion

    (2nd load) Try to load with 1.6A
    4min30s to signal - now maybe 1 ion?
    Blink test looks OK (2 levels) with 400uW of 405nm, levels are bright/dark 60/20 counts in 20 ms

    empty trap by lowering RF, glitch is between 37mVpp and 36mVpp

    go back to 2019-12-19 bias-voltages of (-5.5, 0.2668, 26, 43)
    Optimize beams, A and B-beams give fluorescence

    DC-modulation 1Vpp 500mHz, HV-amp (nominal gain 10x) probe sees 6.9Vpp
    2-ions move side-to-side 'a lot'
    old bias voltages from last time (-5.5V, 0.2668, 10, 17)

    (1st load) Try to load ion with 1.7A dispenser. 2s (!) exposure time on camera (no intensifier)
    loaded 2 (?) ions in 2min50s
    Blink-test seems to indicate 2 ions (blinking visible with lower 1mW 405-power)
    With PMT and still lower 405-power of 500uW see three-level PMT-trace (0, 1, 2 ions)
    ID5377 blink-test, 250uW 405-power. Mostly in 2-ion-fluorescence-state now. 200counts/20ms = 2 ions. ca 100counts/20ms = 1 ion. Dark is 20counts/20ms

    A-beam power levels, through the chamber

    add DC-modulator circuit to helica-resonator. Lower resonance at 16.709 MHz. 12VDC-supply for DC-modulator HV-amp exchanged.
    1.55 Vpp monitor voltage with 70mVpp from Keysight generator

    422nm Rb-signal 4 DIVs at 500mV/DIV. AOM to 185 MHz

    461nm check. R=4.645 kOhm, I=199.3 mA. check dispenser 5A for 2 min -> I=199.32mA

    2020-04-03 Friday

    20:58Z ion lost. storage time roughly 3.5 days (stream: https://www.youtube.com/watch?v=zy_vQ-RwKdc)

    16Z leave live-stream on for the weekend

    13:06Z Coffee...

    Look for secular modes, with the tickler. Removed 2nd shield to check tickler-connection - but it was connected OK to artiq-DDS
    Configuration: RF 70mVpp, monitor 1.71 Vpp, Vbias = (-5.5, 0.2668, 10, 17)
    All beams ABC on
    (see 2019-12-18) DDS att 4dB for the radial modes at 1 MHz, 10 dB for axial at 2 MHz
    Axial mode attenuation can be increased, signal is very clear on PMT. 16..17dB attenuation works
    With too much tickler, the ion heats up and takes a long time to recover into a cooled state
    typical axial scan range 1.9 to 2.1 MHz, 201 points, 40 ms detection time
    typical radial scan range 1.0 to 1.2 MHz, 401 points, 40ms gate, 4dB att

    When no unit is given, assume MHz in the tables below

    How does V12cm change z/r?
    Config radial lo radial hi axial z/r=axial/(0.5*(rhi+rlo))
    70mVpp, -5.5 V12cm 0.9774 1.021 1.994 z/r=2.00
    70mVpp, -4.5 V12cm 1.0055 1.048 1.963 z/r=1.91
    70mVpp, -3.5 V12cm 1.0324 1.0738 1.933 z/r=1.84

    How does RF-drive change seculars?
    Config radial lo radial hi axial z/r=axial/(0.5*(rhi+rlo))
    65mVpp, -5.5 V12cm 0.904 0.9512 1.8408 (broad peak with 18 dB, narrower with 19dB) z/r=1.98
    70mVpp, -5.5 V12cm (repeat of first row above) 0.9775 1.0217 1.992 z/r=1.99
    75mVpp, -5.5 V12cm (monitor 1.83 Vpp) 1.045 1.086 2.1325 (nothing with 20dB, clear peak with 18dB z/r=2.00

    Lower radial mode seems harder to detect? The fluorescence dip is shallower, requires more tickler


    12:01Z ion is still trapped.
    Note: all measurements y-day were with B-beam, it gave the best fluorescence. Beam alignment is very sensitive...

    2020-04-02 Thursday

    run8 (5307): longer 60ms probe, 2 Hz resolution, 400 Hz span, 200pts a 12 s per point
    This run was interrupted because the clock-laser unlocked from the cavity


    Visual fit to run7 data gives ca 60 Hz 1sigma AOM. Roughly 240 Hz optical FWHM.

    run7 (5308): focus on the +C2(?) peak, to get a linewidth
    30ms probe, 10 Hz resolution, 1 kHz span, 101 pts
    randomized scan! ca 9 seconds to probe 100 pulses


    smaller peak at 75.9641 (as before) - this might be +C1
    stronger peak at 75.96721 - this might be +C2
    calculations indicate line-center at 75.9608 MHz

    Try longer scan to see zeeman peaks
    run6: 75.96413 center, 7 kHz span, 20 Hz step. 20ms probe.
    No B-field-coil currents applied (I1,I2,I3 all zero)

    Install 2nd B-shield (NOTE: no collar-metglas around 1st shield, and inner temp-sensor is loose, no door-caps on 2nd shield)

    Try to scan peak with long 50ms probe and 5Hz AOM step
    center 75.96413, span 500 Hz.
    run4, ID5303

    Try Rabi-flopping program
    highest peak was at 75.96413 MHz, so probe there.
    ID 5300, up to 50ms probe, max was 26 jumps at 27ms probe
    ID 5301, shorter probes 1ms-20ms
    no oscillations seen - inconclusive?

    Search for the clock transition
    From PMT-tests set bright_count=160, dark_count=20 (20ms detection window)
    First try center 75.96 MHz, span 100 kHz, step 1 kHz, 101 points. 10ms probe pulse.
    seem to find a peak at 75.964, measure that with 100Hz resolution
    2020-04-02-run3 20ms probe, center 75.96412 MHz, 10Hz step, 1kHz range, data ID 5298
    2020-04-02-run2 10ms probe, center 75.9462 MHz, 10Hz step, 1 kHz range, data ID 5297

    clock-cycle program from 2019-12 "sweep clock cycle in kernel v2". Some adjustments to DDS frequency-setting were required (due to artiq-5 upgrade?)
    The clock cycle has parameters dt_clearout=10, dt_cool=50, dt_repumper=60, dt_detect=20 (in milliseconds)
    clock_probe=10000 (in microseconds
    N_rep=100 (number of pulses, AOM_att=18dB (zeeman AOM)

    (0.2668, 20, 17) gives A 7%, B 29%, C 11%

    Try manual searching/scanning for optimum bias-voltages
    Scan V34diff towards lower values, V34diff=17V seems to improve A-contrast to around 13%

    Contrasts manually, at old optimum point V1 = -5.7668, V2 = -5.2332, V3 = -16.79, V4 = 70.21600000000001
    cm/diff format: (-5.5, 0.2668, 26.713, 43.503) A 31%, B 53%, C 32%

    Try uMotion NPL-optimizer (v3 program), with old ranges (see 6x6x6 result from 2019-12-19). Now to -5.5 V12cm (old setting)
    Algorithm gives strange optimum voltages [13.05783089 -7.38635953 18.00593577]

    Vbias now at (-5, 0.2668, 26.713, 43.503) =(12cm, 12diff, 34cm, 34diff)

    manual umotion contrasts A 30%, B 56%, C 38%
    422-detuning was left at 205MHz yesterday and not touched since

    remove adjustable-iris and try 400um pinhole. Visual inspection indicates it is slightly smaller than the adjustable iris.
    initially gives ca. around 200/40 bright/dark
    In order to move the imaging-objective freely the 1st B-shield needs to be 'loose' i.e. not attached with the plastic M6 screws (but there is a chance of blocking beams if we move shield too much!)

    0713Z PMT SNR - could the fixed pinholes improve this? (we have 50/100/200/400um pinholes to try)
    Start of day bright/dark 300/80 in 20ms gate - with the variable iris set to maybe 1mm (?) diameter - lever-position marked with pencil.

    Tried pixelfly without camera-objective, with sensor placed where the intensifier was previously. But this did not give any better image
    reinstall camera-objective to where it was.

    Clock-laser was unlocked from cavity at around UTC midnight
    Was left on ca 18 local Wed evening

    2020-04-01 April Fools ion

    Tried uMotion histograms, ca 10-20% contrast with all beams

    70/30 splitter for PMT reinstalled. Align by placing fiber-end(red-helper-laser) at PMT position, and look for the fiber-spot on the wide-field camera (IR window). 422-bandpass removed for this
    PMT SNR maybe around 5, i.e. bright/dark in 20ms gate-time approx 300/60. Previously SNR was maybe 10, i.e. 500/50 photons in 20ms gate time

    Image intesifier suddenly broke??
    Intensifier removed, without moving pixelfly+camera-obj the ion can be seen, but image is not so great

    11:15Z single ion loaded with 1.7A dispenser in 2 min 30s
    Optimization of beams with ion-cloud (dispenser at 2.0A continuously during optimization)
    Higher trapping voltage (110mVpp from siggen) may help locate the center of the cloud
    Initially A and C beams cool the ion, after alignment also B-beam
    V12cm = -5 V was set (similar to 2019-12 setting)
    Moving around ASE-beam further increased fluorescence

    Power levels at A-beam output (after chamber)
    461: 230 uW
    422: 0.9 uW
    405: 7 mW

    RF drive 16.773 MHz, 70mVpp gives 1.79 Vpp monitor as usual. 2nd monitor 930mVpp

    Check 461 against Dispenser (5 A for 2 min). 4.645kOhm / 199.3mA

    After optmizations Rb error-signal 4 DIVs peak-peak, about 2x larger than before

    Clock laser
    beat note 23.162593(40) MHz
    Cavity -489.980 MHz SRS

    2020-03-27

    Trapping and optimization efforts.
    The PMT/Camera 70/30 splitter was removed.
    Ca 2 minutes at 2.0A consitently resulted in a diffuse cloud
    Image is of beams, rather than trap. Cloud fluorescence lifetime ~1-2 minutes.
    Both A-beam and ASE-beam were focused to their original (small spot) settings

    The Ando OSA wavelength accuracy was checked against Rb-locked 422 laser
    With 185 MHz AOM detuning the reading is 421.3 nm - Sr+ cooling is 412.67 nm (paper on lab wall)

    2020-03-24 Tuesday


    Try also enlarging the ASE-beam, by defocusing 1/4-turn.

    Power levels at A-beam output
    461: 215 uW
    422: 0.3 uW (ca 4uW before A-fiber)
    aligning fiberlaunch from 422 laser-breadboard helps, to 9uW before A-fiber
    After alignment 422 at exit: 0.8 uW
    405: >5 mW after warmup
    Previously optimized trap frequency: 16.773 MHz
    70 mVpp run, 1.79 Vpp monitor (probe1), 920 mVpp (probe2)

    Coolinglaser. Detuning 185 MHz as usual for loading. Rb error-signal ca 2 DIV-peak-peak (quite small?)
    Usual settings of I_ld 69.6 mA, and I_amp = -1800 mA

    Defocusing A-beam objective by 1/4-turn enlarges the beam to (over)fill the inner-electrode to outer-electrode gap - spot diameter maybe 2x larger than previously.
    461-settings to old values. R_tec=4.645 kOhm, I_ld around 200mA - check against dispenser (run for 2min at 5A): 199.3 mA

    2020-03-20 Friday

    no luck with trap loading...

    re-align camera (B-shield removal requires objective-removal), focus +750/2 um in from electrode focus.

    Measured power-levels, A-beam, through chamber
    422nm: 0.6 uW (Rb error-signal 500mV/DIV on scope, signal is ca 3 divs peak-peak)
    461nm: 250 uW (Rtec=4.603kOhm, Ilaser=198.3 mA)
    405nm: >5 mW (requires adjustment as laser warms up?)

    New 405-pointer warm-up test with OSA. Spectrum does not shift much during 10 minutes.
    Maximum at 405.1 nm - should match Sr ionization transition well?

    Both 1st and 2nd B-shields now removed - to align beams onto ion
    Wide-field camera installed - allows seeing ASE-spot (and blue spots) on electrode surfaces, for aligning

    2020-03-18 Wednesday

    New 405-pointer, measured 404.5 nm

    end-of-day
    Fbeat 23.162962 MHz
    Fcavity -490.006128 Hz

    mid-day:
    Clock-laser, old seed laser
    Beat 23.162984 MHz
    Fcavity -490.006425 MHz SRS

    2020-02-19 Wednesday

    waiting for 405nm replacement...

    clock-laser + beat was ok for almost 16 hours. Then laser unlocked (why?)
    applied drift-correction +20.9 mHz/s is OK.
    shutting down laser for now.

    2020-02-18 Tuesday

    0720Z: Clock laser was locked and beat-note OK over night. Tracker unlocked (why?) in the moring ca 9 local
    Beat note monitor (Siglent SA) shows <-25dBm peak - maybe too low for tracker?
    New applied drift-correction of +20.9 mHz/s is roughly correct.
    clocklaser TEC=17.176 kOhm, I=160.1mA. 1348-amp I=355.1mA, TEC=12.943kOhm
    optimize the beat note with the comb waveplates, maybe 5dB improvement to -21 dBm @ 1kHz RBW now.

    2020-02-17 Monday

    clock-laser TEC 17.175 kOhm. Leave locked

    Could not load ion. Tried up to 2.2 A dispenser. Removed 2nd B-shield to better align A-beam.

    RF drive 70 mVpp as usual, monitor is 1.77Vpp (alternative monitor 920mVpp). Resonance max is at 16.772 MHz

    NOTE: 10 MHz from comb-divider-distributor was in bad shape... needs fixing!
    Camera re-aligned to electrodes + 750um/2 focus.
    Check 422 Rb-lines OK, check 461nm helper-dispenser (I=198.46 mA now after check)
    First loading attempt (before lunch) did not work. Idisp=1.60A for 12 min, but no signal.

    Cooling laser: I=69.6mA, Amp -1850 mA.

    0820Z re-lock laser to cavity mode 1348.052 nm (-9.2 dBm at wavemeter). LD 160.1 mA, TEC 17.176 kOhm
    1348nm Amp TEC=12.943 kOhm, I=355.1 mA

    2020-02-16 Sunday

    1525Z clock-laser has now been locked to cavity, and comb beat-note OK, for 24+ hours. Good.
    Observed cavity drift +20.9 mHz/s
    Applied correction is too large +27.5 mHz/s, by about +6.6 mHz/s
    Apply a new (differential) correction of 2*(-3.3) mHz, to get a new total applied correction of +20.9 mHz/s
    active from around 1525Z
    Factors of 2 here are due to the double-pass AOM. Shift from DDS is applied twice to 674nm entering cavity.
    All AOMs effectively shift up, from the cavity at -490 MHz SRS, to reach line-center at ca 0 Hz SRS at the ion.

    2020-02-15 Saturday

    0850Z comb f0-lock had unlocked overnight. re-align waveplates (30dB f0 peak) and re-lock.
    optimize beat-note, but not much improvement, re-lock tracker.
    laser had apparently stayed locked since friday afternoon. leave running.
    seed TEC now at 17.178 kOhm (this is the one that eventually difts to a limit?).

    2020-02-14

    1520Z leave over weekend

    1350Z Switch back to old seed-laser at 1348nm + fiber pre-stabilizer
    lock to the right cavity-mode, fbeat 23.163 MHz

    The Toptica stays locked to the cavity for short periods. It is not clear how to use the "unlimited slow integrator" analog output from the servo as an input to the digital laser controller. The input can be used for controlling laser current, piezo, or temperature, with adjustable sensitivity. Additionally the Toptica seems sensitive to acoustic vibrations such as the lab-doors, and doors outside the lab.

    2020-02-04

    1355Z try lower gain of +1mA/V for slow servo?

    unlock - why??
    13:40Z FineIn1 has drifted to about -0.29 V
    12:02Z try larger range for ULI, now +/-5V (spec), +/-4V maybe in real life (was on +/-0.5V setting)
    With larger range, need to slow down integrator speed
    Fine-in1 controls current, +5 mA/V as before

    Stays locked for about 2hours. ULI goes to limit of +/-0.4V (this FALC output is connected to "Fine in1" on toptica)
    0839Z re-lock, try ARC 5mA/V.

    0830Z laser has jumped to TEM00 at 1348.053nm

    2020-02-03

    stayed locked for 2-4 hours. ARC was 3mA/V
    17:07Z leave locked overnight... now also logging siglent SA spectra

    Morning: Laser again in TEM02 mode, 1348.056nm
    comb seems ok

    2020-02-02, palindromic Sunday

    beat-note OK for <1 hour...

    12:48Z re-lock clock-laser to TEM00 1348.052nm, beat 23.13MHz
    re-lock tracer at ca 23.1606 MHz

    12:37Z clock-laser now locked to TEM02-mode (1348.045nm - far away!)? (left it Friday locked to TEM00).
    also comb rep-rate lock unlocked?/strange? -> re-lock comb rep-rate

    2020-01-31

    Measurement of Zeeman-AOM attenuation (dB) vs. optical power (W) on the breadboard, and estimated power delivered to ion *TODO results to be inserted here*

    405nm laser seems broken :( only 7 uW output!??

    Beat-note now 23.1602 MHz, up from 23.150 on 2019-12-17 and 23.145 on 2019-11-29
    NOTE comb measures drift-corrected light, not cavity! Residual drift is now:
    10 kHz (4.5e-11) per month = 300Hz (1.5e-12) per day = 4 mHz/s (IR) = 8 mHz/s (RED) roughly?
    Cavity at around -490.0944 MHz SRS. Short-term drift 24 mHz/s
    Drift-corrected light at -330.4160 MHz SRS, residual drfit -2 mHz/s (RED).

    After service-break, all systems seem operational, with imrovements/changes

    For 2019 results see index_2019